### v1

# On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity

Abstract.
We answer three questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [6]. First, we show that the exponent 22 in their unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by an explicit $2\times 2$ counterexample for $p>2$. Second, we construct a compact operator for which the corresponding singular-value inequality fails for every $p>2$ (indeed for all $p>0$ at a fixed pair of indices). Finally, we obtain isometry-orbit refinements of Euler’s quadrilateral identity for matrices and derive several Clarkson–McCarthy type inequalities for Schatten $p$-norms and related consequences.
Key words and phrases:
Symmetric modulus; singular values; unitary orbits; Euler identity; Schatten norms
2020 Mathematics Subject Classification:
15A60, 47A30, 47A57, 15A42
1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space, and let $\mathbb{K}$ denote the ideal of compact operators on $\mathcal{H}$. For $Z\in\mathbb{K}$, the singular values $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ are defined as the eigenvalues (counted with multiplicity) of $|Z|=(Z^{*}Z)^{1/2}$, listed in nonincreasing order. Let $\mathbb{M}_{m,n}$ be the set of all complex $m\times n$ matrices; when $m=n$, we write $\mathbb{M}_{n}$. For a matrix $X\in\mathbb{M}_{n}$, we adopt the compatible convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ may be viewed as an infinite nonincreasing sequence (and this agrees with the compact-operator notation).
We write
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2},\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite.
For $p>0$, the Schatten pp-(quasi)norm on $\mathbb{M}_{n}$ is
$\|X\|_{p}:=\big(\operatorname{Tr}|X|^{p}\big)^{1/p},\qquad X\in\mathbb{M}_{n}.$
When $1<p<\infty$, we denote by $q$ the conjugate exponent, i.e. $1/p+1/q=1$.
Bourin and Lee [6, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus.
Theorem 1.1 (Bourin–Lee).
Let $Z\in\mathbb{M}_{n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\leq U|\operatorname{Re}\,Z|U^{*}+V|\operatorname{Im}\,Z|V^{*}.$
They also posed the following question concerning the optimality of the exponent 22.
Question 1.2 ([6, Question 4.5]).
Let $p>0$. Suppose that, for every $Z\in\mathbb{M}_{n}$, one can find unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq U\,|\operatorname{Re}\,Z|\,U^{*}+V\,|\operatorname{Im}\,Z|\,V^{*}.$ (1.1)
Must we necessarily have $p\leq 2$?
Our first result answers Question 1.2 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
Theorem 1.3.
Let $n\geq 2$ and $p>0$. If for every $Z\in\mathbb{M}_{n}$ there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that (1.1) holds, then necessarily $p\leq 2$.
Remark 1.4.
For $n=1$, the statement of Question 1.2 is different: for a complex number $z$, the left-hand side of (1.1) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.1) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
Bourin and Lee [6] further raised the following question on singular values.
Question 1.5 ([6, Question 4.7]).
Does there exist $Z\in\mathbb{K}$ such that for every $p>2$ one has
$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j+k}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
for some pair of integers $j,k\geq 0$?
We answer this question in the affirmative; in fact, we produce an example where the same pair of indices works for all $p>0$.
Theorem 1.6.
There exists $Z\in\mathbb{K}$ such that for every $p>2$ one can find integers $j,k\geq 0$ for which
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$ (1.2)
Moreover, one may take $j=k=1$, in which case the inequality holds for all $p>0$.
Finally, Bourin and Lee [6] asked for matrix analogues of Euler’s quadrilateral identity.
Question 1.7 ([6, Question 4.8]).
What are the matrix versions, if any, of Euler’s quadrilateral identity
$\|x+y+z\|^{2}+\|x\|^{2}+\|y\|^{2}+\|z\|^{2}=\|x+y\|^{2}+\|y+z\|^{2}+\|z+x\|^{2}$
for three points $x,y,z\in\mathbb{C}^{n}$?
A direct computation yields the following Euler-type operator identity.
Proposition 1.8 (Euler operator identity).
Let $A,B,C\in\mathbb{M}_{n}$. Then
$|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|C+A|^{2}.$ (1.3)
We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$. Bourin and Lee [6] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following main results give isometry-orbit refinements of (1.3).
Theorem 1.9.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist isometries $U_{ij}\in\mathbb{M}_{4n,n}$, $1\leq i,j\leq 4$, such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|C+A|^{2}\oplus 0$ $=\frac{1}{4}\sum_{j=1}^{4}U_{1j}\,|A+B+C|^{2}\,U_{1j}^{*}+\frac{1}{4}\sum_{j=1}^{4}U_{2j}\,|A|^{2}\,U_{2j}^{*}$
$\quad+\frac{1}{4}\sum_{j=1}^{4}U_{3j}\,|B|^{2}\,U_{3j}^{*}+\frac{1}{4}\sum_{j=1}^{4}U_{4j}\,|C|^{2}\,U_{4j}^{*}.$ (1.4)
Theorem 1.10.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist three isometries $U_{1},U_{2},U_{3}\in\mathbb{M}_{3n,n}$ such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}=\frac{1}{3}\sum_{k=1}^{3}U_{k}\,\left(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\right)\,U_{k}^{*}.$
Theorem 1.11.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist four isometries $U_{1},U_{2},U_{3},U_{4}\in\mathbb{M}_{4n,n}$ such that
$|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\frac{1}{4}\sum_{k=1}^{4}U_{k}\,\left(|A+B|^{2}+|B+C|^{2}+|A+C|^{2}\right)\,U_{k}^{*}.$
Based on results obtained by the author in [8], we obtain the following two Clarkson–McCarthy type inequalities.
Corollary 1.12.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$,
$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}$ $\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big),$
$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}$ $\leq 3^{\frac{p}{2}-1}\big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\big).$
For $0<p\leq 2$, both inequalities reverse.
Corollary 1.13.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $1<p\leq 2$ ,
$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
where $q$ is the conjugate exponent of pp. For $p\geq 2$, the inequality is reversed.
Remark 1.14.
Set C=BC=-B in Corollary 1.12 and Corollary 1.13, we derive the classic Clarkson–McCarthy inequality in [7].
Organization of this paper. In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorem 1.6. In Section 4, we prove Theorems 1.9, 1.10, 1.11 and derive further related consequences.
2. Proof of Theorem 1.3
2.1. Trace necessary condition
Assume (1.1) holds for a given $Z\in\mathbb{M}_{n}$ and some unitaries $U,V$. Then, taking traces yields
$\operatorname{Tr}\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq\operatorname{Tr}|\operatorname{Re}\,Z|+\operatorname{Tr}|\operatorname{Im}\,Z|.$ (2.1)
Indeed, all terms in (1.1) are positive semidefinite, and for $0\leq A\leq B$ we have $\operatorname{Tr}A\leq\operatorname{Tr}B$. Moreover, the right-hand side of (1.1) has trace independent of the choice of $U,V$ by unitary invariance. Consequently, if (2.1) fails for some $Z$, then (1.1) fails for that $Z$ for all choices of $U,V$.
Thus, to disprove (1.1) for $p>2$, it suffices to find, for a given $p$, a matrix $Z$ violating (2.1).
$2.$ A $2\times 2$ counterexample for $p>2$
Fix $\theta\in(0,\pi/2)$ and set $c=\cos\theta$, $s=\sin\theta$. Define
$Z_{\theta}=\begin{pmatrix}c&0\\ -s&0\end{pmatrix}\in\mathbb{M}_{2}.$ (2.2)
Note that $Z_{\theta}$ is real, hence $Z_{\theta}^{*}=Z_{\theta}^{\mathsf{T}}$.
Claim 2.1.
For every $p>0$,
$\operatorname{Tr}\left(\frac{|Z_{\theta}|^{p}+|Z_{\theta}^{*}|^{p}}{2}\right)^{1/p}=\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}.$ (2.3)
Proof.
Compute
$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c\\ -s\end{pmatrix}(c,\,-s)=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
Hence $|Z_{\theta}|=(Z_{\theta}^{*}Z_{\theta})^{1/2}=P$. Similarly,
$Z_{\theta}Z_{\theta}^{*}=\begin{pmatrix}c^{2}&-cs\\ -cs&s^{2}\end{pmatrix}=:Q.$
A direct computation shows $Q^{2}=Q$. Moreover $Q=Q^{*}$ (indeed $Q$ is real symmetric), hence $Q$ is an orthogonal projection. Therefore $|Z_{\theta}^{*}|=(Z_{\theta}Z_{\theta}^{*})^{1/2}=Q$.
Since $P$ and $Q$ are projections, $P^{p}=P$ and $Q^{p}=Q$ for every $p>0$. Thus
$\left(\frac{|Z_{\theta}|^{p}+|Z_{\theta}^{*}|^{p}}{2}\right)^{1/p}=\left(\frac{P+Q}{2}\right)^{1/p}.$
We compute the eigenvalues of $(P+Q)/2$. We have
$P+Q=\begin{pmatrix}1+c^{2}&-cs\\ -cs&s^{2}\end{pmatrix},\qquad\operatorname{Tr}(P+Q)=2,\qquad\det(P+Q)=s^{2}.$
Hence the characteristic polynomial is $\lambda^{2}-2\lambda+s^{2}=0$, so the eigenvalues are $\lambda_{\pm}=1\pm\sqrt{1-s^{2}}=1\pm c$. Therefore the eigenvalues of $(P+Q)/2$ are $(1\pm c)/2$, and those of $\bigl((P+Q)/2\bigr)^{1/p}$ are $\bigl((1\pm c)/2\bigr)^{1/p}$. Taking the trace yields (2.3). ∎
Claim $2.$
For $Z_{\theta}$ in (2.2),
$\operatorname{Tr}|\operatorname{Re}\,Z_{\theta}|+\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=1+s.$ (2.4)
Proof.
We compute
$\operatorname{Re}\,Z_{\theta}=\frac{Z_{\theta}+Z_{\theta}^{*}}{2}=\begin{pmatrix}c&-s/2\\ -s/2&0\end{pmatrix},\qquad\operatorname{Im}\,Z_{\theta}=\frac{Z_{\theta}-Z_{\theta}^{*}}{2i}=\begin{pmatrix}0&-is/2\\ is/2&0\end{pmatrix}$.
For $\operatorname{Re}\,Z_{\theta}$, the characteristic polynomial is
$\lambda^{2}-c\lambda-\frac{s^{2}}{4},$
whose roots are
$\lambda_{\pm}=\frac{c\pm\sqrt{c^{2}+s^{2}}}{2}=\frac{c\pm 1}{2}.$
Since $c\in(0,1)$, we have $\lambda_{+}>0$ and $\lambda_{-}<0$, hence
$\operatorname{Tr}|\operatorname{Re}\,Z_{\theta}|=|\lambda_{+}|+|\lambda_{-}|=\frac{1+c}{2}+\frac{1-c}{2}=1.$
For $\operatorname{Im}\,Z_{\theta}$, the eigenvalues are $\pm s/2$, hence $\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=s$. Adding yields (2.4). ∎
Combining (2.1), Claims 2.1 and 2.2, we see that if (1.1) were to hold for $Z_{\theta}$, then necessarily
$\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}\leq 1+s.$ (2.5)
Claim 2.3.
Let $p>2$. Then there exists $\theta_{0}\in(0,\pi/2)$ such that for all $\theta\in(0,\theta_{0})$ the inequality (2.5) fails.
Proof.
Define
$\Phi(\theta):=\left(\frac{1+\cos\theta}{2}\right)^{1/p}+\left(\frac{1-\cos\theta}{2}\right)^{1/p}-(1+\sin\theta).$
As $\theta\to 0^{+}$ we have the standard expansions
$\sin\theta=\theta+O(\theta^{3}),\qquad\cos\theta=1-\frac{\theta^{2}}{2}+O(\theta^{4}).$
Hence
$\frac{1+\cos\theta}{2}=1-\frac{\theta^{2}}{4}+O(\theta^{4})\quad\Longrightarrow\quad\left(\frac{1+\cos\theta}{2}\right)^{1/p}=1+O(\theta^{2}),$
and
$\frac{1-\cos\theta}{2}=\frac{\theta^{2}}{4}+O(\theta^{4})=\frac{\theta^{2}}{4}\bigl(1+O(\theta^{2})\bigr)$
$\Longrightarrow\left(\frac{1-\cos\theta}{2}\right)^{1/p}=4^{-1/p}\theta^{2/p}\bigl(1+O(\theta^{2})\bigr)=4^{-1/p}\theta^{2/p}+O(\theta^{2/p+2}).$
Combining these estimates gives
$\Phi(\theta)=4^{-1/p}\theta^{2/p}-\theta+O(\theta^{2})+O(\theta^{2/p+2})\qquad(\theta\to 0^{+}).$
Dividing by $\theta$ yields
$\frac{\Phi(\theta)}{\theta}=4^{-1/p}\theta^{2/p-1}-1+O(\theta)+O(\theta^{2/p+1})\xrightarrow[\theta\to 0^{+}$$\infty$,
since $p>2$ implies 2/p1<02/p-1<0. Therefore $\Phi(\theta)>0$ for all sufficiently small $\theta$, i.e. (2.5) fails for all $\theta\in(0,\theta_{0})$ for some $\theta_{0}>0$. ∎
Proof of Theorem 1.3.
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (2.2). If (1.1) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (2.1). Claims 2.1 and 2.2 reduce (2.1) to (2.5), which fails for all sufficiently small $\theta$ by Claim 2.3. Hence (1.1) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
If $n>2$, embed $Z_{\theta}$ as $Z_{\theta}\oplus 0_{n-2}\in\mathbb{M}_{n}$; the same trace obstruction applies, so (1.1) fails in $\mathbb{M}_{n}$ as well. This contradicts the hypothesis, so necessarily $p\leq 2$. ∎
3. Proof of Theorem 1.6
Proof of Theorem 1.6.
Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define
$Ze_{1}=e_{1},\qquad Ze_{2}=e_{2},\qquad Ze_{3}=i\,e_{3},\qquad Ze_{n}=0\ \ (n\geq 4).$
Equivalently,
$Z=\operatorname{diag}(1,\,1,\,i,\,0,\,0,\dots).$
This is a rank-33 operator, hence $Z\in\mathbb{K}$.
Since $Z$ is diagonal (hence normal), we have $|Z|=|Z^{*}|$. Indeed,
$Z^{*}=\operatorname{diag}(1,\,1,\,-i,\,0,\,0,\dots),\qquad Z^{*}Z=\operatorname{diag}(1,\,1,\,1,\,0,\,0,\dots),$
and therefore
[…]
Consequently, for every $p>0$,
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=|Z|.$
Hence the singular values of the left-hand operator are
$\mu_{1}^{\downarrow}=1,\ \mu_{2}^{\downarrow}=1,\ \mu_{3}^{\downarrow}=1,\ \mu_{m}^{\downarrow}=0\ (m\geq 4).$
Next,
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\operatorname{diag}(1,\,1,\,0,\,0,\,0,\dots),$
so $\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0$. Also,
$\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}=\operatorname{diag}(0,\,0,\,1,\,0,\,0,\dots),$
so $\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0$.
Choose $j=k=1$, so that $1+j+k=3$ and $1+k=2$. Then for every $p>0$,
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}(|Z|)=1,$
while
$\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0+0=0.$
Therefore (1.2) holds strictly for all $p>0$, and in particular for all $p>2$. ∎
Remark 3.1.
The construction is intentionally minimal and answers the existence question in Question 1.5. Here $Z$ is finite rank and normal, so the left-hand side of (1.2) is independent of $p$ and equals a fixed singular value of $|Z|$. The strict inequality is obtained by selecting indices ($j=k=1$) for which the corresponding singular values of $\operatorname{Re}Z$ and $\operatorname{Im}Z$ vanish, while the chosen singular value of $|Z|$ remains nonzero.
4. Euler-type identities and isometry orbits
4.1. Euler-type identities
Let
$X:=\begin{pmatrix}A+B+C\\ A\\ B\\ C\end{pmatrix},\qquad Y:=\begin{pmatrix}A+B\\ B+C\\ C+A\\ 0\end{pmatrix}\in\mathbb{M}_{4n,n}.$
Consider the $4\times 4$ Hadamard matrix
$H=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ -1&1&1&-1\end{pmatrix},\qquad H^{*}H=I_{4},$
and set $\mathcal{U}:=H\otimes I_{n}\in\mathbb{M}_{4n}$. A direct block computation gives $X=\mathcal{U}\,Y$. Hence $X^{*}X=Y^{*}Y$, i.e. (1.3).
We use the following Pythagoras theorem for partitioned matrices proved by Bourin and Lee [5].
Lemma 4.1 ([5, Theorem 2.1]).
Let $m,n\geq 1$ and let $T=[T_{ij}]_{i,j=1}^{m}$ be an $m\times m$ block matrix with $T_{ij}\in\mathbb{M}_{n}$. Then there exist isometries $W_{ij}\in\mathbb{M}_{mn,n}$ such that
$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Remark 4.2.
The standard $m\times m$ grid partitioning is row/column compatible in the sense of [5], so the above lemma applies directly to block matrices $T=[T_{ij}]$ written with respect to such a partition.
Next, we prove Theorem 1.9.
Proof of Theorem 1.9.
Let
$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n},\qquad\mathcal{U}:=H\otimes I_{n}.$
Set
$T:=\mathcal{U}\,\Delta_{Y}\,\mathcal{U}^{*}=\frac{1}{2}\begin{pmatrix}A+B+C&A&B&C\\ A&A+B+C&-C&-B\\ B&-C&A+B+C&-A\\ C&-B&-A&A+B+C\end{pmatrix}$.
Since $\mathcal{U}$ is unitary,
$|T|^{2}=|\mathcal{U}\Delta_{Y}\mathcal{U}^{*}|^{2}=\mathcal{U}\,|\Delta_{Y}|^{2}\,\mathcal{U}^{*}=\mathcal{U}\,\operatorname{diag}(|A+B|^{2},\ |B+C|^{2},\ |C+A|^{2},\ 0)\,\mathcal{U}^{*}.$
Apply Lemma 4.1 with $m=4$ to the $4\times 4$ block matrix $T=[T_{ij}]$. We obtain isometries $W_{ij}\in\mathbb{M}_{4n,n}$ such that
$|T|^{2}=\sum_{i=1}^{4}\sum_{j=1}^{4}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
Every block $T_{ij}$ equals $(A+B+C)/2$, $\pm A/2$, $\pm B/2$ or $\pm C/2$. Hence
$|T_{ij}|^{2}\in\left\{\tfrac{1}{4}|A+B+C|^{2},\ \tfrac{1}{4}|A|^{2},\ \tfrac{1}{4}|B|^{2},\ \tfrac{1}{4}|C|^{2}\right\}.$
Moreover, each of (A+B+C)/2,A/2,B/2,C/2(A+B+C)/2,\ A/2,\ B/2,\ C/2 appears exactly four times among the sixteen blocks. Grouping the corresponding four terms and renaming the isometries, we obtain isometries $V_{ij}\in\mathbb{M}_{4n,n}$ such that
$|T|^{2}=\frac{1}{4}\sum_{j=1}^{4}V_{1j}\,|A+B+C|^{2}\,V_{1j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{2j}\,|A|^{2}\,V_{2j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{3j}\,|B|^{2}\,V_{3j}^{*}+\frac{1}{4}\sum_{j=1}^{4}V_{4j}\,|C|^{2}\,V_{4j}^{*}.$
Conjugating by $\mathcal{U}^{*}$ yields
$|\Delta_{Y}|^{2}=\frac{1}{4}\sum_{j=1}^{4}(\mathcal{U}^{*}V_{1j})\,|A+B+C|^{2}\,(\mathcal{U}^{*}V_{1j})^{*}+\cdots+\frac{1}{4}\sum_{j=1}^{4}(\mathcal{U}^{*}V_{4j})\,|C|^{2}\,(\mathcal{U}^{*}V_{4j})^{*}.$
Finally, each $\mathcal{U}^{*}V_{ij}$ is still an isometry in $\mathbb{M}_{4n,n}$ since
$(\mathcal{U}^{*}V_{ij})^{*}(\mathcal{U}^{*}V_{ij})=V_{ij}^{*}\mathcal{U}\mathcal{U}^{*}V_{ij}=V_{ij}^{*}V_{ij}=I_{n}.$
Renaming $U_{ij}:=\mathcal{U}^{*}V_{ij}$ gives (1.9). ∎
Naturally, we ask the following question.
Question 4.3.
Does there exist isometries $U_{i}\in\mathbb{M}_{3n,n}$ ($1\leq i\leq 4$) such that
$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}$ $=U_{1}\,|A+B+C|^{2}\,U_{1}^{*}+U_{2}\,|A|^{2}\,U_{2}^{*}$
$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
The answer to Question 4.3 is negative. In fact, take $A=B=C=I_{n}$. Then
$|A+B|^{2}=|2I_{n}|^{2}=4I_{n},\qquad|A+B+C|^{2}=|3I_{n}|^{2}=9I_{n},\qquad|A|^{2}=|B|^{2}=|C|^{2}=I_{n}.$
Hence the desired identity would become
$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
Since all terms on the right-hand side are positive semidefinite, we have
$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|=9$ while $\|4I_{3n}\|=4$, a contradiction. Therefore such isometries cannot exist in general.
Lemma 4.4 ([4, Lemma 3.4]).
Let
$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$
be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}=U\begin{pmatrix}X&0\\ 0&0\end{pmatrix}U^{*}+V_{0}\begin{pmatrix}0&0\\ 0&Z\end{pmatrix}V_{0}^{*}.$
As an immediate corollary of Lemma 4.4, we obtain the following result. For completeness, we provide an alternative proof.
Lemma 4.5.
Let $m\geq 2$ and let $H=[H_{ij}]_{i,j=1}^{m}\in\mathbb{M}_{mn}$ be positive semidefinite, written in $n\times n$ blocks. Then there exist isometries $V_{1},\dots,V_{m}\in\mathbb{M}_{mn,n}$ (i.e., $V_{k}^{*}V_{k}=I_{n}$) such that
$H=\sum_{k=1}^{m}V_{k}\,H_{kk}\,V_{k}^{*}.$ (4.1)
Proof.
Since $H\geq 0$, let $R:=H^{1/2}\in\mathbb{M}_{mn}$, so that $H=R^{*}R$. Partition $R$ into $m$ block columns of width $n$:
$R=\bigl[\,R_{1}\;R_{2}\;\cdots\;R_{m}\,\bigr],\qquad R_{k}\in\mathbb{M}_{mn,n}.$
Then
$H=R^{*}R=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (4.2)
Step 1: Identify the diagonal blocks. By block multiplication, the $(k,k)$ diagonal block of $H=R^{*}R$ is
$H_{kk}=R_{k}^{*}R_{k}\in\mathbb{M}_{n}^{+}.$
In particular, $|R_{k}|:=(R_{k}^{*}R_{k})^{1/2}=H_{kk}^{1/2}$.
Step 2: Polar decomposition and a partial isometry. Take the polar decomposition of each $R_{k}$:
$R_{k}=U_{k}\,|R_{k}|=U_{k}\,H_{kk}^{1/2},$
where $U_{k}\in\mathbb{M}_{mn,n}$ is a partial isometry satisfying
$U_{k}^{*}U_{k}=P_{k},\qquad P_{k}:=\mathrm{supp}(|R_{k}|)=\mathrm{supp}(H_{kk})$
(the orthogonal projection onto $\overline{\mathrm{Ran}}(|R_{k}|)$). Note that $H_{kk}^{1/2}(I_{n}-P_{k})=0$.
Step 3: Extend $U_{k}$ to a genuine isometry $V_{k}$. If $P_{k}=I_{n}$, then $U_{k}^{*}U_{k}=I_{n}$ and $U_{k}$ is already an isometry; set $V_{k}:=U_{k}$.
Assume now $P_{k}\neq I_{n}$. Set $d_{k}:=\mathrm{rank}(P_{k})\leq n$. Then $\dim\mathrm{Ran}(I_{n}-P_{k})=n-d_{k}$. Also, since $U_{k}$ is a partial isometry with initial projection $P_{k}$, its range projection $Q_{k}:=U_{k}U_{k}^{*}$ has rank $d_{k}$, hence
$\dim\mathrm{Ran}(I_{mn}-Q_{k})=mn-d_{k}\ \geq\ n-d_{k},$
using $m\geq 2\Rightarrow mn\geq n$.
Therefore, there exists an isometry
$W_{k}:\mathrm{Ran}(I_{n}-P_{k})\longrightarrow\mathrm{Ran}(I_{mn}-Q_{k}),$
i.e. a matrix $W_{k}\in\mathbb{M}_{mn,n}$ such that
$W_{k}^{*}W_{k}=I_{n}-P_{k},\qquad W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}.$
(For example, choose orthonormal bases of the two subspaces and map one to the other.)
Now define
$V_{k}:=U_{k}+W_{k}\in\mathbb{M}_{mn,n}.$
Then, using $U_{k}^{*}W_{k}=0$ (their ranges are orthogonal since $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$) and $W_{k}^{*}U_{k}=0$ (initial spaces are orthogonal since $W_{k}^{*}W_{k}=I-P_{k}$), we get
$V_{k}^{*}V_{k}=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n}.$
Hence $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$. Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
$V_{k}H_{kk}^{1/2}=(U_{k}+W_{k})H_{kk}^{1/2}=U_{k}H_{kk}^{1/2}=R_{k}.$
Step 4: Conclude the decomposition. From $R_{k}=V_{k}H_{kk}^{1/2}$ we deduce
$R_{k}R_{k}^{*}=V_{k}H_{kk}^{1/2}\,H_{kk}^{1/2}V_{k}^{*}=V_{k}H_{kk}V_{k}^{*}.$
Summing over $k$ and using (4.2) gives
$H=\sum_{k=1}^{m}R_{k}R_{k}^{*}=\sum_{k=1}^{m}V_{k}H_{kk}V_{k}^{*},$
which is exactly (4.1). ∎
Now, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
$W_{3}=\frac{1}{\sqrt{3}}\begin{pmatrix}I&I&I\\ I&\omega I&\omega^{2}I\\ I&\omega^{2}I&\omega I\end{pmatrix}=F_{3}\otimes I_{n}\in\mathbb{M}_{3n}.$
For
$D:=|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}=\operatorname{diag}(D_{1},D_{2},D_{3}),\quad\begin{cases}D_{1}=|A+B|^{2},\\ D_{2}=|B+C|^{2},\\ D_{3}=|A+C|^{2},\end{cases}$
one has
$W_{3}DW_{3}^{*}=\frac{1}{3}\begin{pmatrix}D_{1}+D_{2}+D_{3}&D_{1}+\omega^{2}D_{2}+\omega D_{3}&D_{1}+\omega D_{2}+\omega^{2}D_{3}\\ D_{1}+\omega D_{2}+\omega^{2}D_{3}&D_{1}+D_{2}+D_{3}&D_{1}+\omega^{2}D_{2}+\omega D_{3}\\ D_{1}+\omega^{2}D_{2}+\omega D_{3}&D_{1}+\omega D_{2}+\omega^{2}D_{3}&D_{1}+D_{2}+D_{3}\end{pmatrix}$. (4.3)
In particular, every diagonal block of $W_{3}DW_{3}^{*}$ equals $(D_{1}+D_{2}+D_{3})/3$.
Proof of Theorem 1.10.
Set $D=\operatorname{diag}(|A+B|^{2},|B+C|^{2},|A+C|^{2})$ and let $W_{3}$ be as above. Then $W_{3}DW_{3}^{*}\geq 0$ and, by (4.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$. A direct expansion shows
$D_{1}+D_{2}+D_{3}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
Denote this common sum by $T$. Hence $W_{3}DW_{3}^{*}$ is a $3\times 3$ block positive matrix with constant diagonal block $T/3$. Applying (4.1) to $H=W_{3}DW_{3}^{*}$ yields isometries $V_{1},V_{2},V_{3}\in\mathbb{M}_{3n,n}$ such that
$W_{3}DW_{3}^{*}=\sum_{k=1}^{3}V_{k}\,(T/3)\,V_{k}^{*}.$
Conjugating by $W_{3}^{*}$ and setting $U_{k}=W_{3}^{*}V_{k}$ (still isometries) gives the claim. ∎
Next, let $\eta=e^{2\pi i/4}=i$ and set
$W_{4}=\frac{1}{2}\begin{pmatrix}I&I&I&I\\ I&\eta I&\eta^{2}I&\eta^{3}I\\ I&\eta^{2}I&\eta^{4}I&\eta^{6}I\\ I&\eta^{3}I&\eta^{6}I&\eta^{9}I\end{pmatrix}=F_{4}\otimes I_{n}\in\mathbb{M}_{4n},$
so that $\eta^{2}=-1$, $\eta^{3}=-i$, $\eta^{4}=1$. For
$E:=|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\operatorname{diag}(E_{1},E_{2},E_{3},E_{4}),$
with $E_{1}=|A+B+C|^{2}$, $E_{2}=|A|^{2}$, $E_{3}=|B|^{2}$, $E_{4}=|C|^{2}$, we obtain
$W_{4}EW_{4}^{*}=\frac{1}{4}\bigl[M_{rs}\bigr]_{r,s=1}^{4},\qquad M_{rs}=E_{1}+\eta^{-(r-s)}E_{2}+\eta^{-2(r-s)}E_{3}+\eta^{-3(r-s)}E_{4}.$ (4.4)
In particular, every diagonal block of $W_{4}EW_{4}^{*}$ equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$.
Proof of Theorem 1.11.
Let $E=\operatorname{diag}(|A+B+C|^{2},|A|^{2},|B|^{2},|C|^{2})$ and let $W_{4}$ be the $4n\times 4n$ Fourier matrix above. Then $W_{4}EW_{4}^{*}\geq 0$ and, by (4.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$. As in the previous proof,
$E_{1}+E_{2}+E_{3}+E_{4}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}.$
Denote this sum by $S$. Thus $W_{4}EW_{4}^{*}$ is a $4\times 4$ block positive matrix with constant diagonal block $S/4$. Applying (4.1) to $H=W_{4}EW_{4}^{*}$ yields isometries $V_{1},\dots,V_{4}\in\mathbb{M}_{4n,n}$ such that
$W_{4}EW_{4}^{*}=\sum_{k=1}^{4}V_{k}\,(S/4)\,V_{k}^{*}.$
Conjugating by $W_{4}^{*}$ and setting $U_{k}=W_{4}^{*}V_{k}$ completes the proof. ∎
4.2. Clarkson–McCarthy type inequalities
Let $A,B,C\in\mathbb{M}_{n}$. Set
$Y_{1}:=\begin{pmatrix}A+B\\ B+C\\ C+A\end{pmatrix}\in\mathbb{M}_{3n,n}.$
Consider the $4\times 3$ partial Hadamard (isometry) matrix
$U=\frac{1}{2}\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{pmatrix},\qquad U^{*}U=I_{3},\qquad\mathcal{U}:=U\otimes I_{n}\in\mathbb{M}_{4n,3n}.$
Define
$X:=\mathcal{U}\,Y_{1}=\begin{pmatrix}A+B+C\\ B\\ A\\ -C\end{pmatrix}\in\mathbb{M}_{4n,n}.$
In particular, since |C|=|C||-C|=|C|, we have
$X^{*}X=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
Starting from the Euler-type square-sum relation and its orbit refinements, we now record several consequences for unitarily invariant norms and Schatten norms by using theorems in [8]. It suffices to notice the operator identity (1.3).
Corollary 4.6 (Unitary-invariant norms under Euler’s square-sum constraint).
[…]
For $0<p\leq 2$, both inequalities are reversed.
The preceding norm inequalities admit corresponding unitary-orbit refinements, which we state next.
Corollary 4.7 (Unitary-orbit Euler refinement).
Let $A,B,C\in\mathbb{M}_{n}$. If $p\geq 2$, then there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-1}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
Moreover, there exist unitaries $V_{1},V_{2},V_{3},V_{4}\in\mathbb{M}_{n}$ such that
$V_{1}|A+B+C|^{p}V_{1}^{*}+V_{2}|A|^{p}V_{2}^{*}+V_{3}|B|^{p}V_{3}^{*}+V_{4}|C|^{p}V_{4}^{*}\leq 3^{\frac{p}{2}-1}\big(|A+B|^{p}+|B+C|^{p}+|C+A|^{p}\big).$
For $0<p\leq 2$, both inequalities reverse.
Passing from general unitarily invariant norms to Schatten $p$-norms, we obtain Clarkson–McCarthy type estimates in Corollary 1.12. Using the explicit coefficient matrix $U$ above (hence $\max_{i,j}|u_{ij}|=\tfrac{1}{2}$), we obtain the mixed $\ell^{q}$$\ell^{p}$ Clarkson–McCarthy type estimate in Corollary 1.13.
4.3. A unitary-orbit Euler modulus inequality and singular value consequences
We use the following unitary subadditivity result for concave functions due to Aujla and Bourin [1, Theorem 2.1].
Lemma 4.8 (Aujla–Bourin).
Let $H,K\in\mathbb{M}_{n}$ be positive semidefinite and let $f:[0,\infty)\to[0,\infty)$ be monotone concave with $f(0)\geq 0$. Then there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$f(H+K)\leq Uf(H)U^{*}+Vf(K)V^{*}.$
Theorem 4.9.
Let $A,B,C\in\mathbb{M}_{n}$. Then there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$ (4.5)
Proof.
By (1.3),
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}=\sqrt{|A+B|^{2}+|B+C|^{2}+|C+A|^{2}}.$
Apply Lemma 4.8 with $f(t)=\sqrt{t}$ to $H=|A+B|^{2}$ and $K=|B+C|^{2}+|C+A|^{2}$ to get unitaries $U_{1},V_{1}$ with
$\sqrt{H+K}\leq U_{1}|A+B|U_{1}^{*}+V_{1}\sqrt{K}V_{1}^{*}.$
Apply Lemma 4.8 again to $|B+C|^{2}$ and $|C+A|^{2}$ to obtain unitaries $U_{2},V_{2}$ such that
$\sqrt{K}\leq U_{2}|B+C|U_{2}^{*}+V_{2}|C+A|V_{2}^{*}.$
Substitute and absorb conjugation by $V_{1}$ into $U_{2},V_{2}$ to obtain (4.5). ∎
Corollary 4.10.
[…]
Proof.
[…]
Corollary 4.11.
Let $A,B,C\in\mathbb{M}_{n}$. Then for all integers $j,k,\ell\in\{0,1,\dots,n-1\}$,
$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j}^{\downarrow}(A+B)+\mu_{1+k}^{\downarrow}(B+C)+\mu_{1+\ell}^{\downarrow}(C+A).$
Proof.
By Theorem 4.9, there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$
Set $X:=U|A+B|U^{*}$, $Y:=V|B+C|V^{*}$ and $Z:=W|C+A|W^{*}$. Then $X,Y,Z\geq 0$ and $\mu_{j}^{\downarrow}(X)=\mu_{j}^{\downarrow}(|A+B|)=\mu_{j}^{\downarrow}(A+B)$, and similarly for $Y,Z$.
We use the following singular-value inequality (see, e.g., [2]): for $P,Q\geq 0$ and $r,s\geq 0$,
$\mu_{1+r+s}^{\downarrow}(P+Q)\leq\mu_{1+r}^{\downarrow}(P)+\mu_{1+s}^{\downarrow}(Q).$ (4.6)
Applying (4.6) to $P=X$ and $Q=Y+Z$ gives
$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k+\ell}^{\downarrow}(Y+Z).$
Next apply (4.6) to $P=Y$ and $Q=Z$ to obtain
$\mu_{1+k+\ell}^{\downarrow}(Y+Z)\leq\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
Combining the last two inequalities yields
$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
Finally, since
$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq X+Y+Z$
and both sides are positive semidefinite, the monotonicity of singular values gives
$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z),$
which implies the desired estimate. ∎
$4.$ Symmetric norms, antinorms, and Ky Fan consequences
Corollary 4.12.
Let $A,B,C\in\mathbb{M}_{n}$.
  1. (a)
[…]
  2. (b)
    For every symmetric antinorm $\|\cdot\|_{!}$ on $\mathbb{M}_{n}^{+}$ (in the sense of [3]),
    $\Big\||A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big\|_{!}\geq\big\||A+B|^{2}\big\|_{!}+\big\||B+C|^{2}\big\|_{!}+\big\||C+A|^{2}\big\|_{!}.$
Proof.
Both inequalities follow from (1.3) and the (anti)triangle inequality on $\mathbb{M}_{n}^{+}$: a symmetric norm is subadditive, while a symmetric antinorm is superadditive. ∎
Corollary 4.13.
Let $A,B,C\in\mathbb{M}_{n}$ and $m=1,\dots,n$. Denote by $\mu_{j}^{\downarrow}(\cdot)$ (resp. $\mu_{j}^{\uparrow}(\cdot)$) the singular values in nonincreasing (resp. nondecreasing) order. Then
$\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\leq\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|C+A|^{2}\big),$
$\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\geq\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|C+A|^{2}\big).$
Proof.
Apply Corollary 4.12 to Ky Fan $m$-norms and Ky Fan $m$-antinorms on $\mathbb{M}_{n}^{+}$. ∎
Acknowledgments
Teng Zhang is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No. xzy022024045).

### v2

@@
-On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity
+On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity
 Abstract.
-We answer three questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [6].
-First, we show that the exponent 22 in their unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by an explicit $2\times 2$ counterexample for $p>2$.
-Second, we construct a compact operator for which the corresponding singular-value inequality fails for every $p>2$ (indeed for all $p>0$ at a fixed pair of indices).
-Finally, we obtain isometry-orbit refinements of Euler’s quadrilateral identity for matrices and derive several Clarkson–McCarthy type inequalities for Schatten $p$-norms and related consequences.
+We answer five questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [12], and thereby obtain a sequence of sharp results for matrices and compact operators.
+We first show that the isometry-orbit identity behind the matrix weighted parallelogram law cannot be extended beyond the parameter range $0\leq x\leq 1$, and that a counterexample already exists in dimension one.
+Next, we prove that the exponent $2$ in the Bourin–Lee unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by constructing an explicit $2\times 2$ counterexample for all $p>2$.
+We then construct a compact operator $Z$ for which the associated singular-value inequality fails for every $p>2$, in fact for a fixed choice of indices.
+We also settle a Thompson-type triangle problem for symmetric moduli: the inequality fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
+Finally, we develop isometry-orbit refinements of an Euler operator identity and derive sharp Clarkson–McCarthy type inequalities for Schatten $p$-norms, together with further consequences for unitarily invariant norms and singular values.
 Key words and phrases:
-Symmetric modulus; singular values; unitary orbits; Euler identity; Schatten norms
+Symmetric modulus; unitary orbits; Euler identity; Thompson inequality; Clarkson-McCarthy type inequalities
 2020 Mathematics Subject Classification:
-15A60, 47A30, 47A57, 15A42
+15A60, 47A30, 47A57, 15A42, 47B10
 1. Introduction
-Let $\mathcal{H}$ be a complex Hilbert space, and let $\mathbb{K}$ denote the ideal of compact operators on $\mathcal{H}$.
-For $Z\in\mathbb{K}$, the singular values $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ are defined as the eigenvalues (counted with multiplicity) of $|Z|=(Z^{*}Z)^{1/2}$, listed in nonincreasing order.
-Let $\mathbb{M}_{m,n}$ be the set of all complex $m\times n$ matrices; when $m=n$, we write $\mathbb{M}_{n}$.
-For a matrix $X\in\mathbb{M}_{n}$, we adopt the compatible convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ may be viewed as an infinite nonincreasing sequence (and this agrees with the compact-operator notation).
-We write
-$\operatorname{Re}Z=\frac{Z+Z^{*}}{2},\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
-For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite.
-For $p>0$, the Schatten pp-(quasi)norm on $\mathbb{M}_{n}$ is
-$\|X\|_{p}:=\big(\operatorname{Tr}|X|^{p}\big)^{1/p},\qquad X\in\mathbb{M}_{n}.$
-When $1<p<\infty$, we denote by $q$ the conjugate exponent, i.e. $1/p+1/q=1$.
-Bourin and Lee [6, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus.
+Throughout this paper, $\mathcal{H}$ denotes a complex Hilbert space, and $\mathbb{K}$ the ideal of compact operators on $\mathcal{H}$.
+For $Z\in\mathbb{K}$, we write $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ for the singular values of $Z$, namely the eigenvalues of $|Z|=(Z^{*}Z)^{1/2}$ (counted with multiplicity) arranged in nonincreasing order.
+Let $\mathbb{M}_{m,n}$ be the space of complex $m\times n$ matrices, and write $\mathbb{M}_{n}:=\mathbb{M}_{n,n}$.
+Let $X\in\mathbb{M}_{n}$.
+We adopt the convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ can be regarded as an infinite nonincreasing sequence, consistent with the compact-operator notation.
+We denote by $X^{\mathsf{T}}$ its transpose, and by $X^{*}$ its conjugate transpose (adjoint).
+For scalars $d_{1},\dots,d_{r}\in\mathbb{C}$, we denote by $\operatorname{diag}(d_{1},\dots,d_{r})$ the $r\times r$ diagonal matrix whose diagonal entries are $d_{1},\dots,d_{r}$ (and all off-diagonal entries are 0).
+More generally, if $X_{1}\in\mathbb{M}_{n_{1}},\dots,X_{k}\in\mathbb{M}_{n_{k}}$, then $\operatorname{diag}(X_{1},\dots,X_{k}):=X_{1}\oplus\cdots\oplus X_{k}$ denotes the block-diagonal matrix with diagonal blocks $X_{1},\dots,X_{k}$.
+1.1. A problem on matrix weighted parallelogram law
+For $A,B\in\mathbb{M}_{n}$ and $x\in\mathbb{R}$, we write $A\nabla_{x}B:=(1-x)A+xB$ for the weighted arithmetic mean of $A$ and $B$.
+We begin by recalling a unitary-orbit refinement of the weighted parallelogram identity due to Bourin and Lee [12].
+Theorem 1.1 ([12, Corollary 2.2]).
+Let $A,B\in\mathbb{M}_{n}$ and $0\leq x\leq 1$. Then there exist isometries $U,V,S,T\in\mathbb{M}_{2n,n}$ such that
+$|A\oplus B|^{2}=U\,|A\nabla_{x}B|^{2}U^{*}+V\,|B\nabla_{x}A|^{2}V^{*}+x(1-x)\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\}.$
+Taking traces in Theorem 1.1 yields
+$\|A\oplus B\|_{2}^{2}=\|A\nabla_{x}B\|_{2}^{2}+\|B\nabla_{x}A\|_{2}^{2}+2x(1-x)\|A-B\|_{2}^{2}$
+for $0\leq x\leq 1$. Since the right-hand side of that trace identity is a polynomial in $x$, it extends to all $x\in\mathbb{R}$. This motivates the following question.
+Question 1.2 ([12, Question 2.3]).
+Does Theorem 1.1 hold for every $x\in\mathbb{R}$?
+We first answer Question 1.2 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.1 cannot be removed and therefore is sharp (see Theorem 1.3 below).
+Theorem 1.3.
+Fix $x\in\mathbb{R}\setminus[0,1]$. Then Theorem 1.1 fails for this $x$ and some $A,B\in\mathbb{M}_{n}$. In fact, a counterexample exists already for $n=1$.
+1.2. Symmetric moduli and three related questions
+For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite. While for $A,B\in\mathbb{M}_{n}$ one has the triangle inequality
+$\|A+B\|_{1}\leq\|A\|_{1}+\|B\|_{1},$
+simple examples show that the inequality
+$|A+B|\leq|A|+|B|$
+need not hold in the positive semidefinite order. However, Thompson [20] observed that there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$|A+B|\leq U|A|U^{*}+V|B|V^{*},$ (1.1)
+which, upon taking traces, yields the triangle inequality as a consequence.
+Such “unitary-orbit” inequalities have since been generalized and extended to a variety of settings; see the surveys by Bourin, Harada and Lee [7, 9].
+In this paper, we generalize (1.1) to the setting of rectangular matrices (see Lemma 5.2) in order to address Question 1.11.
+Recall the Cartesian decomposition of a matrix $Z\in\mathbb{M}_{n}$:
+$Z=\operatorname{Re}Z+i\,\operatorname{Im}Z,\qquad\operatorname{Re}Z:=\frac{Z+Z^{*}}{2},\quad\operatorname{Im}Z:=\frac{Z-Z^{*}}{2i}.$
+Note that $|\operatorname{Re}Z|^{2}+|\operatorname{Im}Z|^{2}=\left(|Z|^{2}+|Z^{*}|^{2}\right)/2$.
+Bourin and Lee [12, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
 Theorem 1.4 (Bourin–Lee).
@@
 $\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\leq U|\operatorname{Re}\,Z|U^{*}+V|\operatorname{Im}\,Z|V^{*}.$
-They also posed the following question concerning the optimality of the exponent 22.
-Question 1.2 ([6, Question 4.5]).
+They also posed the following question concerning the optimality of the exponent $2$.
+Question 1.5 ([12, Question 4.6]).
 Let $p>0$. Suppose that, for every $Z\in\mathbb{M}_{n}$, one can find unitary matrices $U,V\in\mathbb{M}_{n}$ such that
@@
 Must we necessarily have $p\leq 2$?
-Our first result answers Question 1.2 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
+Our second result answers Question 1.5 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
 Theorem 1.6.
@@
 Remark 1.7.
-For $n=1$, the statement of Question 1.2 is different: for a complex number $z$, the left-hand side of (1.1) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$.
+For $n=1$, the statement of Question 1.5 is different: for a complex number $z$, the left-hand side of (1.2) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$.
 Hence (1.2) holds for all $p>0$ in dimension 11.
@@
 Bourin and Lee [12] further raised the following question on singular values.
-Question 1.5 ([6, Question 4.7]).
+Remark 1.8.
+In [12, Corollary 4.7], the indices on the right-hand side appear with a typographical slip.
+Starting from the operator inequality in [12, Corollary 4.3] and applying the Weyl inequality for singular values,
+$\mu_{1+j+k}^{\downarrow}(A+B)\leq\mu_{1+j}^{\downarrow}(A)+\mu_{1+k}^{\downarrow}(B)\qquad(A,B\geq 0),$
+one obtains the corrected bound
+$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$ (1.3)
+Accordingly, the natural corrected version of [12, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
+$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
+for some integers $j,k\geq 0$.
+Question 1.9 ([12, Question 4.8]).
 Does there exist $Z\in\mathbb{K}$ such that for every $p>2$ one has
-$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j+k}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
+$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$ (1.4)
 for some pair of integers $j,k\geq 0$?
-We answer this question in the affirmative; in fact, we produce an example where the same pair of indices works for all $p>0$.
+We answer Question 1.9 in the affirmative.
 Theorem 1.10.
-There exists $Z\in\mathbb{K}$ such that for every $p>2$ one can find integers $j,k\geq 0$ for which
-$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$ (1.2)
-Moreover, one may take $j=k=1$, in which case the inequality holds for all $p>0$.
+There exists $Z\in\mathbb{K}$ and fixed integers $j,k\geq 0$ such that for every $p>2$, (1.4) holds. In our construction one may take $(j,k)=(2,0)$.
+Let
+$|Z|_{\mathrm{sym}}:=\frac{|Z|+|Z^{*}|}{2},\qquad|Z|_{\mathrm{qsym}}:=\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{1/2}$
+be the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$, respectively.
+Now, we turn to a related problem raised by Bourin and Lee concerning Thompson-type triangle inequalities for symmetric moduli, compared with (1.1).
+Question 1.11 ([12, Question 4.9]).
+Denote by $|Z|_{\mathrm{(q)sym}}$ either the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$.
+Let $X,Y\in\mathbb{M}_{n}$.
+Does the Thompson triangle type inequality
+$|X+Y|_{\mathrm{(q)sym}}\leq U\,|X|_{\mathrm{(q)sym}}\,U^{*}+V\,|Y|_{\mathrm{(q)sym}}\,V^{*}.$ (1.5)
+hold for some unitary $U,V\in\mathbb{M}_{n}$ ?
+Our next theorem provides a complete answer to Question 1.11: the inequality (1.5) fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
+Theorem 1.12.
+Let $X,Y\in\mathbb{M}_{n}$.
+  1. (i)
+    In general, there do not exist unitaries $U,V$ such that
+    $|X+Y|_{\mathrm{sym}}\leq U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}.$
+  2. (ii)
+    There exist unitaries $U,V$ such that
+    $|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*}.$
+1.3. Matrix analogues of Euler’s quadrilateral identity
 Finally, Bourin and Lee [12] asked for matrix analogues of Euler’s quadrilateral identity.
-Question 1.7 ([6, Question 4.8]).
+Question 1.13 ([12, Question 4.10]).
 What are the matrix versions, if any, of Euler’s quadrilateral identity
@@
 Bourin and Lee [12] suggested exploring isometry/unitary-orbit refinements of Euler-type identities.
-Our following main results give isometry-orbit refinements of (1.3).
+Our following three results give isometry-orbit refinements of (1.6).
 Theorem 1.15.
@@
 $|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\frac{1}{4}\sum_{k=1}^{4}U_{k}\,\left(|A+B|^{2}+|B+C|^{2}+|A+C|^{2}\right)\,U_{k}^{*}.$
-Based on results obtained by the author in [8], we obtain the following two Clarkson–McCarthy type inequalities.
+Naturally, we ask the following question.
+Question 1.18.
+Does there exist isometries $U_{1},U_{2},U_{3},U_{4}\in\mathbb{M}_{3n,n}$ such that
+$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}$ $=U_{1}\,|A+B+C|^{2}\,U_{1}^{*}+U_{2}\,|A|^{2}\,U_{2}^{*}$
+$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
+Unfortunately, the answer to Question 1.18 is negative. In fact, take $A=B=C=I_{n}$. Then
+$|A+B|^{2}=|2I_{n}|^{2}=4I_{n},\qquad|A+B+C|^{2}=|3I_{n}|^{2}=9I_{n},\qquad|A|^{2}=|B|^{2}=|C|^{2}=I_{n}.$
+Hence the desired identity would become
+$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
+Since all terms on the right-hand side are positive semidefinite, we have
+$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
+But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|_{\infty}=9$ while $\|4I_{3n}\|_{\infty}=4$, a contradiction. Therefore such isometries cannot exist in general.
+1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
+For a Hilbert space $\mathcal{H}$, let $S_{p}$ denote the Schatten $p$-class.
+For $p>0$, we say that $X\in S_{p}$ if $X$ is a compact operator on $\mathcal{H}$ whose singular values $(\mu_{k}^{\downarrow}(X))_{k\geq 1}$ satisfy $\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}<\infty$ .
+In this case, the Schatten pp-(quasi)norm is defined by
+$\|X\|_{p}:=\Big(\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}\Big)^{1/p}.$
+(When $0<p<1$, $\|\cdot\|_{p}$ is a quasi-norm.)
+Building on the Euler operator identity (1.6), we obtain the following sharp Clarkson–McCarthy type inequalities.
+Theorem 1.19.
+Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
+$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big).$
+For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
+As a complement to Theorem 1.19, numerical evidence suggests the following sharp inequality. For a weaker bound with coefficient $3^{\frac{p}{2}-1}$, see Theorem 7.10.
+Conjecture 1.20.
+Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
+$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\leq\frac{3^{p-1}+1}{2^{p}}\left(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\right).$
+For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $\frac{3^{p-1}+1}{2^{p}}$ is optimal: equality holds for all $p>0$ by taking $A=B=C\neq 0$.
+Classical Hlawka-type inequalities compare norms (or their powers) of vector sums with those of their pairwise sums.
+A standard form of Hlawka’s inequality [17] asserts that if $V$ is an inner product space and $x,y,z\in V$, then
+$\|x+y\|+\|y+z\|+\|z+x\|\ \leq\ \|x\|+\|y\|+\|z\|+\|x+y+z\|,$
+where $\|\cdot\|$ denotes the norm induced by the inner product; equivalently, the inequality holds for any norm satisfying the parallelogram identity.
+Numerous refinements and variants replace $\|\cdot\|$ by powers $\|\cdot\|^{p}$ and/or adjust the constants.
+Theorem 1.19 and Conjecture 1.20 can be viewed as a Schatten–$p$ (i.e.
+noncommutative $L_{p}$) analogue: it features a sharp $p$-dependent constant and exhibits a change in behavior at $p=2$, as is typical for inequalities governed by the uniform convexity/concavity of $L_{p}$ spaces.
+As a further refinement of Theorem 1.19, we establish the following sharp mixed $\ell_{q}$$\ell_{p}$ Clarkson–McCarthy type inequalities.
+Theorem 1.21.
+Let $A,B,C\in S_{p}$. Then for $1<p\leq 2$ ,
+$\|A+B+C\|_{p}^{q}+\|A\|_{p}^{q}+\|B\|_{p}^{q}+\|C\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Big)^{q/p},$
+$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
+where $q$ is the conjugate exponent of $p$.
+For $p\geq 2$, these two inequalities are reversed.
+Moreover, the constant $2^{1-\frac{q}{p}}$ is optimal: equality holds for all $p>1$ by taking $A=B=-C\neq 0$.
+Remark 1.22.
+Setting C=BC=-B in Theorems 1.19 and 1.21 recovers the classical Clarkson–McCarthy inequality of [16].
+For a complete characterization of the validity of Clarkson–McCarthy type inequalities, see the author’s paper [21, Corollary 1.14] and the following Theorem 1.23 (also see [21, Theorem 1.16]).
+For further results and developments on Clarkson–McCarthy type inequalities, we refer the reader to [5, 6, 9, 10, 15, 21, 22].
+We now record the following mixed-norm inequality (Theorem 1.23), which strengthens the result in the author’s earlier work [21, Theorem 1.16].
+Unlike the approach in [21], which applies the three-lines theorem to a suitable trace functional and then invokes a duality argument, we present a new proof based on a vector-valued Riesz–Thorin type complex interpolation scheme.
+Theorem 1.23 (Mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimates).
+Let $U=[u_{ij}]\in\mathbb{M}_{s,t}$ and set
+$\mu:=\max_{1\leq i\leq s,\ 1\leq j\leq t}|u_{ij}|.$
+For $1<p<\infty$, let $q$ be the conjugate exponent of $p$, i.e. $\frac{1}{p}+\frac{1}{q}=1$. Define the linear map on $t$-tuples
+$T_{U}:\ \ell_{p}^{t}(S_{p})\longrightarrow\ell_{q}^{s}(S_{p}),\qquad T_{U}(Z_{1},\dots,Z_{t}):=\Big(\sum_{j=1}^{t}u_{1j}Z_{j},\ \dots,\ \sum_{j=1}^{t}u_{sj}Z_{j}\Big).$
+We will regard $T_{U}$ as an operator between $\ell_{p}$- and $\ell_{q}$-sums when equipped with the corresponding norms.
+  1. (i)
+    Assume that $U$ is a contraction, i.e. $\|U\|_{\infty}\leq 1$ (in particular, this holds if $U^{*}U=I_{t}$ or $UU^{*}=I_{s}$). Then for $1<p\leq 2$,
+    $\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{p}\Bigg)^{1/p},$ (1.8)
+    and the inequality is reversed for $p\geq 2$ if $U^{*}U=I_{t}$.
+  2. (ii)
+    Assume in addition that $U^{*}U=I_{t}$ (i.e. $U$ is an isometry). Then for $1<p\leq 2$,
+    $\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{p}\Bigg)^{1/p},$ (1.9)
+    and the inequality is reversed for $p\geq 2$.
+Theorem 1.23 has several interesting consequences.
+For example, Audenaert–Kittaneh’s conjecture [1, Section 8.1, Conjecture 7] (proved by the author in [22]) states that if $A_{1},\ldots,A_{n}\in S_{p}$, then for $1<p\leq 2$,
+$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\ \leq\ n\left(\sum_{i=1}^{n}\|A_{i}\|_{p}^{p}\right)^{q/p},$ (1.10)
+where $q$ is the conjugate exponent of $p$; for $p\geq 2$ the inequality is reversed.
+Moreover, the constant $n$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
+This inequality can be recovered from Theorem 1.23 by choosing the isometry
+$U=\frac{1}{\sqrt{n}}\begin{pmatrix}1&1&\cdots&1\\ 1&-1&\cdots&0\\ 1&0&\ddots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1\end{pmatrix}\in\mathbb{M}_{\frac{n(n-1)+2}{2},\,n},$
+whose rows encode the combinations $\sum_{i=1}^{n}A_{i}$ and $A_{i}-A_{j}$. Moreover, Theorem 1.23 yields the following sharp reverse-type companion to (1.10). For more details, see Subsection 7.4.
 Corollary 1.24.
+Let $A_{1},\ldots,A_{n}\in S_{p}$. Then for $1<p\leq 2$,
+$\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\leq n^{-q/p}\left(\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{\frac{q}{p}},$ (1.11)
+where $q$ stands for the conjugate exponent of $p$.
+For $p\geq 2$, (1.11) is reversed.
+Moreover, the constant $n^{-q/p}$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
+As further consequences of Euler’s identity (1.6), we next derive sharp Clarkson–McCarthy type inequalities for general unitarily invariant norms, together with a unitary-orbit refinement.
+Theorem 1.25.
 Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$,
-$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}$ $\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big),$
-$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}$ $\leq 3^{\frac{p}{2}-1}\big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\big).$
-For $0<p\leq 2$, both inequalities reverse.
-Corollary 1.13.
-Let $A,B,C\in\mathbb{M}_{n}$. Then for $1<p\leq 2$ ,
-$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
-where $q$ is the conjugate exponent of pp. For $p\geq 2$, the inequality is reversed.
-Remark 1.14.
-Set C=BC=-B in Corollary 1.12 and Corollary 1.13, we derive the classic Clarkson–McCarthy inequality in [7].
-Organization of this paper. In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorem 1.6. In Section 4, we prove Theorems 1.9, 1.10, 1.11 and derive further related consequences.
+[…]
+For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
+Theorem 1.26.
+Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$, there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
+$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-2}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
+For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
+Organization of this paper.
+Section 2 gives a negative answer to the extension problem for the matrix-weighted parallelogram law: for each $x\notin[0,1]$ we construct a counterexample, proving Theorem 1.3.
+In Section 3, we prove Theorem 1.6 and show that the range $p\leq 2$ in the Bourin–Lee unitary-orbit inequality (Theorem 1.4) is optimal by exhibiting an explicit $2\times 2$ obstruction for every $p>2$.
+Section 4 establishes Theorem 1.10 via a compact-operator construction that forces the failure of the corresponding singular-value inequality for all $p>2$, and hence confirms the optimality of (1.3).
+In Section 5, we resolve the Thompson-type triangle problem for symmetric moduli (Theorem 1.12).
+Finally, Sections 68 develop isometry-orbit refinements of an Euler operator identity (Theorems 1.15, 1.16, 1.17), derive Clarkson–McCarthy type inequalities, and draw further consequences for unitarily invariant norms and singular values.
+2. Proof of Theorem 1.3
+Proof of Theorem 1.3.
+Let $n=1$ and take scalar matrices
+$A=1,\qquad B=\frac{x-1}{x}\quad(\text{note that }x\neq 0\text{ since }x\notin[0,1]).$
+Then
+$A\nabla_{x}B=(1-x)A+xB=(1-x)+x\cdot\frac{x-1}{x}=0,$
+while
+$B\nabla_{x}A=(1-x)B+xA=(1-x)\frac{x-1}{x}+x=\frac{2x-1}{x}\neq 0,$
+because $x\notin[0,1]$ implies $x\neq\tfrac{1}{2}$. Moreover,
+$A-B=1-\frac{x-1}{x}=\frac{1}{x}\neq 0.$
+Assume, for contradiction, that Theorem 1.1 holds for this $x$. Since $x\notin[0,1]$, we have x(1x)<0x(1-x)<0. With $A\nabla_{x}B=0$, the identity in Theorem 1.1 would reduce to
+$|A\oplus B|^{2}=V\,|B\nabla_{x}A|^{2}V^{*}+x(1-x)\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\},$
+for some isometries $V,S,T\in\mathbb{M}_{2,1}$. Moving the last term to the left and using x(1x)>0-x(1-x)>0, we obtain
+$|A\oplus B|^{2}+(-x(1-x))\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\}=V\,|B\nabla_{x}A|^{2}V^{*}.$
+Here $|A\oplus B|^{2}=\operatorname{diag}(|A|^{2},|B|^{2})=\operatorname{diag}\!\bigl(1,|(x-1)/x|^{2}\bigr)$ is positive definite, and the bracketed term is positive semidefinite; hence the whole left-hand side is positive definite, thus has rank $2$.
+On the other hand, since $V\in M_{2,1}$ is an isometry, $VV^{*}$ is a rank-one projection; moreover $|B\nabla_{x}A|^{2}>0$. Therefore the right-hand side $V\,|B\nabla_{x}A|^{2}V^{*}$ has rank $1$.
+This is impossible. Hence Theorem 1.1 cannot hold for this $x$. Since $x\in\mathbb{R}\setminus[0,1]$ was arbitrary, the claim follows. ∎
 3. Proof of Theorem 1.6
@@
 Thus, to disprove (1.2) for $p>2$, it suffices to find, for a given $p$, a matrix $Z$ violating (3.1).
-$2.$ A $2\times 2$ counterexample for $p>2$
+3.2. A $2\times 2$ counterexample for $p>2$
 Fix $\theta\in(0,\pi/2)$ and set $c=\cos\theta$, $s=\sin\theta$. Define
@@
 ∎
-Claim $2.$
+Claim 3.2.
 For $Z_{\theta}$ in (3.2),
@@
 For $\operatorname{Im}\,Z_{\theta}$, the eigenvalues are $\pm s/2$, hence $\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=s$. Adding yields (3.4). ∎
-Combining (2.1), Claims 2.1 and 2.2, we see that if (1.1) were to hold for $Z_{\theta}$, then necessarily
+Combining (3.1), Claims 3.1 and 3.2, we see that if (1.2) were to hold for $Z_{\theta}$, then necessarily
 $\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}\leq 1+s.$ (3.5)
@@
 If (1.2) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (3.1).
-Claims 2.1 and 2.2 reduce (2.1) to (2.5), which fails for all sufficiently small $\theta$ by Claim 2.3.
+Claims 3.1 and 3.2 reduce (3.1) to (3.5), which fails for all sufficiently small $\theta$ by Claim 3.3.
 Hence (1.2) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
 Proof of Theorem 1.10.
-Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define
-$Ze_{1}=e_{1},\qquad Ze_{2}=e_{2},\qquad Ze_{3}=i\,e_{3},\qquad Ze_{n}=0\ \ (n\geq 4).$
-Equivalently,
-$Z=\operatorname{diag}(1,\,1,\,i,\,0,\,0,\dots).$
-This is a rank-33 operator, hence $Z\in\mathbb{K}$.
-Since $Z$ is diagonal (hence normal), we have $|Z|=|Z^{*}|$. Indeed,
-$Z^{*}=\operatorname{diag}(1,\,1,\,-i,\,0,\,0,\dots),\qquad Z^{*}Z=\operatorname{diag}(1,\,1,\,1,\,0,\,0,\dots),$
+Let $H=\ell^{2}(\mathbb{N})$ with its standard orthonormal basis $(e_{n})_{n\geq 1}$, and define a finite-rank operator $Z\in\mathbb{K}$ by
+$Ze_{1}=e_{2},\qquad Ze_{2}=e_{3},\qquad Ze_{n}=0\ \ (n\geq 3).$
+Equivalently, $Z$ is the unilateral shift truncated to the first three coordinates, that is,
+$Z=\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&1&0\end{pmatrix}\oplus 0\oplus 0\oplus\cdots.$
+We first compute $|Z|$ and $|Z^{*}|$. Since
+[…]
+we get
+[…]
+Hence, for every $p>0$,
+$\frac{|Z|^{p}+|Z^{*}|^{p}}{2}=\operatorname{diag}\!\left(\frac{1}{2},\,1,\,\frac{1}{2},\,0,\,0,\dots\right),$
 and therefore
-[…]
-Consequently, for every $p>0$,
-$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=|Z|.$
-Hence the singular values of the left-hand operator are
-$\mu_{1}^{\downarrow}=1,\ \mu_{2}^{\downarrow}=1,\ \mu_{3}^{\downarrow}=1,\ \mu_{m}^{\downarrow}=0\ (m\geq 4).$
-Next,
-$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\operatorname{diag}(1,\,1,\,0,\,0,\,0,\dots),$
-so $\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0$. Also,
-$\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}=\operatorname{diag}(0,\,0,\,1,\,0,\,0,\dots),$
-so $\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0$.
-Choose $j=k=1$, so that $1+j+k=3$ and $1+k=2$. Then for every $p>0$,
-$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}(|Z|)=1,$
+$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}=\operatorname{diag}\!\left(2^{-1/p},\,1,\,2^{-1/p},\,0,\,0,\dots\right).$
+In particular, the singular values of this operator satisfy
+$\mu_{1}^{\downarrow}=1,\qquad\mu_{2}^{\downarrow}=2^{-1/p},\qquad\mu_{3}^{\downarrow}=2^{-1/p},\qquad\mu_{m}^{\downarrow}=0\ (m\geq 4).$
+Next, compute the real and imaginary parts:
+$\operatorname{Re}Z=\frac{Z+Z^{*}}{2}=\begin{pmatrix}0&\frac{1}{2}&0\\[2.0pt] \frac{1}{2}&0&\frac{1}{2}\\[2.0pt] 0&\frac{1}{2}&0\end{pmatrix}\oplus 0\oplus 0\oplus\cdots,\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
+A direct computation shows that $\operatorname{Re}Z$ has eigenvalues 0 and $\pm 2^{-1/2}$, hence
+$\mu^{\downarrow}(\operatorname{Re}Z)=\left(2^{-1/2},\,2^{-1/2},\,0,\,0,\,0,\dots\right).$
+Moreover, $\operatorname{Im}Z$ is unitarily similar to $\operatorname{Re}Z$.
+Indeed, for the $3\times 3$ leading block set $D=\operatorname{diag}(1,-i,-1)$ and extend it by $\oplus I$ on the remaining coordinates.
+Then one checks that
+$\operatorname{Im}Z=D\,(\operatorname{Re}Z)\,D^{*}.$
+Hence $\operatorname{Re}Z$ and $\operatorname{Im}Z$ have the same singular values.
+$\mu^{\downarrow}(\operatorname{Im}Z)=\left(2^{-1/2},\,2^{-1/2},\,0,\,0,\,0,\dots\right).$
+Therefore,
+$\mu_{1}^{\downarrow}(\operatorname{Im}Z)=2^{-1/2},\qquad\mu_{3}^{\downarrow}(\operatorname{Re}Z)=0.$
+Choose $j=2$ and $k=0$, so that $1+j+k=3$, $1+j=3$, and $1+k=1$. Then for every $p>2$,
+$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=\mu_{3}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)=2^{-1/p},$
 while
-$\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{2}^{\downarrow}(\operatorname{Im}Z)=0+0=0.$
-Therefore (1.2) holds strictly for all $p>0$, and in particular for all $p>2$. ∎
+$\mu^{\downarrow}_{1+j}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z)=\mu_{3}^{\downarrow}(\operatorname{Re}Z)+\mu_{1}^{\downarrow}(\operatorname{Im}Z)=0+2^{-1/2}=2^{-1/2}.$
+Since $p>2$ implies $2^{-1/p}>2^{-1/2}$, we conclude that
+$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$
 Remark 4.1.
-The construction is intentionally minimal and answers the existence question in Question 1.5.
-Here $Z$ is finite rank and normal, so the left-hand side of (1.2) is independent of $p$ and equals a fixed singular value of $|Z|$.
-The strict inequality is obtained by selecting indices ($j=k=1$) for which the corresponding singular values of $\operatorname{Re}Z$ and $\operatorname{Im}Z$ vanish, while the chosen singular value of $|Z|$ remains nonzero.
+The construction is intentionally minimal and answers the existence question in Question 1.9.
+Here $Z$ has rank $2$, and the operator $\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}$ is diagonal and easy to read off explicitly.
+The strict inequality is obtained by selecting indices $(j,k)=(2,0)$ for which the corresponding singular value of $\operatorname{Re}Z$ vanishes at level $1+j=3$, while $\operatorname{Im}Z$ contributes only its top singular value $2^{-1/2}$, yet the third singular value on the left-hand side equals $2^{-1/p}>2^{-1/2}$ whenever $p>2$.
+5. Rectangular Thompson inequality and proof of Theorem 1.12
+In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.12 (ii).
+The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$.
+After recording this elementary fact (Lemma 5.1), we apply the classical (square) Thompson inequality to the compressed matrices $W^{*}A$ and $W^{*}B$ and then compare $|W^{*}A|,|W^{*}B|$ with $|A|,|B|$, respectively, to obtain the desired unitary-orbit domination.
+Finally, we use the rectangular Thompson inequality to settle Question 1.11, thereby proving Theorem 1.12.
+Lemma 5.1.
+Let $X\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
+$X=W|X|,\qquad|X|=(X^{*}X)^{1/2},$
+and
+$W^{*}W=\mathrm{supp}(|X|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|X|=|X|,$
+where $\mathrm{supp}(\cdot)$ means the support projection.
+Proof.
+Take a singular value decomposition $X=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
+$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(X).$
+Then
+$X^{*}X=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
+hence
+$|X|=(X^{*}X)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
+Define $W:=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\in\mathbb{M}_{m,n}.$ Clearly, $W^{*}W\leq I_{n},WW^{*}\leq I_{m}$ and
+$W|X|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=X.$
+Moreover,
+$(W^{*}W)|X|=P|X|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|X|.$
+That is, $W^{*}W$ is exactly the support projection of $|X|$. ∎
+Lemma 5.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|X|$.
+This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
+Lemma 5.2 (Rectangular Thompson inequality).
+Let $A,B\in\mathbb{M}_{m,n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
+$|A+B|\ \leq\ U|A|U^{*}+V|B|V^{*}.$ (5.1)
+Proof.
+By using Lemma 5.1, there exists a partial isometry $W$ such that A+B=W|A+B|A+B=W\,|A+B|, $W^{*}W$ is the support projection of $|A+B|$ and $W^{*}W\,|A+B|=|A+B|$.Therefore
+$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
+In particular,
+$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (5.2)
+Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices. Applying Thompson’s inequality (1.1) (the square case) to $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (5.3)
+Since the left-hand side equals $|A+B|$, it remains to compare $|W^{*}A|$ with $|A|$.
+Since $WW^{*}\leq I_{m}$, we have
+$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
+By the operator monotonicity of $t\mapsto t^{1/2}$ on $0,\infty)$, this implies $|W^{*}A|\leq|A|$. Similarly, $|W^{*}B|\leq|B|$. Conjugating preserves the order, hence
+$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
+Substituting into (5.3) and combining with (5.2) yields (5.1). ∎
+With Lemma 5.2 in hand, we now prove Theorem 1.12.
+Part (i) is verified by a concrete counterexample for the symmetric modulus, while part (ii) follows from an application of Lemma 5.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
+Proof of Theorem 1.12.
+(i) Take
+$X=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad Y=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in M_{2}(\mathbb{R}).$
+A direct computation gives
+$|X|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|Y|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|X+Y|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
+Hence
+$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
+so
+$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e.
+}\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
+If the inequality in Question 1.11 held for some unitaries $U,V$, then taking operator norms would yield
+$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty},$
+a contradiction. This proves (i).
+(ii) Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
+$\Gamma(Z):=\begin{pmatrix}Z\\ Z^{*}\end{pmatrix}$.
+Then
+$\Gamma(Z)^{*}\Gamma(Z)=Z^{*}Z+ZZ^{*},\qquad\text{hence}\qquad|\Gamma(Z)|=\bigl(Z^{*}Z+ZZ^{*}\bigr)^{1/2}=\sqrt{2}\,|Z|_{\mathrm{qsym}}.$
+Moreover, $\Gamma(X+Y)=\Gamma(X)+\Gamma(Y)$. Applying rectangular Thompson inequality (Lemma 5.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
+$|\Gamma(X+Y)|\leq U|\Gamma(X)|U^{*}+V|\Gamma(Y)|V^{*}.$
+Dividing by $\sqrt{2}$ and using $|\Gamma(Z)|=\sqrt{2}\,|Z|_{\mathrm{qsym}}$ gives
+$|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*},$
+as desired. ∎
 6. Euler-type identities and isometry orbits
-4.1. Euler-type identities
+In this section, we develop an operator-valued Euler identity and its consequences in the language of isometry orbits.
+Our goal is twofold.
+First, we give a proof of Theorem 1.15 by encoding the Euler-type relation into a unitary conjugation of suitable block matrices and then decomposing the resulting positive operator into isometric compressions of its blocks.
+Second, we collect a convenient isometry decomposition principle for positive block matrices, which will be used later to derive orbit-dominance statements from positivity and symmetry.
+We start with a Pythagoras-type theorem for partitioned matrices due to Bourin and Lee [11], which allows us to express the square modulus of a block matrix as a sum of isometric conjugations of the square moduli of its blocks.
+Lemma 6.1 ([11, Theorem 2.1]).
+Let $m,n\geq 1$ and let $T=[T_{ij}]_{i,j=1}^{m}$ be an $m\times m$ block matrix with $T_{ij}\in\mathbb{M}_{n}$. Then there exist isometries $W_{ij}\in\mathbb{M}_{mn,n}$ such that
+$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
+Remark 6.2.
+The standard $m\times m$ grid partitioning is row/column compatible in the sense of [11], so the above lemma applies directly to block matrices $T=[T_{ij}]$ written with respect to such a partition.
+We now apply Lemma 6.1 to a block matrix obtained from a Hadamard conjugation. This provides an isometric-orbit refinement of the Euler operator identity and yields Theorem 1.15.
+Proof of Theorem 1.15.
 Let
@@
 and set $\mathcal{U}:=H\otimes I_{n}\in\mathbb{M}_{4n}$. A direct block computation gives $X=\mathcal{U}\,Y$. Hence $X^{*}X=Y^{*}Y$, i.e. (1.6).
-We use the following Pythagoras theorem for partitioned matrices proved by Bourin and Lee [5].
-Lemma 4.1 ([5, Theorem 2.1]).
-Let $m,n\geq 1$ and let $T=[T_{ij}]_{i,j=1}^{m}$ be an $m\times m$ block matrix with $T_{ij}\in\mathbb{M}_{n}$. Then there exist isometries $W_{ij}\in\mathbb{M}_{mn,n}$ such that
-$|T|^{2}=\sum_{i=1}^{m}\sum_{j=1}^{m}W_{ij}\,|T_{ij}|^{2}\,W_{ij}^{*}.$
-Remark 4.2.
-The standard $m\times m$ grid partitioning is row/column compatible in the sense of [5], so the above lemma applies directly to block matrices $T=[T_{ij}]$ written with respect to such a partition.
-Next, we prove Theorem 1.9.
-Proof of Theorem 1.9.
 Let
-$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n},\qquad\mathcal{U}:=H\otimes I_{n}.$
+$\Delta_{Y}:=\begin{pmatrix}A+B&0&0&0\\ 0&B+C&0&0\\ 0&0&C+A&0\\ 0&0&0&0\end{pmatrix}\in\mathbb{M}_{4n}.$
 Set
@@
 Renaming $U_{ij}:=\mathcal{U}^{*}V_{ij}$ gives (1.15). ∎
-Naturally, we ask the following question.
-Question 4.3.
-Does there exist isometries $U_{i}\in\mathbb{M}_{3n,n}$ ($1\leq i\leq 4$) such that
-$|A+B|^{2}\oplus|B+C|^{2}\oplus|A+C|^{2}$ $=U_{1}\,|A+B+C|^{2}\,U_{1}^{*}+U_{2}\,|A|^{2}\,U_{2}^{*}$
-$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
-The answer to Question 4.3 is negative. In fact, take $A=B=C=I_{n}$. Then
-$|A+B|^{2}=|2I_{n}|^{2}=4I_{n},\qquad|A+B+C|^{2}=|3I_{n}|^{2}=9I_{n},\qquad|A|^{2}=|B|^{2}=|C|^{2}=I_{n}.$
-Hence the desired identity would become
-$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
-Since all terms on the right-hand side are positive semidefinite, we have
-$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
-But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|=9$ while $\|4I_{3n}\|=4$, a contradiction. Therefore such isometries cannot exist in general.
-Lemma 4.4 ([4, Lemma 3.4]).
-Let
-$\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$
-be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
+Next, we record a useful lemma of Bourin and Lee [9], which decomposes a $2\times 2$ positive block matrix into a sum of two unitary orbits of its diagonal compressions.
+This lemma will serve as a building block for a higher-dimensional isometry decomposition.
+Lemma 6.3 ([9, Lemma 3.4]).
+Let $\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$ be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
 $\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}=U\begin{pmatrix}X&0\\ 0&0\end{pmatrix}U^{*}+V_{0}\begin{pmatrix}0&0\\ 0&Z\end{pmatrix}V_{0}^{*}.$
-As an immediate corollary of Lemma 4.4, we obtain the following result. For completeness, we provide an alternative proof.
+As an immediate consequence of Lemma 6.3, one obtains an isometry decomposition for positive block matrices with an arbitrary number of blocks.
+For the reader’s convenience, we include a self-contained proof.
 Lemma 6.4.
@@
 Proof.
-Since $H\geq 0$, let $R:=H^{1/2}\in\mathbb{M}_{mn}$, so that $H=R^{*}R$. Partition $R$ into $m$ block columns of width $n$:
-$R=\bigl[\,R_{1}\;R_{2}\;\cdots\;R_{m}\,\bigr],\qquad R_{k}\in\mathbb{M}_{mn,n}.$
-Then
-$H=R^{*}R=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (4.2)
+Since $H\geq 0$, let $R:=H^{1/2}$, so $R=R^{*}$ and $H=R^{2}=RR^{*}=R^{*}R$. Partition $R$ into block columns $R=[R_{1}\ \cdots\ R_{m}]$. Then
+$H=RR^{*}=\sum_{k=1}^{m}R_{k}R_{k}^{*}.$ (6.2)
 Step 1: Identify the diagonal blocks. By block multiplication, the $(k,k)$ diagonal block of $H=R^{*}R$ is
@@
 $V_{k}:=U_{k}+W_{k}\in\mathbb{M}_{mn,n}.$
-Then, using $U_{k}^{*}W_{k}=0$ (their ranges are orthogonal since $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$) and $W_{k}^{*}U_{k}=0$ (initial spaces are orthogonal since $W_{k}^{*}W_{k}=I-P_{k}$), we get
-$V_{k}^{*}V_{k}=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n}.$
-Hence $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$. Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
+Then we claim the cross terms vanish.
+First, since $Q_{k}:=U_{k}U_{k}^{*}$ is the range projection of $U_{k}$, we have $Q_{k}U_{k}=U_{k}$, hence $U_{k}^{*}(I_{mn}-Q_{k})=0$.
+On the other hand, $W_{k}W_{k}^{*}\leq I_{mn}-Q_{k}$ implies $\operatorname{Ran}(W_{k})\subseteq\operatorname{Ran}(I_{mn}-Q_{k})$, equivalently $(I_{mn}-Q_{k})W_{k}=W_{k}$.
+Therefore
+$U_{k}^{*}W_{k}=U_{k}^{*}(I_{mn}-Q_{k})W_{k}=0.$
+Similarly, $U_{k}$ has initial projection $P_{k}=U_{k}^{*}U_{k}$, so $U_{k}P_{k}=U_{k}$ and hence $(I_{n}-P_{k})U_{k}^{*}=0$.
+Since $W_{k}^{*}W_{k}=I_{n}-P_{k}$, we have $\operatorname{Ran}(W_{k}^{*})\subseteq\operatorname{Ran}(I_{n}-P_{k})$, equivalently $W_{k}^{*}=W_{k}^{*}(I_{n}-P_{k})$.
+Thus
+$W_{k}^{*}U_{k}=W_{k}^{*}(I_{n}-P_{k})U_{k}=0.$
+Consequently,
+$V_{k}^{*}V_{k}=(U_{k}+W_{k})^{*}(U_{k}+W_{k})=U_{k}^{*}U_{k}+W_{k}^{*}W_{k}=P_{k}+(I_{n}-P_{k})=I_{n},$
+so $V_{k}$ is an isometry in $\mathbb{M}_{mn,n}$.
+Moreover, since $H_{kk}^{1/2}(I_{n}-P_{k})=0$, we still have
 $V_{k}H_{kk}^{1/2}=(U_{k}+W_{k})H_{kk}^{1/2}=U_{k}H_{kk}^{1/2}=R_{k}.$
@@
 which is exactly (6.1). ∎
-Now, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
+We are now ready to prove Theorem 1.16 and Theorem 1.17.
+In both cases, we conjugate an appropriate direct sum by a Fourier matrix $F_{m}\otimes I_{n}$ to obtain a positive block matrix with constant diagonal blocks, and then apply Lemma 6.4 to conclude the desired isometry-orbit domination.
+Proof of Theorem 1.16.
+First, we compute Fourier conjugations of two direct sums. Let $\omega=e^{2\pi i/3}$ and set
 $W_{3}=\frac{1}{\sqrt{3}}\begin{pmatrix}I&I&I\\ I&\omega I&\omega^{2}I\\ I&\omega^{2}I&\omega I\end{pmatrix}=F_{3}\otimes I_{n}\in\mathbb{M}_{3n}.$
@@
 In particular, every diagonal block of $W_{3}DW_{3}^{*}$ equals $(D_{1}+D_{2}+D_{3})/3$.
-Proof of Theorem 1.10.
-Set $D=\operatorname{diag}(|A+B|^{2},|B+C|^{2},|A+C|^{2})$ and let $W_{3}$ be as above.
-Then $W_{3}DW_{3}^{*}\geq 0$ and, by (4.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$.
-A direct expansion shows
+Clearly, $W_{3}DW_{3}^{*}\geq 0$ and, by (6.3), each diagonal block equals $(D_{1}+D_{2}+D_{3})/3$. A direct expansion shows
 $D_{1}+D_{2}+D_{3}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
@@
 Conjugating by $W_{3}^{*}$ and setting $U_{k}=W_{3}^{*}V_{k}$ (still isometries) gives the claim. ∎
-Next, let $\eta=e^{2\pi i/4}=i$ and set
+Proof of Theorem 1.17.
+First, let $\eta=e^{2\pi i/4}=i$ and set
 $W_{4}=\frac{1}{2}\begin{pmatrix}I&I&I&I\\ I&\eta I&\eta^{2}I&\eta^{3}I\\ I&\eta^{2}I&\eta^{4}I&\eta^{6}I\\ I&\eta^{3}I&\eta^{6}I&\eta^{9}I\end{pmatrix}=F_{4}\otimes I_{n}\in\mathbb{M}_{4n},$
@@
 In particular, every diagonal block of $W_{4}EW_{4}^{*}$ equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$.
-Proof of Theorem 1.11.
-Let $E=\operatorname{diag}(|A+B+C|^{2},|A|^{2},|B|^{2},|C|^{2})$ and let $W_{4}$ be the $4n\times 4n$ Fourier matrix above.
-Then $W_{4}EW_{4}^{*}\geq 0$ and, by (4.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$.
-As in the previous proof,
+Clearly, $W_{4}EW_{4}^{*}\geq 0$ and, by (6.4), each diagonal block equals $(E_{1}+E_{2}+E_{3}+E_{4})/4$. As in the previous proof,
 $E_{1}+E_{2}+E_{3}+E_{4}=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|A+C|^{2}.$
@@
 Conjugating by $W_{4}^{*}$ and setting $U_{k}=W_{4}^{*}V_{k}$ completes the proof. ∎
-4.2. Clarkson–McCarthy type inequalities
-Let $A,B,C\in\mathbb{M}_{n}$. Set
-$Y_{1}:=\begin{pmatrix}A+B\\ B+C\\ C+A\end{pmatrix}\in\mathbb{M}_{3n,n}.$
-Consider the $4\times 3$ partial Hadamard (isometry) matrix
-$U=\frac{1}{2}\begin{pmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{pmatrix},\qquad U^{*}U=I_{3},\qquad\mathcal{U}:=U\otimes I_{n}\in\mathbb{M}_{4n,3n}.$
-Define
-$X:=\mathcal{U}\,Y_{1}=\begin{pmatrix}A+B+C\\ B\\ A\\ -C\end{pmatrix}\in\mathbb{M}_{4n,n}.$
-In particular, since |C|=|C||-C|=|C|, we have
-$X^{*}X=|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}.$
-Starting from the Euler-type square-sum relation and its orbit refinements, we now record several consequences for unitarily invariant norms and Schatten norms by using theorems in [8].
-It suffices to notice the operator identity (1.3).
-Corollary 4.6 (Unitary-invariant norms under Euler’s square-sum constraint).
+7. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
+In this section, we exploit Euler’s operator identity (1.6) and the associated isometry-/unitary-orbit refinements obtained in Section 6 to derive sharp Clarkson–McCarthy type inequalities.
+We treat two parallel frameworks: first, inequalities for general unitarily invariant norms; second, stronger unitary-orbit dominations, from which Schatten $p$-norm inequalities follow by taking traces.
+Finally, we introduce a vector-valued complex interpolation scheme and deduce a mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimate, which will be used to prove the mixed-exponent Euler-type inequalities later in the paper.
+7.1. Unitarily invariant norm case
+We begin with a Jensen-type principle for unitarily invariant norms, which allows us to convert the square-sum identity (1.6) into $p$-power estimates via the function $t\mapsto t^{p/2}$.
+We will use the following Jensen-type inequalities for unitarily invariant norms.
+Lemma 7.1 ([14, Lemma 2.1]).
+Let $m\geq 1$. Let $A_{1},\dots,A_{m}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{m}$ be positive real numbers such that $\sum_{j=1}^{m}\alpha_{j}=1$. Then
+  1. (1)
+    for every convex function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
+[…]
+[…]
+  2. (2)
+    for every concave function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
+[…]
+[…]
+Proof of Theorem 1.25.
+Set $r:=p/2$ and $f(t):=t^{r}$ on $0,\infty)$. Note that $f(0)=0$.
+Define the positive semidefinite matrices
+$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
+and
+$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
+By Euler’s identity (1.6),
+$X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}=:S.$
+Moreover, $|A+B|^{p}=X_{1}^{r}=f(X_{1})$, etc., and $|A|^{p}=Y_{2}^{r}=f(Y_{2})$, etc.
+Step 1: case 1: $p\geq 2$. That is, $r\geq 1$ Then $f$ is convex.
+Applying Lemma 7.1 (1) the second inequality to $X_{1},X_{2},X_{3}$ yields
+[…]
+Next, applying Lemma 7.1 (1) the first inequality with $n=4$ and $\alpha_{j}=1/4$ to $Y_{1},Y_{2},Y_{3},Y_{4}$ gives
 […]
@@
 […]
-For $0<p\leq 2$, both inequalities are reversed.
-The preceding norm inequalities admit corresponding unitary-orbit refinements, which we state next.
-Corollary 4.7 (Unitary-orbit Euler refinement).
-Let $A,B,C\in\mathbb{M}_{n}$. If $p\geq 2$, then there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
-$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-1}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
-Moreover, there exist unitaries $V_{1},V_{2},V_{3},V_{4}\in\mathbb{M}_{n}$ such that
-$V_{1}|A+B+C|^{p}V_{1}^{*}+V_{2}|A|^{p}V_{2}^{*}+V_{3}|B|^{p}V_{3}^{*}+V_{4}|C|^{p}V_{4}^{*}\leq 3^{\frac{p}{2}-1}\big(|A+B|^{p}+|B+C|^{p}+|C+A|^{p}\big).$
-For $0<p\leq 2$, both inequalities reverse.
-Passing from general unitarily invariant norms to Schatten $p$-norms, we obtain Clarkson–McCarthy type estimates in Corollary 1.12.
-Using the explicit coefficient matrix $U$ above (hence $\max_{i,j}|u_{ij}|=\tfrac{1}{2}$), we obtain the mixed $\ell^{q}$$\ell^{p}$ Clarkson–McCarthy type estimate in Corollary 1.13.
-4.3. A unitary-orbit Euler modulus inequality and singular value consequences
-We use the following unitary subadditivity result for concave functions due to Aujla and Bourin [1, Theorem 2.1].
+Hence
+[…]
+Combining the above inequalities yields the desired estimate for $p\geq 2$.
+Step 2: $0<p\leq 2$. That is, $0<r\leq 1$. Then $f$ is concave.
+Applying Lemma 7.1 (2) the second inequality to $X_{1},X_{2},X_{3}$ yields
+[…]
+Next, applying Lemma 7.1 (2) the first inequality with $n=4$ and $\alpha_{j}=1/4$ to $Y_{1},Y_{2},Y_{3},Y_{4}$ gives
+[…]
+Using homogeneity again,
+[…]
+and therefore
+[…]
+[…]
+Step 3: optimality. Taking $A=B=-C\neq 0$, we have $A+B=2A$, $B+C=0$, $C+A=0$, and $A+B+C=A$. Thus
+[…]
+while
+[…]
+Hence equality holds for all $p>0$, and the constant $2^{p-2}$ is best possible. ∎
+7.2. Unitary orbit case
+We next turn to a unitary-orbit version of Jensen’s inequality for monotone convex/concave functions.
+These orbit inequalities yield matrix-order dominations, and hence deduce Schatten $p$-norm consequences.
+The following unitary-orbit inequalities on monotone concave/convex functions will be used in the sequel.
+Lemma 7.2 (1) the first inequality is obtained by a straightforward induction on $n$ from the corresponding two-variable inequality [9, Corollary 3.2].
+Lemma 7.2 (2) the first inequality follows similarly from [9, Theorem 3.1] (equivalently, [2, Theorem 2.1]).
+Lemma 7.2 (1) the second inequality follows from the isometric-column Jensen inequality [9, Corollary 2.4] by choosing $Z_{i}=\sqrt{\alpha_{i}}\,I$, so that $\sum_{i=1}^{n}Z_{i}^{*}Z_{i}=I$ and $\sum_{i=1}^{n}Z_{i}^{*}A_{i}Z_{i}=\sum_{i=1}^{n}\alpha_{i}A_{i}$.
+Similarly, Lemma 7.2 (2) the second inequality follows from the same result, using the fact that the inequality in [9, Corollary 2.4] reverses for concave functions.
+Lemma 7.2.
+Let $A_{1},\dots,A_{n}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{n}$ be positive real numbers such that $\sum_{j=1}^{n}\alpha_{j}=1$. Then
+  1. (1)
+    for every monotone convex function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\leq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
+    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\geq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
+    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\leq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*};$
+  2. (2)
+    for every monotone concave function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\geq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
+    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\leq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
+    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\geq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*}.$
+We remark that Theorems 1.19, 1.26, and 7.10 can in fact be derived directly from the results of the author in [21, Theorem 1.12 and Corollary 1.14].
+For the reader’s convenience, and to keep the present paper self-contained, we nevertheless include complete proofs.
+Proof of Theorem 1.26.
+Let $r:=p/2$ and define $f(t):=t^{r}$ on $0,\infty)$.
+Note that $f$ is monotone.
+If $p\geq 2$, then $r\geq 1$ and $f$ is convex with $f(0)=0\leq 0$.
+If $0<p\leq 2$, then $0<r\leq 1$ and $f$ is concave with $f(0)=0\geq 0$.
+Set
+$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
+and
+$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
+By Euler’s identity (1.6),
+$S:=X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}.$
+Moreover, $f(X_{1})=|A+B|^{p}$, $f(X_{2})=|B+C|^{p}$, $f(X_{3})=|C+A|^{p}$, and $f(Y_{1})=|A+B+C|^{p}$, $f(Y_{2})=|A|^{p}$, $f(Y_{3})=|B|^{p}$, $f(Y_{4})=|C|^{p}$.
+Step 1: case $p\geq 2$. Apply Lemma 7.2 (1) the first inequality to $X_{1},X_{2},X_{3}$. Then there exist unitaries $W_{1},W_{2},W_{3}\in\mathbb{M}_{n}$ such that
+$\sum_{j=1}^{3}W_{j}f(X_{j})W_{j}^{*}\ \leq\ f(S).$ (7.1)
+Next apply Lemma 7.2 (1) the second inequality to $Y_{1},Y_{2},Y_{3},Y_{4}$ with weights $\alpha_{j}=1/4$. Then there exists a unitary $U\in\mathbb{M}_{n}$ such that
+$f\!\Bigl(\frac{1}{4}S\Bigr)\ \leq\ U\Bigl(\frac{1}{4}\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
+Using $f(tX)=t^{r}f(X)$ for $t>0$, we obtain
+$f(S)=4^{r}f\!\Bigl(\frac{1}{4}S\Bigr)\ \leq\ 4^{r-1}\,U\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
+Conjugating by $U^{*}$ yields
+$U^{*}f(S)U\ \leq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k}).$ (7.2)
+Now conjugate (7.1) by $U^{*}$ and set $U_{j}:=U^{*}W_{j}$ to get
+$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \leq\ U^{*}f(S)U.$
+Combining with (7.2) gives
+$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \leq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k})=2^{p-2}\bigl(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\bigr),$
+i.e.
+$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-2}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
+Step 2: case $0<p\leq 2$. Now apply Lemma 7.2 (2) the first inequality to $X_{1},X_{2},X_{3}$ to get
+[…]
+for some unitaries $W_{j}$. Also apply Lemma 7.2 (2) the second inequality to $Y_{1},\dots,Y_{4}$ with $\alpha_{k}=1/4$ to get a unitary $U$ such that
+$f(S)=4^{r}f\!\Bigl(\frac{1}{4}S\Bigr)\ \geq\ 4^{r-1}\,U\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
+Conjugating and combining exactly as above (setting $U_{j}:=U^{*}W_{j}$) yields the reversed inequality
+$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \geq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k})=2^{p-2}\bigl(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\bigr).$
+Step 3: optimality. Take $A=B=-C\neq 0$. Then $A+B=2A$, $B+C=0$, $C+A=0$, and $A+B+C=A$. Hence the inequality becomes an equality for all $p>0$, and the constant $2^{p-2}=4^{r-1}$ is best possible. ∎
+Taking traces in the unitary-orbit inequality (or, equivalently, taking the trace norm in the unitarily invariant norm statement) immediately yields the Schatten $p$-norm formulation.
+Proof of Theorem 1.19.
+Taking trace norms in Theorem 1.25, or simply taking traces in Theorem 1.26, and then passing to general Schatten class operators by approximating with finite-rank operators, yields Theorem 1.19. ∎
+7.3. Complex interpolation theory
+We now introduce the tools from complex interpolation needed in this paper; for the basic theory of interpolation, we refer the interested reader to [3, Chapters 1–5].
+Definition 7.3.
+Let $A_{0}$ and $A_{1}$ be two topological vector spaces.
+We say that $A_{0}$ and $A_{1}$ are compatible if there exists a Hausdorff topological vector space $\mathcal{Z}$ such that $A_{0}$ and $A_{1}$ are subspaces of $\mathcal{Z}$.
+In this case, we can form their sum $A_{0}+A_{1}$ and their intersection $A_{0}\cap A_{1}$.
+Definition 7.4.
+Let $X$ and $Y$ be normed spaces, and let $T:X\to Y$ be a bounded linear operator. The operator norm of $T$ is defined by
+$\|T\|_{X\to Y}:=\sup_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\frac{\|Tx\|_{Y}}{\|x\|_{X}}.$
+Definition 7.5.
+Let $(X_{0},X_{1})$ be a compatible Banach couple and $0<\theta<1$. The space $(X_{0},X_{1})_{[\theta]}$ is defined as
+$(X_{0},X_{1})_{[\theta]}:=\{f(\theta):f\in\mathcal{F}(X_{0},X_{1})\},$
+where $\mathcal{F}(X_{0},X_{1})$ denotes the space of all bounded continuous functions $f:\{z\in\mathbb{C}:0\leq\Re z\leq 1\}\to X_{0}+X_{1}$ which are analytic on $\{z\in\mathbb{C}:0<\Re z<1\}$, satisfy $f(it)\in X_{0}$ for all $t\in\mathbb{R}$ and $f(1+it)\in X_{1}$ for all $t\in\mathbb{R}$, and are bounded on the boundary lines in the respective norms of $X_{0}$ and $X_{1}$.
+The norm on $(X_{0},X_{1})_{[\theta]}$ is given by
+$\|x\|_{(X_{0},X_{1})_{[\theta]}}:=\inf\left\{\|f\|_{\mathcal{F}(X_{0},X_{1})}:f(\theta)=x\right\}.$
+We recall the abstract complex interpolation theorem for Banach couples; see [3, Theorem 4.1.2].
+Theorem 7.6 ([3]).
+Let $(X_{0},X_{1})$ and $(Y_{0},Y_{1})$ be two compatible couples of Banach spaces, and let $0<\theta<1$. Suppose that
+$T:X_{0}+X_{1}\longrightarrow Y_{0}+Y_{1}$
+is a linear operator such that $T:X_{j}\to Y_{j}$ is bounded for $j=0,1$. Then $T$ is bounded from $(X_{0},X_{1})_{[\theta]}$ to $(Y_{0},Y_{1})_{[\theta]}$, and
+$\|T\|_{(X_{0},X_{1})_{[\theta]}\to(Y_{0},Y_{1})_{[\theta]}}\leq\|T\|_{X_{0}\to Y_{0}}^{\,1-\theta}\,\|T\|_{X_{1}\to Y_{1}}^{\,\theta}.$
+We will also use the classical Riesz–Thorin interpolation theorem.
+For $L^{p}$ spaces, see [3, Theorem 5.1.1].
+For certain noncommutative $L^{p}$ spaces, such as the Schatten $p$-classes $S_{p}$, a Riesz–Thorin interpolation theorem is also known to hold; see [18, Eq.
+(2.1)], since matrix algebras are semifinite von Neumann algebras.
+See also [13, Theorem 13.1], [19, Theorems 2.9–2.10], and [19, Remark 1, p.
+23].
+Theorem 7.7 ([18]).
+Assume that $1\leq p_{0},p_{1}\leq\infty$ and $0<\theta<1$. Then
+$(S_{p_{0}},S_{p_{1}})_{[\theta]}=S_{p}\quad\text{(with equal norms)},$
+if
+$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}.$
+We shall also use the complex interpolation formula for $\ell_{p}$-sums; see [3, Theorem 5.6.3, p.
+123].
+[^1] Following [3, p.
+121, Section 5.6], let $A$ be a Banach space, let $s\in\mathbb{R}$ and $q>0$.
+Denote by $\dot{\ell}_{q}^{\,s}(A)$ the space of all bi-infinite sequences $a=(a_{\nu})_{\nu\in\mathbb{Z}}$ with $a_{\nu}\in A$ such that $\|a\|_{\dot{\ell}_{q}^{\,s}(A)}:=\Bigg(\sum_{\nu\in\mathbb{Z}}\bigl(2^{\nu s}\|a_{\nu}\|_{A}\bigr)^{q}\Bigg)^{1/q}<\infty.$ In particular, taking $s=0$ and fixing $m\in\mathbb{N}$, we may identify an $m$-tuple $(a_{1},\dots,a_{m})$ with an element of this sequence space by extending it to a bi-infinite sequence $a=(a_{\nu})_{\nu\in\mathbb{Z}}$ via $a_{\nu}=0$ for all $\nu\notin\{1,\dots,m\}$; this agrees with our definition in Theorem 7.8.
+Theorem 7.8 ([3]).
+Let $(X_{0},X_{1})$ be a compatible couple of Banach spaces.
+Assume that $1\leq p_{0},p_{1}\leq\infty$ and $0<\theta<1$.
+For each $m\in\mathbb{N}$ and $k=0,1$, define $\ell_{p_{k}}^{m}(X_{k})$ as the space of $m$-tuples $a=(a_{1},\dots,a_{m})$ with $a_{j}\in X_{k}$, equipped with the norm
+$\|a\|_{\ell_{p_{k}}^{m}(X_{k})}=\begin{cases}\big(\sum_{j=1}^{m}\|a_{j}\|_{X_{k}}^{p_{k}}\big)^{1/p_{k}},&p_{k}<\infty,\\[5.69054pt] \max_{1\leq j\leq m}\|a_{j}\|_{X_{k}},&p_{k}=\infty.\end{cases}$
+Then
+[…]
+if
+$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}.$
+Combining Theorem 7.7 and Theorem 7.8 gives the following.
+Theorem 7.9.
+Assume that $1\leq p_{0},p_{1},q_{0},q_{1}\leq\infty$ and $0<\theta<1$. Then for each $m\in\mathbb{N}$,
+[…]
+if
+$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}\quad\text{ and }\quad\frac{1}{q}=\frac{1-\theta}{q_{0}}+\frac{\theta}{q_{1}}.$
+With the interpolation tools in hand, we proceed to prove Theorem 1.23.
+We first obtain (1.8) for $1<p\leq 2$ via complex interpolation between the endpoints $(p,q)=(1,\infty)$ and $(2,2)$, and then deduce (1.9) from (1.8) applied to $U^{*}$.
+Proof of Theorem 1.23.
+We first prove (1.8) for $1<p\leq 2$ by complex interpolation between the endpoints $(p,q)=(1,\infty)$ and $(p,q)=(2,2)$, and then deduce (1.9) from (1.8) applied to $U^{*}$.
+Step 1: endpoint 1: $\ell_{1}^{t}(S_{1})\to\ell_{\infty}^{s}(S_{1})$. For each $i$,
+$\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{1}\leq\sum_{j=1}^{t}|u_{ij}|\,\|Z_{j}\|_{1}\leq\mu\sum_{j=1}^{t}\|Z_{j}\|_{1}.$
+Taking the maximum over $i$ yields
+$\|T_{U}(Z)\|_{\ell_{\infty}^{s}(S_{1})}\leq\mu\,\|Z\|_{\ell_{1}^{t}(S_{1})},$
+hence
+$\|T_{U}\|_{\ell_{1}^{t}(S_{1})\to\ell_{\infty}^{s}(S_{1})}\leq\mu.$ (7.3)
+Step 2: endpoint 2: $\ell_{2}^{t}(S_{2})\to\ell_{2}^{s}(S_{2})$.
+Equip $S_{2}$ with the Hilbert–Schmidt inner product $\langle X,Y\rangle=\mathrm{Tr}(Y^{*}X)$, so that $\|X\|_{2}^{2}=\langle X,X\rangle$.
+A direct computation gives
+$\|T_{U}(Z)\|_{\ell_{2}^{s}(S_{2})}^{2}=\sum_{j,k=1}^{t}(U^{*}U)_{jk}\,\langle Z_{k},Z_{j}\rangle=\operatorname{Tr}(U^{*}U\,G),\qquad G=[\langle Z_{k},Z_{j}\rangle]_{j,k=1}^{t}\geq 0.$
+If $\|U\|_{\infty}\leq 1$, then $U^{*}U\leq I_{t}$, and therefore
+$\|T_{U}(Z)\|_{\ell_{2}^{s}(S_{2})}^{2}\leq\sum_{j=1}^{t}\|Z_{j}\|_{2}^{2}=\|Z\|_{\ell_{2}^{t}(S_{2})}^{2},$
+so that
+$\|T_{U}\|_{\ell_{2}^{t}(S_{2})\to\ell_{2}^{s}(S_{2})}\leq 1.$ (7.4)
+Step 3: interpolation. Fix $0<\theta<1$. Applying Theorem 7.6 to the couples
+$X_{0}=\ell_{1}^{t}(S_{1}),\quad X_{1}=\ell_{2}^{t}(S_{2}),\qquad Y_{0}=\ell_{\infty}^{s}(S_{1}),\quad Y_{1}=\ell_{2}^{s}(S_{2}),$
+and using (7.3)–(7.4), we obtain
+$\|T_{U}\|_{(X_{0},X_{1})_{[\theta]}\to(Y_{0},Y_{1})_{[\theta]}}\leq\mu^{\,1-\theta}.$ (7.5)
+By Theorem 7.9,
+$(X_{0},X_{1})_{[\theta]}=\ell_{p}^{t}(S_{p}),\qquad(Y_{0},Y_{1})_{[\theta]}=\ell_{q}^{s}(S_{p}),$
+where
+$\frac{1}{p}=1-\frac{\theta}{2},\qquad\frac{1}{q}=\frac{\theta}{2},$
+so that $q$ is the conjugate exponent of $p$ and $1<p\leq 2$. Thus (7.5) yields
+$\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\leq\mu^{\,1-\theta}\,\|Z\|_{\ell_{p}^{t}(S_{p})}=\mu^{\,\frac{2}{p}-1}\,\|Z\|_{\ell_{p}^{t}(S_{p})},$
+which is exactly (1.8).
+Step 4: the case $p\geq 2$ (reverse inequality in (1.8)).
+Assume in addition that $U^{*}U=I_{t}$ (as required in the statement for the reverse direction).
+Let $q$ be the conjugate exponent of $p$, so that $1<q\leq 2$.
+Let $Z=(Z_{1},\dots,Z_{t})\in\ell_{p}^{t}(S_{p})$. By duality,
+$\|Z\|_{\ell_{p}^{t}(S_{p})}=\sup\Big\{\,|\langle Z,W\rangle|:\ W\in\ell_{q}^{t}(S_{q}),\ \|W\|_{\ell_{q}^{t}(S_{q})}\leq 1\,\Big\},\qquad\langle Z,W\rangle:=\sum_{j=1}^{t}\operatorname{Tr}(W_{j}^{*}Z_{j}).$
+For such a $W$, since $U^{*}U=I_{t}$ we have $T_{U^{*}}\circ T_{U}=\mathrm{Id}$, hence $Z=T_{U^{*}}(T_{U}(Z))$ and therefore
+$\langle Z,W\rangle=\langle T_{U^{*}}(T_{U}(Z)),\,W\rangle=\langle T_{U}(Z),\,T_{U}(W)\rangle.$
+By Hölder’s inequality for Schatten classes and for $\ell_{p}$-sums,
+$|\langle T_{U}(Z),T_{U}(W)\rangle|\leq\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\ \|T_{U}(W)\|_{\ell_{p}^{s}(S_{q})}.$
+Since $1<q\leq 2$ and $\|U\|_{\infty}\leq 1$, we may apply the already proved estimate (1.8) (with $q$ in place of $p$ and $p$ in place of $q$) to obtain
+$\|T_{U}(W)\|_{\ell_{p}^{s}(S_{q})}\leq\mu^{\frac{2}{q}-1}\,\|W\|_{\ell_{q}^{t}(S_{q})}=\mu^{\,1-\frac{2}{p}}\,\|W\|_{\ell_{q}^{t}(S_{q})}.$
+Hence, for $\|W\|_{\ell_{q}^{t}(S_{q})}\leq 1$,
+$|\langle Z,W\rangle|\leq\mu^{\,1-\frac{2}{p}}\,\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}.$
+Taking the supremum over all such $W$ yields
+$\|Z\|_{\ell_{p}^{t}(S_{p})}\leq\mu^{\,1-\frac{2}{p}}\,\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})},$
+equivalently,
+$\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\geq\mu^{\,\frac{2}{p}-1}\,\|Z\|_{\ell_{p}^{t}(S_{p})}.$
+This is exactly the reversed inequality in (1.8) for $p\geq 2$.
+Step 5: the reverse-direction estimate. Assume now that $U^{*}U=I_{t}$. Consider the adjoint matrix $U^{*}\in\mathbb{M}_{t,s}$ and the associated map
+$T_{U^{*}}:\ \ell_{p}^{s}(S_{p})\longrightarrow\ell_{p}^{t}(S_{p}),\qquad T_{U^{*}}(W_{1},\dots,W_{s}):=\Big(\sum_{i=1}^{s}\overline{u_{i1}}\,W_{i},\ \dots,\ \sum_{i=1}^{s}\overline{u_{it}}\,W_{i}\Big).$
+Then $T_{U^{*}}\circ T_{U}=\mathrm{Id}$ on $\ell_{p}^{t}(S_{p})$. Applying (1.8) to $U^{*}$ (note that $\max_{i,j}|(U^{*})_{ij}|=\mu$) gives, for $1<p\leq 2$,
+$\|T_{U^{*}}(W)\|_{\ell_{q}^{t}(S_{p})}\leq\mu^{\,\frac{2}{p}-1}\,\|W\|_{\ell_{p}^{s}(S_{p})}\qquad(W\in\ell_{p}^{s}(S_{p})).$
+Taking $W=T_{U}(Z)$ and using $T_{U^{*}}(T_{U}(Z))=Z$ yields (1.9). The case $p\geq 2$ again follows by duality. ∎
+Proof of Theorem 1.21.
+Fix $1<p\leq 2$ and let $q$ be the conjugate exponent of $p$, i.e. $1/p+1/q=1$. Set
+$X_{1}:=A+B,\quad X_{2}:=B+C,\quad X_{3}:=C+A,\qquad Y_{1}:=A+B+C,\ Y_{2}:=A,\ Y_{3}:=B,\ Y_{4}:=C.$
+Step 1: A suborthogonal coefficient matrix with $U^{*}$ isometric. Consider the real $3\times 4$ matrix
+$U:=\frac{1}{2}\begin{pmatrix}1&1&1&-1\\ 1&-1&1&1\\ 1&1&-1&1\end{pmatrix}$.
+Then
+$\begin{pmatrix}X_{1}\\ X_{2}\\ X_{3}\end{pmatrix}=U\begin{pmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\\ Y_{4}\end{pmatrix},\qquad\text{i.e.}\qquad X_{i}=\sum_{j=1}^{4}u_{ij}Y_{j}\ (i=1,2,3).$
+Moreover, a direct computation gives $UU^{*}=I_{3}$; hence $U^{*}$ is an isometry. Also $\max_{i,j}|u_{ij}|=1/2=:\mu$.
+Step 2: The relations $X=UY$ and $Y=U^{*}X$. By construction,
+$\begin{pmatrix}X_{1}\\ X_{2}\\ X_{3}\end{pmatrix}=U\begin{pmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\\ Y_{4}\end{pmatrix}$.
+Moreover, a direct computation gives $Y=U^{*}X$ and explicitly
+$Y_{1}=\frac{X_{1}+X_{2}+X_{3}}{2},\quad Y_{2}=\frac{X_{1}-X_{2}+X_{3}}{2},\quad Y_{3}=\frac{X_{1}+X_{2}-X_{3}}{2},\quad Y_{4}=\frac{-X_{1}+X_{2}+X_{3}}{2},$
+which are easily verified from $X_{1}=A+B$, $X_{2}=B+C$, $X_{3}=C+A$.
+Step 3: Apply Theorem 1.23. Since $\|U\|_{\infty}\leq 1$, Theorem 1.23 (i) gives for $1<p\leq 2$,
+$\Big(\|X_{1}\|_{p}^{q}+\|X_{2}\|_{p}^{q}+\|X_{3}\|_{p}^{q}\Big)^{1/q}\leq\mu^{\frac{2}{p}-1}\Big(\|Y_{1}\|_{p}^{p}+\|Y_{2}\|_{p}^{p}+\|Y_{3}\|_{p}^{p}+\|Y_{4}\|_{p}^{p}\Big)^{1/p}.$
+Raising both sides to the power $q$ and using $\mu=1/2$ yields
+$\|X_{1}\|_{p}^{q}+\|X_{2}\|_{p}^{q}+\|X_{3}\|_{p}^{q}\leq 2^{\,1-\frac{q}{p}}\Big(\|Y_{1}\|_{p}^{p}+\|Y_{2}\|_{p}^{p}+\|Y_{3}\|_{p}^{p}+\|Y_{4}\|_{p}^{p}\Big)^{q/p},$
+which is the second inequality.
+Applying the same estimate to $U^{*}$ and using $Y=U^{*}X$ gives
+$\Big(\|Y_{1}\|_{p}^{q}+\|Y_{2}\|_{p}^{q}+\|Y_{3}\|_{p}^{q}+\|Y_{4}\|_{p}^{q}\Big)^{1/q}\leq\mu^{\frac{2}{p}-1}\Big(\|X_{1}\|_{p}^{p}+\|X_{2}\|_{p}^{p}+\|X_{3}\|_{p}^{p}\Big)^{1/p},$
+and raising to $q$ again yields the first inequality with the same constant $2^{\,1-\frac{q}{p}}$.
+Step 4: optimality. Taking $A=B=-C\neq 0$ gives $A+B=2A$ and $B+C=C+A=0$, while $A+B+C=A$. Thus equality holds for all $p>1$, and the constant $2^{1-\frac{q}{p}}$ is optimal.
+7.4. Derivation of Audenaert–Kittaneh’s inequality and its sharp complement
+In this subsection, we show that the Audenaert–Kittaneh inequality (1.10) and the complementary estimate (1.11) both follow from Theorem 1.23 by a suitable choice of the coefficient matrix $U$.
+Let $n\geq 2$ and set
+$s:=1+\binom{n}{2}$.
+Fix an enumeration of the pairs $(i,j)$ with $1\leq i<j\leq n$:
+$(i_{1},j_{1}),\ (i_{2},j_{2}),\ \ldots,\ (i_{\binom{n}{2}},j_{\binom{n}{2}})$.
+Define
+$U=\frac{1}{\sqrt{n}}\begin{pmatrix}\mathbf{1}^{\top}\\ (e_{i_{1}}^{\top}-e_{j_{1}}^{\top})\\ (e_{i_{2}}^{\top}-e_{j_{2}}^{\top})\\ \vdots\\ (e_{i_{\binom{n}{2}}}^{\top}-e_{j_{\binom{n}{2}}}^{\top})\end{pmatrix}\in\mathbb{M}_{s,n},$
+where $\mathbf{1}=(1,\dots,1)^{\top}\in\mathbb{C}^{n}$ and $e_{k}$ denotes the $k$th standard basis vector in $\mathbb{C}^{n}$. Then
+$\mu:=\max_{\alpha,j}|u_{\alpha j}|=\frac{1}{\sqrt{n}}.$
+We claim $U$ is an isometry, i.e. $U^{*}U=I_{n}$ (hence $\|U\|_{\infty}=1$ and $U$ is a contraction). Indeed,
+$U^{*}U=\frac{1}{n}\left(\mathbf{1}\mathbf{1}^{*}+\sum_{1\leq i<j\leq n}(e_{i}-e_{j})(e_{i}-e_{j})^{*}\right),$
+and one checks the identity
+$\sum_{1\leq i<j\leq n}(e_{i}-e_{j})(e_{i}-e_{j})^{*}=nI_{n}-\mathbf{1}\mathbf{1}^{*}.$
+Therefore $U^{*}U=\frac{1}{n}(\mathbf{1}\mathbf{1}^{*}+nI_{n}-\mathbf{1}\mathbf{1}^{*})=I_{n}$.
+Given $A_{1},\dots,A_{n}\in S_{p}$, consider the $n$-tuple $Z:=(A_{1},\dots,A_{n})\in\ell_{p}^{n}(S_{p})$. By construction,
+$T_{U}(Z)=\Big(\tfrac{1}{\sqrt{n}}\sum_{k=1}^{n}A_{k},\ \tfrac{1}{\sqrt{n}}(A_{i_{1}}-A_{j_{1}}),\ \dots,\ \tfrac{1}{\sqrt{n}}(A_{i_{\binom{n}{2}}}-A_{j_{\binom{n}{2}}})\Big)\in\ell_{q}^{s}(S_{p}).$
+Consequently, for $r\in\{p,q\}$ we have
+$\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{r}=\frac{1}{n^{r/2}}\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{r}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{r}\right).$ (7.6)
+In particular, (7.6) with $r=q$ gives the left side of (1.8), and with $r=p$ gives the right side of (1.9).
+Step 1: Audenaert–Kittaneh inequality from Theorem 1.23(i). Let $1<p\leq 2$ and let $q$ be the conjugate exponent of $p$. Since $U$ is a contraction, we may apply (1.8) to $Z=(A_{1},\dots,A_{n})$:
+$\Bigg(\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\leq\mu^{\frac{2}{p}-1}\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{1/p}.$
+Using (7.6) with $r=q$ and $\mu=1/\sqrt{n}$, we obtain
+$\frac{1}{\sqrt{n}}\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\right)^{1/q}\leq\Big(\frac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{1/p}.$
+Multiplying by $\sqrt{n}$ and noting that
+$\sqrt{n}\Big(\tfrac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}=n^{1-\frac{1}{p}},\qquad\text{and}\qquad q\Bigl(1-\frac{1}{p}\Bigr)=1,$
+we get
+$\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{q/p},$
+which is exactly (1.10). For $p\geq 2$, the inequality is reversed by Theorem 1.23(i) since $U^{*}U=I_{n}$.
+Step 2: the complementary estimate from Theorem 1.23(ii). Since $U^{*}U=I_{n}$, we may apply (1.9) to the same $U$ and $Z$. For $1<p\leq 2$,
+$\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{q}\Bigg)^{1/q}\leq\mu^{\frac{2}{p}-1}\Bigg(\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{p}\Bigg)^{1/p}.$
+Using (7.6) with $r=p$ and $\mu=1/\sqrt{n}$ yields
+$\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{q}\Bigg)^{1/q}\leq\Big(\frac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}\cdot\Big(\frac{1}{\sqrt{n}}\Big)\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{1/p}.$
+Since $\big(\tfrac{1}{\sqrt{n}}\big)^{\frac{2}{p}-1}\cdot\tfrac{1}{\sqrt{n}}=n^{-1/p}$, raising both sides to the power $q$ gives (1.11). For $p\geq 2$, the inequality is reversed by Theorem 1.23(ii).
+Step 3: optimality of the constant in (1.11). Let $A_{1}=\cdots=A_{n}=A\neq 0$. Then $\sum_{k=1}^{n}A_{k}=nA$ and $A_{i}-A_{j}=0$, hence
+[…]
+Thus equality holds for all $1<p<\infty$, and the constant $n^{-q/p}$ is best possible.
+7.5. A weaker bound on Conjecture 1.20
+We record a non-sharp estimate with the explicit constant $3^{\frac{p}{2}-1}$, which follows directly from the Euler-type identities; however, this coefficient is not optimal.
+Theorem 7.10.
+Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
+$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\ \leq\ 3^{\frac{p}{2}-1}\Bigl(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Bigr).$
+For $0<p\leq 2$, the inequality is reversed.
+Proof.
+Set $r:=p/2$ and $f(t):=t^{r}$ on $0,\infty)$. Let
+$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
+and
+$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
+By Euler’s identity (1.6),
+$S:=X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}.$
+Assume first that $p\geq 2$, i.e.
+$r\geq 1$, so that $f$ is convex and $f(0)=0$.
+Using the trace form of the Jensen/Rotfel’d type inequality for convex $f$ (equivalently, Lemma 7.1 (1) with the trace norm), we have
+$\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)\ \leq\ \operatorname{Tr}\bigl(f(S)\bigr).$ (7.7)
+Next apply Jensen’s inequality with weights $\alpha_{j}=1/3$ to $X_{1},X_{2},X_{3}$:
+$\operatorname{Tr}\Bigl(f\!\bigl(\tfrac{1}{3}S\bigr)\Bigr)\leq\operatorname{Tr}\Bigl(\tfrac{1}{3}\sum_{j=1}^{3}f(X_{j})\Bigr).$
+Since $f(tX)=t^{r}f(X)$ for $t>0$, this becomes
+$3^{-r}\operatorname{Tr}\bigl(f(S)\bigr)\ \leq\ \tfrac{1}{3}\,\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr),$
+or equivalently
+$\operatorname{Tr}\bigl(f(S)\bigr)\ \leq\ 3^{r-1}\,\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr).$ (7.8)
+Combining (7.7)–(7.8) gives
+$\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}|Y_{k}|^{r}\Bigr)=\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)\leq 3^{r-1}\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr)=3^{r-1}\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}|X_{j}|^{r}\Bigr).$
+Finally, passing to general Schatten class operators by approximating with finite-rank operators, we obtain exactly the desired inequality with constant $3^{r-1}=3^{\frac{p}{2}-1}$.
+For $0<p\leq 2$ (i.e. $0<r\leq 1$), the function $f$ is concave and all inequalities above reverse, yielding the reversed statement. ∎
+Remark 7.11.
+The constant $3^{\frac{p}{2}-1}$ in Theorem 7.10 is generally not optimal; numerical evidence suggests that the sharp constant should be the one proposed in Conjecture 1.20.
+8. Unitary-orbit refinements of Euler’s identity and spectral consequences
+8.1. A unitary-orbit bound for the Euler modulus
+To convert the square-sum identity (1.6) into a linear (modulus) estimate in unitary orbits, we use the following unitary subadditivity result for concave functions due to Aujla and Bourin [2, Theorem 2.1], which is also included in Lemma 7.2 (2) the first inequality.
 Lemma 8.1 (Aujla–Bourin).
@@
 $f(H+K)\leq Uf(H)U^{*}+Vf(K)V^{*}.$
+We now apply Lemma 8.1 with the concave function $f(t)=t^{1/2}$ to the Euler square-sum relation (1.6). This yields a three-term unitary-orbit domination of the Euler modulus by the pairwise moduli.
 Theorem 8.2.
@@
 Substitute and absorb conjugation by $V_{1}$ into $U_{2},V_{2}$ to obtain (8.1). ∎
+As a first immediate consequence, taking any unitarily invariant norm in (8.1) and using the triangle inequality yields a norm inequality.
 Corollary 8.3.
+Let $A,B,C\in\mathbb{M}_{n}$. Then
 […]
-[…]
-Proof.
-[…]
+We next record a singular-value refinement, obtained by combining Theorem 8.2 with the Weyl inequality for sums of positive semidefinite matrices.
 Corollary 8.4.
@@
 which implies the desired estimate. ∎
-$4.$ Symmetric norms, antinorms, and Ky Fan consequences
+8.2. Symmetric norms, antinorms, and Ky Fan consequences
+Finally, since Euler’s identity (1.6) is an exact equality of positive operators, it also yields further consequences for symmetric norms and antinorms (in the sense of [8]) on the positive semidefinite cone of $\mathbb{M}_{n}$, and in particular for Ky Fan sums of singular values.
 Corollary 8.5.
@@
   2. (b)
-    For every symmetric antinorm $\|\cdot\|_{!}$ on $\mathbb{M}_{n}^{+}$ (in the sense of [3]),
+    For every symmetric antinorm $\|\cdot\|_{!}$,
     $\Big\||A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big\|_{!}\geq\big\||A+B|^{2}\big\|_{!}+\big\||B+C|^{2}\big\|_{!}+\big\||C+A|^{2}\big\|_{!}.$
-Proof.
-Both inequalities follow from (1.3) and the (anti)triangle inequality on $\mathbb{M}_{n}^{+}$: a symmetric norm is subadditive, while a symmetric antinorm is superadditive. ∎
+Applying Corollary 8.5 to Ky Fan $m$-norms and Ky Fan $m$-antinorms gives the following eigenvalue-sum inequalities.
 Corollary 8.6.
-Let $A,B,C\in\mathbb{M}_{n}$ and $m=1,\dots,n$. Denote by $\mu_{j}^{\downarrow}(\cdot)$ (resp. $\mu_{j}^{\uparrow}(\cdot)$) the singular values in nonincreasing (resp. nondecreasing) order. Then
+Let $A,B,C\in\mathbb{M}_{n}$. Then for $m=1,\dots,n$,
 $\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\leq\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|C+A|^{2}\big),$
 $\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\geq\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|C+A|^{2}\big).$
-Proof.
-Apply Corollary 4.12 to Ky Fan $m$-norms and Ky Fan $m$-antinorms on $\mathbb{M}_{n}^{+}$. ∎
 Acknowledgments
-Teng Zhang is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No.
+The author is very grateful to Professor J.-C.
+Bourin for reading this manuscript and for his valuable comments.
+This work is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No.
 xzy022024045).

### v3

@@
-On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity
+An operator triangle inequality for the quadratic symmetric modulus
 Abstract.
-We answer five questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [12], and thereby obtain a sequence of sharp results for matrices and compact operators.
-We first show that the isometry-orbit identity behind the matrix weighted parallelogram law cannot be extended beyond the parameter range $0\leq x\leq 1$, and that a counterexample already exists in dimension one.
-Next, we prove that the exponent $2$ in the Bourin–Lee unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by constructing an explicit $2\times 2$ counterexample for all $p>2$.
-We then construct a compact operator $Z$ for which the associated singular-value inequality fails for every $p>2$, in fact for a fixed choice of indices.
-We also settle a Thompson-type triangle problem for symmetric moduli: the inequality fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
-Finally, we develop isometry-orbit refinements of an Euler operator identity and derive sharp Clarkson–McCarthy type inequalities for Schatten $p$-norms, together with further consequences for unitarily invariant norms and singular values.
+50 years after Thompson’s famous triangle inequality for the operator right modulus, we establish a triangle inequality for the quadratic symmetric modulus.
+We also discuss the corresponding equality cases as well as the infinite-dimensional setting.
+In addition, we obtain Clarkson–McCarthy type inequalities for the quadratic symmetric modulus.
+Moreover, we answer several questions raised by Bourin and Lee in [Bull.
+Lond.
+Math.
+Soc.
+44 (2012), no.
+6, 1085–1102] and [Internat.
+J.
+Math.
+31 (2026), no.
+6, 2650018].
 Key words and phrases:
-Symmetric modulus; unitary orbits; Euler identity; Thompson inequality; Clarkson-McCarthy type inequalities
+Quadratic symmetric modulus, triangle inequality, unitary orbits, equality cases, Clarkson–McCarthy type inequalities, Euler’s identity
 2020 Mathematics Subject Classification:
-15A60, 47A30, 47A57, 15A42, 47B10
+Primary 47A63; Secondary 47A30, 47A05
 1. Introduction
-Throughout this paper, $\mathcal{H}$ denotes a complex Hilbert space, and $\mathbb{K}$ the ideal of compact operators on $\mathcal{H}$.
-For $Z\in\mathbb{K}$, we write $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ for the singular values of $Z$, namely the eigenvalues of $|Z|=(Z^{*}Z)^{1/2}$ (counted with multiplicity) arranged in nonincreasing order.
-Let $\mathbb{M}_{m,n}$ be the space of complex $m\times n$ matrices, and write $\mathbb{M}_{n}:=\mathbb{M}_{n,n}$.
-Let $X\in\mathbb{M}_{n}$.
-We adopt the convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ can be regarded as an infinite nonincreasing sequence, consistent with the compact-operator notation.
-We denote by $X^{\mathsf{T}}$ its transpose, and by $X^{*}$ its conjugate transpose (adjoint).
-For scalars $d_{1},\dots,d_{r}\in\mathbb{C}$, we denote by $\operatorname{diag}(d_{1},\dots,d_{r})$ the $r\times r$ diagonal matrix whose diagonal entries are $d_{1},\dots,d_{r}$ (and all off-diagonal entries are 0).
-More generally, if $X_{1}\in\mathbb{M}_{n_{1}},\dots,X_{k}\in\mathbb{M}_{n_{k}}$, then $\operatorname{diag}(X_{1},\dots,X_{k}):=X_{1}\oplus\cdots\oplus X_{k}$ denotes the block-diagonal matrix with diagonal blocks $X_{1},\dots,X_{k}$.
-1.1. A problem on matrix weighted parallelogram law
+Let $\mathbb{M}_{m,n}$ denote the space of all $m\times n$ complex matrices, and write $\mathbb{M}_{n}:=\mathbb{M}_{n,n}$.
+Let $\mathcal{B}(\mathcal{H},\mathcal{K})$ denote the space of all bounded linear operators from a complex Hilbert space $\mathcal{H}$ to a complex Hilbert space $\mathcal{K}$; when $\mathcal{H}=\mathcal{K}$, we simply write $\mathcal{B}(\mathcal{H})$.
+We denote by $\mathbb{K}(\mathcal{H})$ the ideal of compact operators on $\mathcal{H}$; when the underlying Hilbert space is clear, we simply write $\mathbb{K}$.
+1.1. Thompson’s inequality and its development
+In 1976, Thompson [16, 17] obtained his famous triangle inequality for the operator right modulus.
+Theorem 1.1 (Thompson).
+Let $A,B\in\mathbb{M}_{n}$. Then there exist two unitaries $U,V$ such that
+$|A+B|\leq U|A|U^{*}+V|B|V^{*}.$
+Thompson [18] also characterized the equality case in Theorem 1.1.
+Theorem 1.2 (Thompson).
+Let $A,B\in\mathbb{M}_{n}$. The following are equivalent:
+  1. (i)
+    there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+    $|A+B|=U|A|U^{*}+V|B|V^{*};$
+  2. (ii)
+    there exists a unitary $W\in\mathbb{M}_{n}$ such that
+    $A=W|A|,\qquad B=W|B|.$
+Passing from matrices to operators on an infinite-dimensional Hilbert space, the triangle inequality persists with unitaries replaced by isometries thanks to Akemann, Anderson and Pedersen [1]:
+Theorem 1.3 (Akemann–Anderson–Pedersen).
+Let $A,B\in\mathcal{B}(\mathcal{H})$. Then there exist two isometries $U,V$ such that
+$|A+B|\leq U|A|U^{*}+V|B|V^{*}.$
+Naturally, we also ask when equality can occur in Theorem 1.3.
+However, the equality characterization in Theorem 1.2 is genuinely finite-dimensional.
+Indeed, Ando and Hayashi [2] remarked that Theorem 1.2 fails in the infinite-dimensional setting: one can find $A,B\in\mathcal{B}(\mathcal{H})$ and unitaries $U,V$ such that
+$|A+B|=U|A|U^{*}+V|B|V^{*},$
+while there is no partial isometry $W$ satisfying simultaneously
+$A=W|A|\quad\text{and}\quad B=W|B|.$
+For instance, let $P$ be an orthogonal projection and let $Q$ be a subprojection of $P$.
+Assume that $P$, $Q$, and PQP-Q all have infinite rank.
+Then there exist (at least countably many) unitaries $U,V$ for which
+$|P+(-Q)|=U|P|U^{*}+V|{-}Q|V^{*}$
+holds, whereas there is no partial isometry $W$ such that
+$P=W|P|\quad\text{and}\quad(-Q)=W|{-}Q|.$
+Even though Theorem 1.2 breaks down in infinite dimension, Ando and Hayashi [2] showed that the intrinsic triangle equality for the usual modulus still admits a clean description.
+Theorem 1.4 (Ando–Hayashi).
+Let $A,B\in\mathcal{B}(\mathcal{H})$. If
+$\left|A+B\right|=\left|A\right|+\left|B\right|,$
+then there exists a partial isometry $U\in\mathcal{B}(\mathcal{H})$ such that
+$A=U\left|A\right|,\qquad B=U\left|B\right|.$
+Aujla and Bourin [3] established a matrix subadditivity inequality for concave functions.
+Theorem 1.5 (Aujla–Uchiyama).
+Let $A,B\geq 0$ and $f:[0,\infty)\to\mathbb{R}$ be monotone concave with $f(0)\geq 0$. Then there exist unitaries $U,V$ such that
+$f(A+B)\leq Uf(A)U^{*}+Vf(B)V^{*}.$
+Combining Theorems 1.1 and 1.5 with the standard fact that, whenever $0\leq X\leq Y$ and $f$ is nondecreasing,
+$\lambda_{j}\bigl(f(X)\bigr)\leq\lambda_{j}\bigl(f(Y)\bigr),\qquad 1\leq j\leq n,$
+where $\lambda_{j}(\cdot)$ denotes the $j$-th largest eigenvalue, and hence
+$f(X)\leq Wf(Y)W^{*}$
+for some unitary $W$, we obtain the following: for every monotone concave function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\geq 0$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$f(|A+B|)\leq Uf(|A|)U^{*}+Vf(|B|)V^{*},\qquad A,B\in\mathbb{M}_{n},$ (1.1)
+see [12]. Thus Thompson’s inequality extends from the identity function $f(t)=t$ to all monotone concave functions with $f(0)\geq 0$.
+Since $f$ is monotone and $f(0)\geq 0$, we automatically have $f(t)\geq 0$ for all $t\geq 0$.
+Conversely, every concave map $f:[0,\infty)\to[0,\infty)$ is automatically nondecreasing.
+Indeed, if $f$ were not nondecreasing, then there would exist $0\leq s<t$ such that $f(s)>f(t)$, and hence
+$\frac{f(t)-f(s)}{t-s}<0.$
+By concavity, for every $x>t$,
+$\frac{f(x)-f(t)}{x-t}\leq\frac{f(t)-f(s)}{t-s}<0,$
+so that
+$f(x)\leq f(t)+\frac{x-t}{t-s}\bigl(f(t)-f(s)\bigr).$
+Since f(t)f(s)<0f(t)-f(s)<0, the right-hand side tends to -$\infty$ as $x\to\infty$, contradicting the assumption that $f(x)\geq 0$ for all $x\geq 0$.
+Therefore every concave map $f:[0,\infty)\to[0,\infty)$ is nondecreasing.
+This observation shows that, for concave functions on $0,\infty)$, nonnegativity of the range is equivalent to monotonicity.
+Bourin and Lee [7, Remark 3.13], and later Bourin, Harada, and Lee [6, Question 1.3], asked whether the monotonicity assumption in Theorem 1.5 can be removed if one only assumes concavity together with $f(0)\geq 0$.
+Question 1.6 (Bourin–Harada–Lee).
+It is not known whether the monotonicity assumption in Theorem 1.5 can be deleted, that is, whether the theorem remains valid for every concave function $f:[0,\infty)\to\mathbb{R}$ satisfying $f(0)\geq 0$.
+We shall show that the monotonicity assumption in Theorem 1.5 is indispensable.
+More precisely, we construct a counterexample using the concave function $f(t)=t-t^{2}$, thereby answering Question 1.6 in the negative.
+For further developments regarding the technique of unitary orbits, we refer the reader to a good survey [7].
+1.2. A triangle inequality for the quadratic symmetric modulus
+To obtain new triangle inequalities for the newly introduced operator moduli, let us first recall the definition of the modulus in the scalar case.
+For a complex number $z=\Re z+i\,\Im z\in\mathbb{C}$, the modulus is classically defined by
+$|z|:=\sqrt{(\Re z)^{2}+(\Im z)^{2}},$ (1.2)
+which immediately yields
+$|z|=(\bar{z}z)^{\frac{1}{2}}$ (1.3)
+and
+$|z|=\frac{|z|+|\bar{z}|}{2}.$ (1.4)
+For $Z\in\mathbb{M}_{n}$, the right modulus, in analogy with (1.3), is defined by
+$|Z|:=(Z^{*}Z)^{\frac{1}{2}},$
+where $Z^{*}$ denotes the conjugate transpose of $Z$. On the other hand, if we write the Cartesian decomposition $Z=\Re Z+i\,\Im Z$, where
+$\Re Z:=\frac{Z+Z^{*}}{2},\qquad\Im Z:=\frac{Z-Z^{*}}{2i},$
+then a more geometric analogue of (1.2) motivates the definition of the quadratic symmetric modulus:
+$|Z|_{\mathrm{qsym}}:=\sqrt{(\Re Z)^{2}+(\Im Z)^{2}}=\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{\frac{1}{2}}.$
+We also define the arithmetic symmetric modulus, following (1.4), by
+$|Z|_{\mathrm{sym}}:=\frac{|Z|+|Z^{*}|}{2}.$
+We now turn to a related problem raised by Bourin and Lee concerning Thompson-type triangle inequalities for symmetric moduli, in comparison with Theorem 1.1.
+Question 1.7 ([10, Question 4.9]).
+Let $|Z|_{\mathrm{(q)sym}}$ denote either the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$.
+For $A,B\in\mathbb{M}_{n}$, does there exist a pair of unitaries $U,V\in\mathbb{M}_{n}$ such that the following Thompson-type triangle inequality holds?
+$|A+B|_{\mathrm{(q)sym}}\leq U\,|A|_{\mathrm{(q)sym}}\,U^{*}+V\,|B|_{\mathrm{(q)sym}}\,V^{*}.$ (1.5)
+Our next two theorems give a complete answer to Question 1.7: the inequality (1.5) fails for the arithmetic symmetric modulus, whereas it does hold for the quadratic symmetric modulus.
+Theorem 1.8.
+Let $A,B\in\mathbb{M}_{n}$. In general, there do not exist unitaries $U,V$ such that
+$|A+B|_{\mathrm{sym}}\leq U|A|_{\mathrm{sym}}U^{*}+V|B|_{\mathrm{sym}}V^{*}.$
+Indeed, there exist $A,B\in\mathbb{M}_{2}$ such that
+$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty}.$
+Theorem 1.9.
+Let $A,B\in\mathbb{M}_{n}$. Then there exist two unitaries $U,V$ such that
+$|A+B|_{\mathrm{qsym}}\leq U|A|_{\mathrm{qsym}}U^{*}+V|B|_{\mathrm{qsym}}V^{*}.$
+We remark that Theorem 1.9 can be regarded as an operator version of the classical triangle inequality in the complex plane, $|a+b|\leq|a|+|b|$ for $a,b\in\mathbb{C}$, a result of historical importance.
+In contrast to Theorem 1.1, here the modulus is defined via the real and imaginary parts of the operator itself.
+In this sense, it may be viewed as a more faithful operator analogue of the triangle inequality.
+We also characterize the equality case in Theorem 1.9 as follows.
+Theorem 1.10.
+Let $A,B\in\mathbb{M}_{n}$. Assume that there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$\left|A+B\right|_{\mathrm{qsym}}=U\,\left|A\right|_{\mathrm{qsym}}\,U^{*}+V\,\left|B\right|_{\mathrm{qsym}}\,V^{*}.$
+Define the block matrix
+$\Gamma(T):=\frac{1}{\sqrt{2}}\begin{pmatrix}T\\ T^{*}\end{pmatrix}\in\mathbb{M}_{2n,n}.$
+Then there exists a partial isometry $W\in\mathbb{M}_{2n,n}$ such that
+$\Gamma(A)=W\,\left|A\right|_{\mathrm{qsym}},\qquad\Gamma(B)=W\,\left|B\right|_{\mathrm{qsym}}.$
+Equivalently, $\Gamma(A)$ and $\Gamma(B)$ admit polar decompositions sharing a common partial isometry factor.
+The next question is whether Theorem 1.9 remains true in the infinite-dimensional setting.
+Our first observation is that it does: the quadratic symmetric Thompson type inequality extends to $\mathcal{B}(\mathcal{H})$, provided one replaces unitaries by isometries, exactly as in Theorem 1.3.
+Theorem 1.11.
+Let $A,B\in\mathcal{B}(\mathcal{H})$. Then there exist isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
+$\left|A+B\right|_{\mathrm{qsym}}\ \leq\ U\,\left|A\right|_{\mathrm{qsym}}\,U^{*}\;+\;V\,\left|B\right|_{\mathrm{qsym}}\,V^{*}.$
+Having established the triangle inequality, we turn to its equality case.
+Motivated by Theorem 1.4, we prove an Ando–Hayashi type characterization for the quadratic symmetric modulus.
+Since $\left|T\right|_{\mathrm{qsym}}$ is not, in general, the modulus of $T$ itself, the correct formulation uses a canonical lifting.
+Theorem 1.12.
+Let $A,B\in\mathcal{B}(\mathcal{H})$. Assume that
+$\left|A+B\right|_{\mathrm{qsym}}=\left|A\right|_{\mathrm{qsym}}+\left|B\right|_{\mathrm{qsym}}.$
+Define the linear lifting $\Gamma:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H})$ by
+$\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}$.
+Then there exists a partial isometry $W\in\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H})$ such that
+$\Gamma(A)=W\,\left|A\right|_{\mathrm{qsym}},\qquad\Gamma(B)=W\,\left|B\right|_{\mathrm{qsym}}.$
+Equivalently, $\Gamma(A)$ and $\Gamma(B)$ admit polar decompositions sharing a common partial isometry factor.
+Besides triangle inequalities, we also establish a concave-function subadditivity theorem; compare this with (1.1).
+Theorem 1.13.
+Let $A,B\in\mathbb{M}_{n}$ and let $f:[0,\infty)\to[0,\infty)$ be concave. Then there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$f\bigl(\left|A+B\right|_{\mathrm{qsym}}\bigr)\leq Uf\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)U^{*}+Vf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)V^{*}.$
+A number of elegant operator inequalities can be derived from Theorem 1.13. For example, taking $f(t)=t^{p}$ with $0<p<1$ in Theorem 1.13 yields
+$\left|A+B\right|_{\mathrm{qsym}}^{p}\leq U\left|A\right|_{\mathrm{qsym}}^{p}U^{*}+V\left|B\right|_{\mathrm{qsym}}^{p}V^{*},\qquad 0<p<1.$
+[…]
+Corollary 1.14.
+Let $A,B\in\mathbb{M}_{n}$ and let $f:[0,\infty)\to[0,\infty)$ be concave. Then for any unitarily invariant norm, we have
+[…]
+Taking traces in Theorem 1.13 yields a Rotfel’d type trace inequality for the quadratic symmetric modulus.
+Corollary 1.15.
+Let $A,B\in\mathbb{M}_{n}$ and let $f:[0,\infty)\to[0,\infty)$ be concave. Then
+$\operatorname{Tr}\,f\bigl(\left|A+B\right|_{\mathrm{qsym}}\bigr)\leq\operatorname{Tr}\,f\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)+\operatorname{Tr}\,f\bigl(\left|B\right|_{\mathrm{qsym}}\bigr).$
+We can also derive the following corollary, which may be regarded as an operator version of the scalar inequality $|a|-|b|\leq|a-b|$ for $a,b\in\mathbb{C}$.
+Corollary 1.16.
+Let $A,B\in\mathbb{M}_{n}$ and let $f:[0,\infty)\to[0,\infty)$ be concave. Then there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$Uf\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)U^{*}-Vf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)V^{*}\leq f\bigl(\left|A-B\right|_{\mathrm{qsym}}\bigr).$
+If $A\in\mathbb{M}_{n}$ is invertible, its condition number is defined by
+$\kappa(A):=\|A\|_{\infty}\|A^{-1}\|_{\infty},$
+where $\|\cdot\|_{\infty}$ denotes the spectral norm.
+Lee [14] proved the following triangle inequality: if $A_{1},\dots,A_{m}\in\mathbb{M}_{n}$ are invertible matrices whose condition numbers are bounded by $\omega>0$, then
+$\left|A_{1}+\cdots+A_{m}\right|\leq\frac{\omega+1}{2\sqrt{\omega}}\,\bigl(\left|A_{1}\right|+\cdots+\left|A_{m}\right|\bigr).$ (1.6)
+Bourin and Lee [7, p.
+1096, line 4] remarked that the bound (1.6) is independent of the number of operators, although it is a rather small bound, it is not known whether it is sharp, see also [14, Question 2].
+In this paper, we will give an example to illustrate its sharpness.
+Moreover, we obtain the following more refined versions of (1.6) in terms of the three introduced moduli.
+Theorem 1.17.
+Let $A_{1},\dots,A_{m}\in\mathcal{B}(\mathcal{H})$ be invertible operators whose condition numbers are $\kappa_{1},\ldots,\kappa_{m}$. Then
+$\bigl|A_{1}+\cdots+A_{m}\bigr|_{\mathrm{((q)sym)}}\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\,|A_{1}|_{\mathrm{((q)sym)}}+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\,|A_{m}|_{\mathrm{((q)sym)}},$
+where the coefficient function $c(\kappa):=\frac{\kappa+1}{2\sqrt{\kappa}}$ is individually sharp for each term: for every $\kappa\geq 1$, the coefficient $c(\kappa)$ in front of a term of condition number $\kappa$ cannot be replaced by any smaller constant, even if all the other coefficients are kept unchanged.
+1.3. Clarkson–McCarthy type inequalities for the quadratic symmetric modulus
+Let $\|\cdot\|_{p}$ denote the Schatten $p$-norm for $1\leq p\leq\infty$, and the quasi-Schatten $p$-norm for $0<p<1$.
+The classical Clarkson–McCarthy inequalities [15] state that for $A,B\in\mathbb{M}_{n}$ and $p\geq 2$,
+$\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2^{p-1}\bigl(\|A\|_{p}^{p}+\|B\|_{p}^{p}\bigr).$ (1.7)
+For $0<p\leq 2$, the inequality is reversed.
+In particular, (1.7) shows that the space $\mathbb{M}_{n}$ is $p$-uniformly convex for $p\geq 2$.
+Bourin and Lee [8] established the following unitary-orbit refinement of (1.7): for $p\geq 2$, there exist unitaries $U,V$ such that
+$U|A+B|^{p}U^{*}+V|A-B|^{p}V^{*}\leq 2^{p-1}\bigl(|A|^{p}+|B|^{p}\bigr).$ (1.8)
+For $0<p\leq 2$, the inequality is reversed.
+The Clarkson–McCarthy inequalities have since been extended in various directions [4], including to multivariable inequalities involving several operators [5, 20]; for a unified framework, see the author’s recent paper [19].
+We now turn from triangle inequalities to Clarkson–McCarthy type estimates for quadratic symmetric modulus.
+The starting point is an exact square-sum identity, which plays the role of a Parseval formula for the quadratic symmetric modulus.
+Theorem 1.18.
+For a fixed $m\times n$ matrix $U=(u_{ij})$, the operator square-sum identity
+$\sum_{i=1}^{m}\left|\sum_{j=1}^{n}u_{ij}A_{j}\right|_{\mathrm{qsym}}^{2}=\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}$ (1.9)
+holds for all $A_{1},\ldots,A_{n}\in\mathbb{M}_{N}$ if and only if $U$ is an isometry, that is, $U^{*}U=I_{n}.$
+With the square-sum identity in hand, we may combine it with Jensen-type orbit inequalities to derive the desired Clarkson–McCarthy type estimates.
+Theorem 1.19.
+Let $B_{i},A_{j}\in\mathbb{M}_{N}$ for $1\leq i\leq m,1\leq j\leq n$ satisfy $(B_{1},\ldots,B_{m})^{T}=U(A_{1},\ldots,A_{n})^{T}$ for some isometry $U=(u_{ij})$.
+Then for $p\geq 2$, there exist unitaries $U_{i},V_{j}$ such that
+$\sum_{i=1}^{m}U_{i}\left|B_{i}\right|_{\mathrm{qsym}}^{p}U_{i}^{*}$ $\leq n^{\frac{p}{2}-1}\sum_{j=1}^{n}\left|A_{j}\right|_{\mathrm{qsym}}^{p},$
+$\sum_{j=1}^{n}V_{j}\left|A_{j}\right|_{\mathrm{qsym}}^{p}V_{j}^{*}$ $\leq m^{\frac{p}{2}-1}\sum_{i=1}^{m}\left|B_{i}\right|_{\mathrm{qsym}}^{p}.$
+For $0<p\leq 2$, the inequalities are reversed.
+By choosing the normalized Hadamard isometry
+$\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$
+in Theorem 1.19 yields the following two-variable Clarkson–McCarthy type inequality.
+Corollary 1.20.
+Let $A,B\in\mathbb{M}_{n}$. Then for $p\geq 2$, there exist unitaries $U,V$ such that
+$U\left|A+B\right|_{\mathrm{qsym}}^{p}U^{*}+V\left|A-B\right|_{\mathrm{qsym}}^{p}V^{*}\leq 2^{p-1}(\left|A\right|_{\mathrm{qsym}}^{p}+\left|B\right|_{\mathrm{qsym}}^{p}).$
+For $0<p\leq 2$, the inequality is reversed.
+We next combine Theorem 1.18 with certain Jensen-type inequalities for unitarily invariant norms to derive the following Clarkson–McCarthy type estimates.
+Theorem 1.21.
+Let $B_{i},A_{j}\in\mathbb{M}_{N}$ for $1\leq i\leq m,1\leq j\leq n$ satisfy $(B_{1},\ldots,B_{m})^{T}=U(A_{1},\ldots,A_{n})^{T}$ for some isometry $U=(u_{ij})$. Then for $p\geq 2$,
+[…]
+For $0<p\leq 2$, the inequalities are reversed.
+Again, choosing the $2\times 2$ normalized Hadamard isometry gives the familiar two-variable form.
+Corollary 1.22.
+Let $A,B\in\mathbb{M}_{n}$. Then for $p\geq 2$,
+[…]
+For $0<p\leq 2$, the inequality is reversed.
+Taking the trace in Theorem 1.21 or the trace norm in Theorem 1.21 recovers the following Schatten $p$-norm version.
+Corollary 1.23.
+Let $B_{i},A_{j}\in\mathbb{M}_{N}$ for $1\leq i\leq m,1\leq j\leq n$ satisfy $(B_{1},\ldots,B_{m})^{T}=U(A_{1},\ldots,A_{n})^{T}$ for some isometry $U=(u_{ij})$. Then for $p\geq 2$,
+$m^{1-\frac{p}{2}}\sum_{j=1}^{n}\||A_{j}|_{\mathrm{qsym}}\|_{p}^{p}\leq\sum_{i=1}^{m}\||B_{i}|_{\mathrm{qsym}}\|_{p}^{p}\leq n^{\frac{p}{2}-1}\sum_{j=1}^{n}\||A_{j}|_{\mathrm{qsym}}\|_{p}^{p}.$
+For $0<p\leq 2$, the inequalities are reversed.
+Corollary 1.24.
+Let $A,B\in\mathbb{M}_{n}$. Then for $p\geq 2$,
+$\|\left|A+B\right|_{\mathrm{qsym}}\|_{p}^{p}+\|\left|A-B\right|_{\mathrm{qsym}}\|_{p}^{p}\leq 2^{p-1}(\|\left|A\right|_{\mathrm{qsym}}\|_{p}^{p}+\|\left|B\right|_{\mathrm{qsym}}\|_{p}^{p}).$
+For $0<p\leq 2$, the inequality is reversed.
+1.4. Questions of Bourin and Lee on symmetric moduli and Euler-type identities
 For $A,B\in\mathbb{M}_{n}$ and $x\in\mathbb{R}$, we write $A\nabla_{x}B:=(1-x)A+xB$ for the weighted arithmetic mean of $A$ and $B$.
 We begin by recalling a unitary-orbit refinement of the weighted parallelogram identity due to Bourin and Lee [10].
-Theorem 1.1 ([12, Corollary 2.2]).
+Theorem 1.25 ([10, Corollary 2.2]).
 Let $A,B\in\mathbb{M}_{n}$ and $0\leq x\leq 1$. Then there exist isometries $U,V,S,T\in\mathbb{M}_{2n,n}$ such that
@@
 for $0\leq x\leq 1$. Since the right-hand side of that trace identity is a polynomial in $x$, it extends to all $x\in\mathbb{R}$. This motivates the following question.
-Question 1.2 ([12, Question 2.3]).
+Question 1.26 ([10, Question 2.3]).
 Does Theorem 1.25 hold for every $x\in\mathbb{R}$?
-We first answer Question 1.2 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.1 cannot be removed and therefore is sharp (see Theorem 1.3 below).
+We first answer Question 1.26 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.25 cannot be removed and therefore is sharp (see Theorem 1.27 below).
 Theorem 1.27.
 Fix $x\in\mathbb{R}\setminus[0,1]$. Then Theorem 1.25 fails for this $x$ and some $A,B\in\mathbb{M}_{n}$. In fact, a counterexample exists already for $n=1$.
-1.2. Symmetric moduli and three related questions
-For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite. While for $A,B\in\mathbb{M}_{n}$ one has the triangle inequality
-$\|A+B\|_{1}\leq\|A\|_{1}+\|B\|_{1},$
-simple examples show that the inequality
-$|A+B|\leq|A|+|B|$
-need not hold in the positive semidefinite order. However, Thompson [20] observed that there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
-$|A+B|\leq U|A|U^{*}+V|B|V^{*},$ (1.1)
-which, upon taking traces, yields the triangle inequality as a consequence.
-Such “unitary-orbit” inequalities have since been generalized and extended to a variety of settings; see the surveys by Bourin, Harada and Lee [7, 9].
-In this paper, we generalize (1.1) to the setting of rectangular matrices (see Lemma 5.2) in order to address Question 1.11.
-Recall the Cartesian decomposition of a matrix $Z\in\mathbb{M}_{n}$:
-$Z=\operatorname{Re}Z+i\,\operatorname{Im}Z,\qquad\operatorname{Re}Z:=\frac{Z+Z^{*}}{2},\quad\operatorname{Im}Z:=\frac{Z-Z^{*}}{2i}.$
-Note that $|\operatorname{Re}Z|^{2}+|\operatorname{Im}Z|^{2}=\left(|Z|^{2}+|Z^{*}|^{2}\right)/2$.
-Bourin and Lee [12, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
+Bourin and Lee [10, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
 Theorem 1.28 (Bourin–Lee).
@@
 $\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\leq U|\operatorname{Re}\,Z|U^{*}+V|\operatorname{Im}\,Z|V^{*}.$
-They also posed the following question concerning the optimality of the exponent $2$.
-Question 1.5 ([12, Question 4.6]).
+Clearly, since the function $t\mapsto t^{p/2}$ is operator concave on $0,\infty)$ for $0<p<2$, we have
+$\frac{|Z|^{p}+|Z^{*}|^{p}}{2}=\frac{(|Z|^{2})^{p/2}+(|Z^{*}|^{2})^{p/2}}{2}\leq\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{p/2}.$
+Since both sides are positive semidefinite and the function $t\mapsto t^{1/p}$ is increasing on $0,\infty)$, Weyl’s monotonicity theorem yields
+$\lambda_{j}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)\leq\lambda_{j}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)$
+for all $j$, where $\lambda_{j}(\cdot)$ denotes the $j$-th largest eigenvalue. Therefore,
+$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq W\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\,W^{*}$
+for some unitary matrix $W$. Therefore, by Theorem 1.28, for any $0<p<2$, we obtain
+$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq U|\operatorname{Re}Z|U^{*}+V|\operatorname{Im}Z|V^{*}.$
+Bourin and Lee [10] then raised the following question concerning whether the exponent $2$ is optimal.
+Question 1.29 ([10, Question 4.6]).
 Let $p>0$. Suppose that, for every $Z\in\mathbb{M}_{n}$, one can find unitary matrices $U,V\in\mathbb{M}_{n}$ such that
@@
 Must we necessarily have $p\leq 2$?
-Our second result answers Question 1.5 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
+Our next result answers Question 1.29 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
 Theorem 1.30.
@@
 Remark 1.31.
-For $n=1$, the statement of Question 1.5 is different: for a complex number $z$, the left-hand side of (1.2) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$.
+For $n=1$, the statement of Question 1.29 is different: for a complex number $z$, the left-hand side of (1.10) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$.
 Hence (1.10) holds for all $p>0$ in dimension 11.
@@
 Remark 1.32.
-In [12, Corollary 4.7], the indices on the right-hand side appear with a typographical slip.
-Starting from the operator inequality in [12, Corollary 4.3] and applying the Weyl inequality for singular values,
+In [10, Corollary 4.7], the indices on the right-hand side appear with a typographical slip.
+Starting from the operator inequality in [10, Corollary 4.3] and applying the Weyl’s inequality for singular values,
 $\mu_{1+j+k}^{\downarrow}(A+B)\leq\mu_{1+j}^{\downarrow}(A)+\mu_{1+k}^{\downarrow}(B)\qquad(A,B\geq 0),$
-one obtains the corrected bound
-$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$ (1.3)
-Accordingly, the natural corrected version of [12, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
+where $\mu_{j}^{\downarrow}(\cdot)$ denotes the $j$-th largest singular value, one obtains the corrected bound
+$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$
+Accordingly, the natural corrected version of [10, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
 $\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
 for some integers $j,k\geq 0$.
-Question 1.9 ([12, Question 4.8]).
+Question 1.33 ([10, Question 4.8]).
 Does there exist $Z\in\mathbb{K}$ such that for every $p>2$ one has
@@
 for some pair of integers $j,k\geq 0$?
-We answer Question 1.9 in the affirmative.
+We answer Question 1.33 in the affirmative.
 Theorem 1.34.
 There exists $Z\in\mathbb{K}$ and fixed integers $j,k\geq 0$ such that for every $p>2$, (1.11) holds. In our construction one may take $(j,k)=(2,0)$.
-Let
-$|Z|_{\mathrm{sym}}:=\frac{|Z|+|Z^{*}|}{2},\qquad|Z|_{\mathrm{qsym}}:=\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{1/2}$
-be the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$, respectively.
-Now, we turn to a related problem raised by Bourin and Lee concerning Thompson-type triangle inequalities for symmetric moduli, compared with (1.1).
-Question 1.11 ([12, Question 4.9]).
-Denote by $|Z|_{\mathrm{(q)sym}}$ either the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$.
-Let $X,Y\in\mathbb{M}_{n}$.
-Does the Thompson triangle type inequality
-$|X+Y|_{\mathrm{(q)sym}}\leq U\,|X|_{\mathrm{(q)sym}}\,U^{*}+V\,|Y|_{\mathrm{(q)sym}}\,V^{*}.$ (1.5)
-hold for some unitary $U,V\in\mathbb{M}_{n}$ ?
-Our next theorem provides a complete answer to Question 1.11: the inequality (1.5) fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
-Theorem 1.12.
-Let $X,Y\in\mathbb{M}_{n}$.
-  1. (i)
-    In general, there do not exist unitaries $U,V$ such that
-    $|X+Y|_{\mathrm{sym}}\leq U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}.$
-  2. (ii)
-    There exist unitaries $U,V$ such that
-    $|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*}.$
-1.3. Matrix analogues of Euler’s quadrilateral identity
-Finally, Bourin and Lee [12] asked for matrix analogues of Euler’s quadrilateral identity.
-Question 1.13 ([12, Question 4.10]).
+Bourin and Lee [10] finally asked for matrix analogues of Euler’s quadrilateral identity.
+Question 1.35 ([10, Question 4.10]).
 What are the matrix versions, if any, of Euler’s quadrilateral identity
-$\|x+y+z\|^{2}+\|x\|^{2}+\|y\|^{2}+\|z\|^{2}=\|x+y\|^{2}+\|y+z\|^{2}+\|z+x\|^{2}$
+$\|x+y+z\|_{2}^{2}+\|x\|_{2}^{2}+\|y\|_{2}^{2}+\|z\|_{2}^{2}=\|x+y\|_{2}^{2}+\|y+z\|_{2}^{2}+\|z+x\|_{2}^{2}$
 for three points $x,y,z\in\mathbb{C}^{n}$?
@@
 $|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}=|A+B|^{2}+|B+C|^{2}+|C+A|^{2}.$ (1.12)
-We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$.
-Bourin and Lee [12] suggested exploring isometry/unitary-orbit refinements of Euler-type identities.
-Our following three results give isometry-orbit refinements of (1.6).
+Bourin and Lee [10] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following three results give isometry-orbit refinements of (1.12).
 Theorem 1.37.
@@
 $|A+B+C|^{2}\oplus|A|^{2}\oplus|B|^{2}\oplus|C|^{2}=\frac{1}{4}\sum_{k=1}^{4}U_{k}\,\left(|A+B|^{2}+|B+C|^{2}+|A+C|^{2}\right)\,U_{k}^{*}.$
+Remark 1.40.
+The Euler-type orbit identities in Proposition 1.36 and Theorems 1.38 and 1.39 also admit quadratic symmetric modulus versions.
+Indeed, these two results extend verbatim to rectangular matrices with a fixed number of columns, since their proofs only involve the positive matrices $|A|^{2}=A^{*}A$ and the Euler identity (1.12), which remains valid in the rectangular setting.
+Applying these rectangular versions to the lifting $\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}\in\mathbb{M}_{2n,n},$ we obtain the corresponding analogues for quadratic symmetric modulus.
 Naturally, we ask the following question.
-Question 1.18.
+Question 1.41.
 Does there exist isometries $U_{1},U_{2},U_{3},U_{4}\in\mathbb{M}_{3n,n}$ such that
@@
 $\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
-Unfortunately, the answer to Question 1.18 is negative. In fact, take $A=B=C=I_{n}$. Then
+Unfortunately, the answer to Question 1.41 is negative. In fact, take $A=B=C=I_{n}$. Then
 $|A+B|^{2}=|2I_{n}|^{2}=4I_{n},\qquad|A+B+C|^{2}=|3I_{n}|^{2}=9I_{n},\qquad|A|^{2}=|B|^{2}=|C|^{2}=I_{n}.$
@@
 But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|_{\infty}=9$ while $\|4I_{3n}\|_{\infty}=4$, a contradiction. Therefore such isometries cannot exist in general.
-1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
-For a Hilbert space $\mathcal{H}$, let $S_{p}$ denote the Schatten $p$-class.
-For $p>0$, we say that $X\in S_{p}$ if $X$ is a compact operator on $\mathcal{H}$ whose singular values $(\mu_{k}^{\downarrow}(X))_{k\geq 1}$ satisfy $\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}<\infty$ .
-In this case, the Schatten pp-(quasi)norm is defined by
-$\|X\|_{p}:=\Big(\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}\Big)^{1/p}.$
-(When $0<p<1$, $\|\cdot\|_{p}$ is a quasi-norm.)
-Building on the Euler operator identity (1.6), we obtain the following sharp Clarkson–McCarthy type inequalities.
-Theorem 1.19.
-Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
-$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big).$
-For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
-As a complement to Theorem 1.19, numerical evidence suggests the following sharp inequality. For a weaker bound with coefficient $3^{\frac{p}{2}-1}$, see Theorem 7.10.
-Conjecture 1.20.
-Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
-$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\leq\frac{3^{p-1}+1}{2^{p}}\left(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\right).$
-For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $\frac{3^{p-1}+1}{2^{p}}$ is optimal: equality holds for all $p>0$ by taking $A=B=C\neq 0$.
-Classical Hlawka-type inequalities compare norms (or their powers) of vector sums with those of their pairwise sums.
-A standard form of Hlawka’s inequality [17] asserts that if $V$ is an inner product space and $x,y,z\in V$, then
-$\|x+y\|+\|y+z\|+\|z+x\|\ \leq\ \|x\|+\|y\|+\|z\|+\|x+y+z\|,$
-where $\|\cdot\|$ denotes the norm induced by the inner product; equivalently, the inequality holds for any norm satisfying the parallelogram identity.
-Numerous refinements and variants replace $\|\cdot\|$ by powers $\|\cdot\|^{p}$ and/or adjust the constants.
-Theorem 1.19 and Conjecture 1.20 can be viewed as a Schatten–$p$ (i.e.
-noncommutative $L_{p}$) analogue: it features a sharp $p$-dependent constant and exhibits a change in behavior at $p=2$, as is typical for inequalities governed by the uniform convexity/concavity of $L_{p}$ spaces.
-As a further refinement of Theorem 1.19, we establish the following sharp mixed $\ell_{q}$$\ell_{p}$ Clarkson–McCarthy type inequalities.
-Theorem 1.21.
-Let $A,B,C\in S_{p}$. Then for $1<p\leq 2$ ,
-$\|A+B+C\|_{p}^{q}+\|A\|_{p}^{q}+\|B\|_{p}^{q}+\|C\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Big)^{q/p},$
-$\|A+B\|_{p}^{q}+\|B+C\|_{p}^{q}+\|C+A\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\Big)^{q/p},$
-where $q$ is the conjugate exponent of $p$.
-For $p\geq 2$, these two inequalities are reversed.
-Moreover, the constant $2^{1-\frac{q}{p}}$ is optimal: equality holds for all $p>1$ by taking $A=B=-C\neq 0$.
+1.5. Future directions
+After early versions of this paper were completed, some related works appeared.
+Bourin and Lee [11] established triangle inequalities for unitarily invariant norms on arithmetic symmetric modulus, while the author [21] studied triangle inequalities for unitarily invariant norms and Schatten $p$-norms on two symmetric moduli.
+Future directions for further study include the Lipschitz continuity of these moduli and related problems.
+1.6. Organization of this paper
+The paper is organized as follows.
+Section 2 answers Question 1.6 and proves Theorem 1.8.
+Section 3 establishes a rectangular Thompson’s inequality and uses it to prove Theorems 1.9 and 1.10.
+Section 4 extends the quadratic symmetric Thompson inequality to operators on Hilbert spaces and proves the corresponding equality characterization.
+Section 5 proves the concave-function subadditivity theorem for the quadratic symmetric modulus and derives its norm and trace consequences.
+Section 6 obtains condition-number estimates for the symmetric moduli and shows that the coefficient is sharp.
+Section 7 develops Clarkson–McCarthy type inequalities for the quadratic symmetric modulus, both in operator order and for unitarily invariant norms.
+Sections 8, 9, and 10 answer three questions raised in [10].
+Finally, Section 11 establishes Euler-type operator identities together with several isometry-orbit refinements.
+Acknowledgments
+The author is very grateful to Professor Bourin for his valuable suggestions.
+This work is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No.
+xzy022024045).
+2. A counterexample to Question 1.6 and proof of Theorem 1.8
+In this section, we answer Question 1.6 and establish Theorem 1.8.
+Answer to Question 1.6. The answer is negative. Indeed, consider
+$f(t)=t-t^{2},\qquad t\geq 0.$
+Then $f$ is concave on $0,\infty)$ and $f(0)=0$. Let
+$A=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\qquad B=\begin{pmatrix}c^{2}&cs\\ cs&s^{2}\end{pmatrix},\qquad 0<c<1,\quad s=\sqrt{1-c^{2}}.$
+Since $A$ and $B$ are rank-one projections, we have $A^{2}=A$ and $B^{2}=B$, and therefore
+$f(A)=A-A^{2}=0,\qquad f(B)=B-B^{2}=0.$
+Thus, for all unitaries $U,V$,
+$Uf(A)U^{*}+Vf(B)V^{*}=0.$
+If the conclusion of Theorem 1.5 were valid under the weaker assumption $f(0)=0$ and concavity alone, then we would have $f(A+B)\leq 0$.
+However, $A+B$ has trace $2$ and determinant $1-c^{2}$, so its eigenvalues are $1\pm c$. Consequently, the eigenvalues of $f(A+B)$ are
+$f(1+c)=-c(1+c)\qquad\text{and}\qquad f(1-c)=c(1-c).$
+Since $0<c<1$, the latter is strictly positive. Hence $f(A+B)$ has a positive eigenvalue, so $f(A+B)\nleq 0$. Therefore there do not exist unitaries $U,V$ such that
+$f(A+B)\leq Uf(A)U^{*}+Vf(B)V^{*}.$
+This proves that Theorem 1.5 does not extend to arbitrary concave functions satisfying only $f(0)=0$. In particular, the monotonicity assumption cannot be removed.∎
+Proof of Theorem 1.8.
+Take
+$A=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad B=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in\mathbb{M}_{2}.$
+A direct computation gives
+$|A|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|B|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|A+B|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
+Hence
+$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
+so
+$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e.
+}\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
+If the inequality in Theorem 1.8 held for some unitaries $U,V$, then taking operator norms would yield
+$\bigl\||A+B|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|A|_{\mathrm{sym}}U^{*}+V|B|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||A|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||B|_{\mathrm{sym}}\bigr\|_{\infty},$
+a contradiction. ∎
+3. A rectangular Thompson’s inequality and proof of Theorem 1.9
+In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.9.
+The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$.
+We then use the rectangular Thompson’s inequality to prove Theorem 1.9.
+Lemma 3.1.
+Let $A\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
+$A=W|A|,\qquad|A|=(A^{*}A)^{1/2},$
+and
+$W^{*}W=\mathrm{supp}(|A|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|A|=|A|,$
+where $\mathrm{supp}(\cdot)$ means the support projection.
+Proof.
+Take a singular value decomposition $A=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
+$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(A).$
+Then
+$A^{*}A=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
+hence
+$|A|=(A^{*}A)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
+Define $W:=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\in\mathbb{M}_{m,n}.$ Clearly, $W^{*}W\leq I_{n},WW^{*}\leq I_{m}$ and
+$W|A|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=A.$
+Moreover,
+$(W^{*}W)\,|A|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|A|.$
+That is, $W^{*}W$ is exactly the support projection of $|A|$. ∎
+Lemma 3.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|A|$.
+This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
+Proposition 3.2.
+Let $A,B\in\mathbb{M}_{m,n}$. Then there exist two unitaries $U,V$ such that
+$|A+B|\leq U|A|U^{*}+V|B|V^{*}.$
+Proof.
+By using Lemma 3.1, there exists a partial isometry $W$ such that A+B=W|A+B|A+B=W\,|A+B|, $W^{*}W$ is the support projection of $|A+B|$ and $W^{*}W\,|A+B|=|A+B|$. Therefore
+$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
+In particular,
+$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (3.1)
+Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices.
+Applying Thompson’s inequality (Theorem 1.1) to the square matrices $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
+$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (3.2)
+Since the left-hand side equals $|A+B|$, it remains to compare $|W^{*}A|$ with $|A|$.
+Since $WW^{*}\leq I_{m}$, we have
+$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
+By the operator monotonicity of $t\mapsto t^{1/2}$ on $0,\infty)$, this implies $|W^{*}A|\leq|A|$. Similarly, $|W^{*}B|\leq|B|$. Conjugating preserves the order, hence
+$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
+Substituting into (3.2) and combining with (3.1) yields our desired result. ∎
+We now prove Theorem 1.9, it follows from an application of Proposition 3.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
+Proof of Theorem 1.9.
+Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
+$\Gamma(T):=\frac{1}{\sqrt{2}}\begin{pmatrix}T\\ T^{*}\end{pmatrix}$.
+Then
+$\Gamma(A+B)=\Gamma(A)+\Gamma(B)\quad\text{ and }\quad|\Gamma(T)|=\,|T|_{\mathrm{qsym}}.$
+Applying rectangular Thompson’s inequality (Proposition 3.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
+$|\Gamma(A+B)|$ $=|\Gamma(A)+\Gamma(B)|$
+$\leq U|\Gamma(A)|U^{*}+V|\Gamma(B)|V^{*}.$
+That is,
+$|A+B|_{\mathrm{qsym}}\leq U|A|_{\mathrm{qsym}}U^{*}+V|B|_{\mathrm{qsym}}V^{*}.$
+We now turn to the equality case. The argument is most transparent after embedding the rectangular problem into a square one, where Thompson’s equality theorem is available.
+Proof of Theorem 1.10.
+Define
+$A_{0}:=\Gamma(A)\in\mathbb{M}_{2n,n},\qquad B_{0}:=\Gamma(B)\in\mathbb{M}_{2n,n}.$
+Then
+$|A_{0}|=\left|A\right|_{\mathrm{qsym}},\qquad|B_{0}|=\left|B\right|_{\mathrm{qsym}},\qquad|A_{0}+B_{0}|=\left|A+B\right|_{\mathrm{qsym}}.$
+Hence the assumption
+$\left|A+B\right|_{\mathrm{qsym}}=U\left|A\right|_{\mathrm{qsym}}U^{*}+V\left|B\right|_{\mathrm{qsym}}V^{*}$
+is equivalent to
+$|A_{0}+B_{0}|=U|A_{0}|U^{*}+V|B_{0}|V^{*}.$ (3.3)
+Step 1: Embed the rectangular problem into a square one. Consider the block embedding of $\mathbb{M}_{2n,n}$ into $\mathbb{M}_{3n}$ given by
+$\widetilde{A}:=\begin{pmatrix}0&0\\ A_{0}&0\end{pmatrix},\qquad\widetilde{B}:=\begin{pmatrix}0&0\\ B_{0}&0\end{pmatrix}\quad\text{in }\mathbb{M}_{3n},$
+with respect to the decomposition $\mathbb{C}^{3n}\cong\mathbb{C}^{n}\oplus\mathbb{C}^{2n}$. A direct computation shows
+$|\widetilde{A}|=\operatorname{diag}(|A_{0}|,0),\qquad|\widetilde{B}|=\operatorname{diag}(|B_{0}|,0),\qquad|\widetilde{A}+\widetilde{B}|=\operatorname{diag}(|A_{0}+B_{0}|,0).$
+Let
+$\widetilde{U}:=\operatorname{diag}(U,I_{2n}),\qquad\widetilde{V}:=\operatorname{diag}(V,I_{2n})\quad\text{in }\mathbb{M}_{3n}.$
+Then (3.3) implies
+$|\widetilde{A}+\widetilde{B}|=\operatorname{diag}(|A_{0}+B_{0}|,0)=\operatorname{diag}(U|A_{0}|U^{*}+V|B_{0}|V^{*},0)=\widetilde{U}|\widetilde{A}|\widetilde{U}^{*}+\widetilde{V}|\widetilde{B}|\widetilde{V}^{*}$.
+Step 2: Apply Thompson’s equality theorem in $\mathbb{M}_{3n}$. By Theorem 1.2, there exists a unitary $\widetilde{W}\in\mathbb{M}_{3n}$ such that
+[…]
+Write $\widetilde{W}$ in block form (with respect to $\mathbb{C}^{n}\oplus\mathbb{C}^{2n}$):
+$\widetilde{W}=\begin{pmatrix}W_{11}&W_{12}\\ W_{21}&W_{22}\end{pmatrix}$.
+[…]
+$\begin{pmatrix}0&0\\ A_{0}&0\end{pmatrix}=\begin{pmatrix}W_{11}&W_{12}\\ W_{21}&W_{22}\end{pmatrix}\begin{pmatrix}|A_{0}|&0\\ 0&0\end{pmatrix}=\begin{pmatrix}W_{11}|A_{0}|&0\\ W_{21}|A_{0}|&0\end{pmatrix}$.
+Hence
+$A_{0}=W_{21}|A_{0}|\qquad\text{and}\qquad W_{11}|A_{0}|=0.$
+[…]
+$B_{0}=W_{21}|B_{0}|\qquad\text{and}\qquad W_{11}|B_{0}|=0.$
+Step 3: Construct the common partial isometry. Let
+$P:=\operatorname{supp}(|A_{0}|+|B_{0}|)\in\mathbb{M}_{n}.$
+Then $P|A_{0}|=|A_{0}|$ and $P|B_{0}|=|B_{0}|$, so if we define
+$W:=W_{21}P\in\mathbb{M}_{2n,n},$
+then
+$W|A_{0}|=W_{21}P|A_{0}|=W_{21}|A_{0}|=A_{0},\qquad W|B_{0}|=W_{21}P|B_{0}|=W_{21}|B_{0}|=B_{0}.$
+Moreover, since $W_{11}|A_{0}|=0$ and $W_{11}|B_{0}|=0$, we have
+$W_{11}(|A_{0}|+|B_{0}|)=0,$
+hence $W_{11}P=0$. Therefore
+$\widetilde{W}\,\operatorname{diag}(P,0)=\begin{pmatrix}W_{11}P&0\\ W_{21}P&0\end{pmatrix}=\begin{pmatrix}0&0\\ W&0\end{pmatrix}$.
+Because $\widetilde{W}$ is unitary, $\widetilde{W}\,\operatorname{diag}(P,0)$ is a partial isometry with initial projection $\operatorname{diag}(P,0)$. Taking adjoints and multiplying yields
+$\begin{pmatrix}0&W^{*}\\ 0&0\end{pmatrix}\begin{pmatrix}0&0\\ W&0\end{pmatrix}=\operatorname{diag}(P,0),$
+and hence
+$W^{*}W=P.$
+So $W$ is a partial isometry in $\mathbb{M}_{2n,n}$.
+Finally, since $A_{0}=\Gamma(A)$, $B_{0}=\Gamma(B)$, $|A_{0}|=\left|A\right|_{\mathrm{qsym}}$, and $|B_{0}|=\left|B\right|_{\mathrm{qsym}}$, we conclude that
+$\Gamma(A)=W\,\left|A\right|_{\mathrm{qsym}},\qquad\Gamma(B)=W\,\left|B\right|_{\mathrm{qsym}}.$
+Equivalently, the polar decompositions of $\Gamma(A)$ and $\Gamma(B)$ share the same partial isometry factor $W$. ∎
+4. Infinite-dimensional extensions and equality cases
+In this section, we discuss the infinite-dimensional case of the triangle inequality for the quadratic symmetric modulus.
+We begin by recording a rectangular version of Theorem 1.3 for operators in $\mathcal{B}(\mathcal{H},\mathcal{K})$.
+Proposition 4.1.
+Let $\mathcal{H},\mathcal{K}$ be complex Hilbert spaces and let $A,B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Then there exist isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
+$\left|A+B\right|\ \leq\ U\,\left|A\right|\,U^{*}\;+\;V\,\left|B\right|\,V^{*}.$
+Proof.
+Let $A+B=W\,\left|A+B\right|$ be the polar decomposition, where $W\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is a partial isometry.
+Recall that for an operator $S$ we write $\operatorname{Ran}(S)$ for its range.
+For a positive operator $S$ we denote by $\operatorname{supp}(S)$ its support projection, i.e.
+the orthogonal projection onto $\overline{\operatorname{Ran}(S)}$.
+Then
+$W^{*}(A+B)=W^{*}W\,\left|A+B\right|=\left|A+B\right|,$
+since $W^{*}W=\operatorname{supp}(\left|A+B\right|)$. Hence
+$\left|A+B\right|=\left|W^{*}(A+B)\right|=\left|W^{*}A+W^{*}B\right|.$
+Now $W^{*}A,W^{*}B\in\mathcal{B}(\mathcal{H})$ are operators on $\mathcal{H}$, so by Akemann–Anderson–Pedersen’s inequality (Theorem 1.3) there exist isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
+$\left|W^{*}A+W^{*}B\right|\ \leq\ U\,\left|W^{*}A\right|\,U^{*}\;+\;V\,\left|W^{*}B\right|\,V^{*}.$
+Finally, since $WW^{*}\leq I_{\mathcal{K}}$ we have
+$\left|W^{*}A\right|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=\left|A\right|^{2},$
+hence $\left|W^{*}A\right|\leq\left|A\right|$ by operator monotonicity of $t\mapsto t^{1/2}$.
+Similarly $\left|W^{*}B\right|\leq\left|B\right|$.
+Conjugation preserves order, so $U\left|W^{*}A\right|U^{*}\leq U\left|A\right|U^{*}$ and $V\left|W^{*}B\right|V^{*}\leq V\left|B\right|V^{*}$.
+Combining the last three displays yields the desired inequality.
+∎
+Now, we give a proof of Theorem 1.11.
+Proof of Theorem 1.11.
+Define the linear lifting
+$\Gamma:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H}),\qquad\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}$.
+Then
+$\Gamma(A+B)=\Gamma(A)+\Gamma(B)\quad\text{ and }\quad|\Gamma(T)|=\,|T|_{\mathrm{qsym}}.$
+Applying Proposition 4.1 yields isometries $U,V\in\mathcal{B}(\mathcal{H})$ such that
+$|\Gamma(A+B)|$ $=|\Gamma(A)+\Gamma(B)|$
+$\leq U|\Gamma(A)|U^{*}+V|\Gamma(B)|V^{*}.$
+That is,
+$|A+B|_{\mathrm{qsym}}\leq U|A|_{\mathrm{qsym}}U^{*}+V|B|_{\mathrm{qsym}}V^{*}.$
 Remark 4.2.
-Setting C=BC=-B in Theorems 1.19 and 1.21 recovers the classical Clarkson–McCarthy inequality of [16].
-For a complete characterization of the validity of Clarkson–McCarthy type inequalities, see the author’s paper [21, Corollary 1.14] and the following Theorem 1.23 (also see [21, Theorem 1.16]).
-For further results and developments on Clarkson–McCarthy type inequalities, we refer the reader to [5, 6, 9, 10, 15, 21, 22].
-We now record the following mixed-norm inequality (Theorem 1.23), which strengthens the result in the author’s earlier work [21, Theorem 1.16].
-Unlike the approach in [21], which applies the three-lines theorem to a suitable trace functional and then invokes a duality argument, we present a new proof based on a vector-valued Riesz–Thorin type complex interpolation scheme.
-Theorem 1.23 (Mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimates).
-Let $U=[u_{ij}]\in\mathbb{M}_{s,t}$ and set
-$\mu:=\max_{1\leq i\leq s,\ 1\leq j\leq t}|u_{ij}|.$
-For $1<p<\infty$, let $q$ be the conjugate exponent of $p$, i.e. $\frac{1}{p}+\frac{1}{q}=1$. Define the linear map on $t$-tuples
-$T_{U}:\ \ell_{p}^{t}(S_{p})\longrightarrow\ell_{q}^{s}(S_{p}),\qquad T_{U}(Z_{1},\dots,Z_{t}):=\Big(\sum_{j=1}^{t}u_{1j}Z_{j},\ \dots,\ \sum_{j=1}^{t}u_{sj}Z_{j}\Big).$
-We will regard $T_{U}$ as an operator between $\ell_{p}$- and $\ell_{q}$-sums when equipped with the corresponding norms.
-  1. (i)
-    Assume that $U$ is a contraction, i.e. $\|U\|_{\infty}\leq 1$ (in particular, this holds if $U^{*}U=I_{t}$ or $UU^{*}=I_{s}$). Then for $1<p\leq 2$,
-    $\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{p}\Bigg)^{1/p},$ (1.8)
-    and the inequality is reversed for $p\geq 2$ if $U^{*}U=I_{t}$.
-  2. (ii)
-    Assume in addition that $U^{*}U=I_{t}$ (i.e. $U$ is an isometry). Then for $1<p\leq 2$,
-    $\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{p}\Bigg)^{1/p},$ (1.9)
-    and the inequality is reversed for $p\geq 2$.
-Theorem 1.23 has several interesting consequences.
-For example, Audenaert–Kittaneh’s conjecture [1, Section 8.1, Conjecture 7] (proved by the author in [22]) states that if $A_{1},\ldots,A_{n}\in S_{p}$, then for $1<p\leq 2$,
-$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\ \leq\ n\left(\sum_{i=1}^{n}\|A_{i}\|_{p}^{p}\right)^{q/p},$ (1.10)
-where $q$ is the conjugate exponent of $p$; for $p\geq 2$ the inequality is reversed.
-Moreover, the constant $n$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
-This inequality can be recovered from Theorem 1.23 by choosing the isometry
-$U=\frac{1}{\sqrt{n}}\begin{pmatrix}1&1&\cdots&1\\ 1&-1&\cdots&0\\ 1&0&\ddots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1\end{pmatrix}\in\mathbb{M}_{\frac{n(n-1)+2}{2},\,n},$
-whose rows encode the combinations $\sum_{i=1}^{n}A_{i}$ and $A_{i}-A_{j}$. Moreover, Theorem 1.23 yields the following sharp reverse-type companion to (1.10). For more details, see Subsection 7.4.
-Corollary 1.24.
-Let $A_{1},\ldots,A_{n}\in S_{p}$. Then for $1<p\leq 2$,
-$\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\leq n^{-q/p}\left(\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{\frac{q}{p}},$ (1.11)
-where $q$ stands for the conjugate exponent of $p$.
-For $p\geq 2$, (1.11) is reversed.
-Moreover, the constant $n^{-q/p}$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
-As further consequences of Euler’s identity (1.6), we next derive sharp Clarkson–McCarthy type inequalities for general unitarily invariant norms, together with a unitary-orbit refinement.
-Theorem 1.25.
-Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$,
+In finite dimension, every isometry is unitary. Hence Theorem 1.11 implies Theorem 1.9.
+Before proving Theorem 1.12, we need the standard two-space extension of Theorem 1.4 to operators between Hilbert spaces.
+Proposition 4.3.
+Let $\mathcal{H},\mathcal{K}$ be complex Hilbert spaces and let $A,B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. If
+$\left|A+B\right|=\left|A\right|+\left|B\right|,$
+then there exists a partial isometry $W\in\mathcal{B}(\mathcal{H},\mathcal{K})$ such that
+$A=W\left|A\right|,\qquad B=W\left|B\right|.$
+Proof.
+Consider $\mathcal{H}\oplus\mathcal{K}$ and define block operators
+$\widetilde{A}:=\begin{pmatrix}0&0\\ A&0\end{pmatrix},\qquad\widetilde{B}:=\begin{pmatrix}0&0\\ B&0\end{pmatrix}\quad\text{in }\mathcal{B}(\mathcal{H}\oplus\mathcal{K}).$
 […]
-For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
-Theorem 1.26.
-Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$, there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
-$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-2}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
-For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
-Organization of this paper.
-Section 2 gives a negative answer to the extension problem for the matrix-weighted parallelogram law: for each $x\notin[0,1]$ we construct a counterexample, proving Theorem 1.3.
-In Section 3, we prove Theorem 1.6 and show that the range $p\leq 2$ in the Bourin–Lee unitary-orbit inequality (Theorem 1.4) is optimal by exhibiting an explicit $2\times 2$ obstruction for every $p>2$.
-Section 4 establishes Theorem 1.10 via a compact-operator construction that forces the failure of the corresponding singular-value inequality for all $p>2$, and hence confirms the optimality of (1.3).
-In Section 5, we resolve the Thompson-type triangle problem for symmetric moduli (Theorem 1.12).
-Finally, Sections 68 develop isometry-orbit refinements of an Euler operator identity (Theorems 1.15, 1.16, 1.17), derive Clarkson–McCarthy type inequalities, and draw further consequences for unitarily invariant norms and singular values.
+[…]
+Write $\widetilde{U}=\begin{pmatrix}U_{11}&U_{12}\\ U_{21}&U_{22}\end{pmatrix}$. From
+$\widetilde{A}=\widetilde{U}\,\operatorname{diag}(\left|A\right|,0)\quad\text{and}\quad\widetilde{B}=\widetilde{U}\,\operatorname{diag}(\left|B\right|,0)$
+we obtain
+$U_{21}\left|A\right|=A,\qquad U_{21}\left|B\right|=B,\qquad U_{11}\left|A\right|=0,\qquad U_{11}\left|B\right|=0.$
+Let $P$ be the support projection of $\left|A\right|+\left|B\right|$ on $\mathcal{H}$ (equivalently, the orthogonal projection onto $\overline{\operatorname{Ran}(\left|A\right|+\left|B\right|)}$).
+Then $P\left|A\right|=\left|A\right|$ and $P\left|B\right|=\left|B\right|$, and moreover $U_{11}P=0$ since $U_{11}$ vanishes on $\operatorname{Ran}(\left|A\right|)$ and $\operatorname{Ran}(\left|B\right|)$.
+Set $W:=U_{21}P\in\mathcal{B}(\mathcal{H},\mathcal{K})$.
+Then
+$A=W\left|A\right|,\qquad B=W\left|B\right|.$
+[…]
+$\widetilde{U}\widetilde{P}=\begin{pmatrix}U_{11}P&0\\ U_{21}P&0\end{pmatrix}=\begin{pmatrix}0&0\\ W&0\end{pmatrix}$,
+it follows that $W^{*}W=P$, and therefore $W$ is a partial isometry. ∎
+Next, we prove Theorem 1.12.
+Proof of Theorem 1.12.
+Clearly,
+$\left|\Gamma(A)\right|=\left|A\right|_{\mathrm{qsym}},\qquad\left|\Gamma(B)\right|=\left|B\right|_{\mathrm{qsym}},\qquad\left|\Gamma(A+B)\right|=\left|A+B\right|_{\mathrm{qsym}},$
+and
+$\Gamma(A+B)=\Gamma(A)+\Gamma(B).$
+Thus the assumption
+$\left|A+B\right|_{\mathrm{qsym}}=\left|A\right|_{\mathrm{qsym}}+\left|B\right|_{\mathrm{qsym}}$
+is equivalent to
+$\left|\Gamma(A)+\Gamma(B)\right|=\left|\Gamma(A)\right|+\left|\Gamma(B)\right|.$
+Applying Proposition 4.3 to the operators $\Gamma(A),\Gamma(B)\in\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H}),$ we obtain a partial isometry $W\in\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H})$ such that
+$\Gamma(A)=W\,\left|\Gamma(A)\right|=W\,\left|A\right|_{\mathrm{qsym}},\qquad\Gamma(B)=W\,\left|\Gamma(B)\right|=W\,\left|B\right|_{\mathrm{qsym}}.$
+Remark 4.4 (Block form).
+Write $W=\binom{W_{1}}{W_{2}}$ with $W_{1},W_{2}\in\mathcal{B}(\mathcal{H})$. Then $\Gamma(A)=W\left|A\right|_{\mathrm{qsym}}$ and $\Gamma(B)=W\left|B\right|_{\mathrm{qsym}}$ are equivalent to
+$A=\sqrt{2}\,W_{1}\,\left|A\right|_{\mathrm{qsym}},\quad A^{*}=\sqrt{2}\,W_{2}\,\left|A\right|_{\mathrm{qsym}},\qquad B=\sqrt{2}\,W_{1}\,\left|B\right|_{\mathrm{qsym}},\quad B^{*}=\sqrt{2}\,W_{2}\,\left|B\right|_{\mathrm{qsym}}.$
+Thus the equality $\left|A+B\right|_{\mathrm{qsym}}=\left|A\right|_{\mathrm{qsym}}+\left|B\right|_{\mathrm{qsym}}$ forces the same pair $(W_{1},W_{2})$ to control both $(A,A^{*})$ and $(B,B^{*})$ through their quadratic symmetric moduli.
+Remark 4.5 (Why one should not expect $A=U\left|A\right|_{\mathrm{qsym}}$ in general).
+Even for a single operator $A$, a factorization $A=U\left|A\right|_{\mathrm{qsym}}$ with $U\in\mathcal{B}(\mathcal{H})$ a partial isometry is generally impossible unless $A$ is very special.
+For example, if $\mathcal{H}=\mathbb{C}^{n}$ and $A=U\left|A\right|_{\mathrm{qsym}}$, then
+$A^{*}A=\left|A\right|_{\mathrm{qsym}}\,U^{*}U\,\left|A\right|_{\mathrm{qsym}}\leq\left|A\right|_{\mathrm{qsym}}^{2}=\frac{A^{*}A+AA^{*}}{2},$
+so $A^{*}A\leq AA^{*}$.
+Taking traces forces $A^{*}A=AA^{*}$, hence $A$ is normal.
+Therefore non-normal matrices (e.g.
+a Jordan block) cannot satisfy $A=U\left|A\right|_{\mathrm{qsym}}$.
+This explains why the natural equality theorem for the quadratic symmetric modulus is formulated for the lifted operators $\Gamma(A)$.
+5. Concave-function subadditivity for the quadratic symmetric modulus
+We begin by proving the concave-function subadditivity theorem, namely Theorem 1.13.
+Proof of Theorem 1.13.
+Since $f:[0,\infty)\to[0,\infty)$ is concave, it is nondecreasing on $0,\infty)$. By Theorem 1.9, there exist unitaries $U_{0},V_{0}\in\mathbb{M}_{n}$ such that
+$\left|A+B\right|_{\mathrm{qsym}}\leq U_{0}\,\left|A\right|_{\mathrm{qsym}}\,U_{0}^{*}+V_{0}\,\left|B\right|_{\mathrm{qsym}}\,V_{0}^{*}.$
+Set
+$X:=\left|A+B\right|_{\mathrm{qsym}},\qquad Y:=U_{0}\,\left|A\right|_{\mathrm{qsym}}\,U_{0}^{*}+V_{0}\,\left|B\right|_{\mathrm{qsym}}\,V_{0}^{*}.$
+Then $0\leq X\leq Y$, and hence
+$\lambda_{j}(X)\leq\lambda_{j}(Y),\qquad 1\leq j\leq n.$
+Since $f$ is nondecreasing,
+$\lambda_{j}\bigl(f(X)\bigr)=f\bigl(\lambda_{j}(X)\bigr)\leq f\bigl(\lambda_{j}(Y)\bigr)=\lambda_{j}\bigl(f(Y)\bigr),\qquad 1\leq j\leq n.$
+Therefore there exists a unitary $W\in\mathbb{M}_{n}$ such that
+$f(X)\leq W\,f(Y)\,W^{*}.$
+Now apply Theorem 1.5 to the positive matrices $U_{0}\,\left|A\right|_{\mathrm{qsym}}\,U_{0}^{*}$ and $V_{0}\,\left|B\right|_{\mathrm{qsym}}\,V_{0}^{*}$.
+Then there exist unitaries $W_{1},W_{2}\in\mathbb{M}_{n}$ such that
+$f(Y)$ $\leq W_{1}\,f\!\bigl(U_{0}\,\left|A\right|_{\mathrm{qsym}}\,U_{0}^{*}\bigr)\,W_{1}^{*}+W_{2}\,f\!\bigl(V_{0}\,\left|B\right|_{\mathrm{qsym}}\,V_{0}^{*}\bigr)\,W_{2}^{*}$
+$=W_{1}U_{0}\,f\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)\,U_{0}^{*}W_{1}^{*}+W_{2}V_{0}\,f\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)\,V_{0}^{*}W_{2}^{*}.$
+Combining the last two inequalities and setting
+$U:=WW_{1}U_{0},\qquad V:=WW_{2}V_{0},$
+we obtain
+$f\bigl(\left|A+B\right|_{\mathrm{qsym}}\bigr)\leq Uf\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)U^{*}+Vf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)V^{*}.$
+Next, we prove Corollary 1.16.
+Proof of Corollary 1.16.
+Apply Theorem 1.13 to the pair ABA-B and $B$. Then there exist unitaries $W,T\in\mathbb{M}_{n}$ such that
+$f\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)=f\bigl(\left|(A-B)+B\right|_{\mathrm{qsym}}\bigr)\leq Wf\bigl(\left|A-B\right|_{\mathrm{qsym}}\bigr)W^{*}+Tf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)T^{*}.$
+Conjugating by $W^{*}$, we obtain
+$W^{*}f\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)W\leq f\bigl(\left|A-B\right|_{\mathrm{qsym}}\bigr)+W^{*}Tf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)T^{*}W.$
+Hence
+$W^{*}f\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)W-W^{*}Tf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)T^{*}W\leq f\bigl(\left|A-B\right|_{\mathrm{qsym}}\bigr).$
+Therefore, setting
+$U:=W^{*},\qquad V:=W^{*}T,$
+we get
+$Uf\bigl(\left|A\right|_{\mathrm{qsym}}\bigr)U^{*}-Vf\bigl(\left|B\right|_{\mathrm{qsym}}\bigr)V^{*}\leq f\bigl(\left|A-B\right|_{\mathrm{qsym}}\bigr),$
+as desired. ∎
+6. Condition-number estimates for symmetric moduli
+First, we construct two $2\times 2$ matrices to illustrate the sharpness of the coefficient of (1.6).
+Sharpness of the coefficient of (1.6).
+To see that the coefficient in (1.6) is best possible, it is enough to consider the case $m=2$ and $n=2$.
+Write $t=\sqrt{\omega},$ in particular, $t\geq 1$.
+Define
+$A_{1}=\begin{pmatrix}t&0\\[2.84526pt] 0&t^{-1}\end{pmatrix},\qquad R=\frac{1}{t^{2}+1}\begin{pmatrix}-2t&t^{2}-1\\ 1-t^{2}&-2t\end{pmatrix},\qquad J=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$,
+and set
+$A_{2}:=J\begin{pmatrix}t^{-1}&0\\ 0&t\end{pmatrix}R.$
+Since both $J$ and $R$ are orthogonal, the singular values of $A_{2}$ are exactly $t$ and $t^{-1}$. The same is clearly true for $A_{1}$. Hence
+$\kappa(A_{1})=\kappa(A_{2})=\frac{t}{t^{-1}}=t^{2}=\omega.$
+Now let $x=\binom{1}{-t}$. Since $A_{1}$ is positive definite, we have $|A_{1}|=A_{1}$. Moreover,
+$|A_{2}|=R^{T}\begin{pmatrix}t^{-1}&0\\ 0&t\end{pmatrix}R.$
+A direct calculation then shows that
+$\bigl(|A_{1}|+|A_{2}|\bigr)x=\frac{4t}{t^{2}+1}\,x.$ (6.1)
+On the other hand,
+$A_{1}+A_{2}=\frac{2}{t^{2}+1}\begin{pmatrix}t&-t^{2}\\ -1&t\end{pmatrix}$,
+and therefore
+$(A_{1}+A_{2})^{*}(A_{1}+A_{2})=\frac{4}{t^{2}+1}\begin{pmatrix}1&-t\\ -t&t^{2}\end{pmatrix}$.
+If we denote
+$P=\begin{pmatrix}1&-t\\ -t&t^{2}\end{pmatrix}$,
+then $P^{2}=(t^{2}+1)P$, and thus
+$|A_{1}+A_{2}|=\frac{2}{t^{2}+1}P.$
+Since $Px=(t^{2}+1)x$, it follows that
+$|A_{1}+A_{2}|x=2x.$ (6.2)
+Combining the identities (6.1) and (6.2), we obtain
+$|A_{1}+A_{2}|x=\frac{t^{2}+1}{2t}\,\bigl(|A_{1}|+|A_{2}|\bigr)x=\frac{\omega+1}{2\sqrt{\omega}}\,\bigl(|A_{1}|+|A_{2}|\bigr)x.$
+Consequently, if one replaced the coefficient $\frac{\omega+1}{2\sqrt{\omega}}$ in (1.6) by any smaller constant $c$, then
+$x^{*}\Bigl(c\bigl(|A_{1}|+|A_{2}|\bigr)-|A_{1}+A_{2}|\Bigr)x<0,$
+which is impossible if the corresponding operator inequality were true. This proves that the coefficient $(\omega+1)/(2\sqrt{\omega})$ is sharp.∎
+Before giving a proof of Theorem 1.17, we need some lemmas.
+Lemma 6.1.
+Let $X\in\mathcal{B}(\mathcal{H})$ be positive and assume that
+$\alpha I\leq X\leq\beta I\qquad(0<\alpha\leq\beta).$
+If $K\in\mathcal{B}(\mathcal{H})$ is a contraction, i.e. $\|K\|_{\infty}\leq 1$, then
+$\operatorname{Re}(KX)\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}\,X.$
+Proof.
+Define $Y:=X^{-1/2}KX^{1/2}.$ Then
+$X^{1/2}YX^{1/2}=KX,\qquad X^{1/2}Y^{*}X^{1/2}=XK^{*},$
+and hence
+$\operatorname{Re}(KX)=X^{1/2}\,\operatorname{Re}(Y)\,X^{1/2}.$
+Therefore it suffices to show that
+$\operatorname{Re}(Y)\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}\,I.$
+Let $x\in\mathcal{H}$ be a unit vector. Then
+$\langle\operatorname{Re}(Y)x,x\rangle=\operatorname{Re}\langle Yx,x\rangle\leq|\langle Yx,x\rangle|.$
+By the Cauchy–Schwarz inequality and the assumption $\|K\|_{\infty}\leq 1$,
+$|\langle Yx,x\rangle|=\bigl|\langle KX^{1/2}x,X^{-1/2}x\rangle\bigr|\leq\|KX^{1/2}x\|_{2}\,\|X^{-1/2}x\|_{2}\leq\|X^{1/2}x\|_{2}\,\|X^{-1/2}x\|_{2}.$
+On the other hand, for every $t\in[\alpha,\beta]$ one has
+$(t-\alpha)(\beta-t)\geq 0,$
+which is equivalent to
+$t+\alpha\beta\,t^{-1}\leq\alpha+\beta.$
+By functional calculus,
+$X+\alpha\beta\,X^{-1}\leq(\alpha+\beta)I.$
+Taking inner products with $x$ gives
+$\langle Xx,x\rangle+\alpha\beta\,\langle X^{-1}x,x\rangle\leq\alpha+\beta.$
+Using the arithmetic–geometric mean inequality,
+$2\sqrt{\alpha\beta}\,\sqrt{\langle Xx,x\rangle\langle X^{-1}x,x\rangle}\leq\langle Xx,x\rangle+\alpha\beta\,\langle X^{-1}x,x\rangle\leq\alpha+\beta.$
+Hence
+$\|X^{1/2}x\|_{2}\,\|X^{-1/2}x\|_{2}=\sqrt{\langle Xx,x\rangle\langle X^{-1}x,x\rangle}\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}.$
+Therefore
+$\langle\operatorname{Re}(Y)x,x\rangle\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}.$
+Since this holds for every unit vector $x$, we obtain
+$\operatorname{Re}(Y)\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}\,I.$
+Multiplying on both sides by $X^{1/2}$ yields
+$\operatorname{Re}(KX)\leq\frac{\alpha+\beta}{2\sqrt{\alpha\beta}}\,X.$
+Proposition 6.2.
+Let $A_{1},\dots,A_{m}\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Assume that for each $i$ there exist real numbers $0<\alpha_{i}\leq\beta_{i}$ such that
+$\alpha_{i}I\leq\left|A_{i}\right|\leq\beta_{i}I.$
+Then
+$\left|\sum_{i=1}^{m}A_{i}\right|\leq\sum_{i=1}^{m}\frac{\alpha_{i}+\beta_{i}}{2\sqrt{\alpha_{i}\beta_{i}}}\,\left|A_{i}\right|.$
+Proof.
+Set $S:=\sum_{i=1}^{m}A_{i}$ and write the polar decomposition
+$S=W\left|S\right|,$
+where $W\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is a partial isometry. Then
+$W^{*}S=W^{*}W\left|S\right|=\left|S\right|.$
+For each $i$, write $A_{i}=U_{i}\left|A_{i}\right|$ for the polar decomposition of $A_{i}$.
+Since $\left|A_{i}\right|\geq\alpha_{i}I>0$, the support projection of $\left|A_{i}\right|$ is $I$, and therefore $U_{i}^{*}U_{i}=I$; that is, each $U_{i}$ is an isometry.
+In particular, $W^{*}U_{i}$ is a contraction.
+Now
+$\left|S\right|=W^{*}S=\sum_{i=1}^{m}W^{*}U_{i}\left|A_{i}\right|.$
+Since $\left|S\right|$ is self-adjoint, taking real parts gives
+$\left|S\right|=\sum_{i=1}^{m}\operatorname{Re}\,\!\bigl(W^{*}U_{i}\left|A_{i}\right|\bigr).$
+Applying Lemma 6.1 to
+$X=\left|A_{i}\right|,\qquad K=W^{*}U_{i},$
+we obtain
+$\operatorname{Re}\!\bigl(W^{*}U_{i}\left|A_{i}\right|\bigr)\leq\frac{\alpha_{i}+\beta_{i}}{2\sqrt{\alpha_{i}\beta_{i}}}\,\left|A_{i}\right|$
+for every $i$. Summing these inequalities yields
+$\left|S\right|\leq\sum_{i=1}^{m}\frac{\alpha_{i}+\beta_{i}}{2\sqrt{\alpha_{i}\beta_{i}}}\,\left|A_{i}\right|.$
+Lemma 6.3.
+Let $c(\kappa):=\frac{\kappa+1}{2\sqrt{\kappa}}$ with $\kappa\geq 1.$ Then for every $a<c(\kappa)$, there exist invertible $2\times 2$ matrices $A,B$ such that
+$\kappa(A)=\kappa,\qquad\kappa(B)=1,$
+but
+$|A+B|\nleq a\,|A|+|B|.$
+In particular, the coefficient $c(\kappa)$ in front of $|A|$ is sharp even when the coefficient in front of $|B|$ is fixed at $1=c(1)$.
+Proof.
+If $\kappa=1$, then $c(1)=1$. Let $a<1$, take
+$A=I_{2},\qquad B=\varepsilon I_{2}$
+with $0<\varepsilon<1-a$. Then $\kappa(A)=\kappa(B)=1$, while
+$|A+B|=(1+\varepsilon)I_{2}\nleq(a+\varepsilon)I_{2}=a|A|+|B|.$
+So the claim holds when $\kappa=1$.
+Assume now that $\kappa>1$, and write $t:=\sqrt{\kappa}>1$. Set
+$m:=\frac{t+t^{-1}}{2}=c(\kappa),\qquad d:=\frac{t-t^{-1}}{2}.$
+Define
+$A:=\begin{pmatrix}0&t^{-1}\\ t&0\end{pmatrix},\qquad B_{\mu}:=\begin{pmatrix}\mu&d\\ -d&\mu\end{pmatrix},\qquad\mu\geq m.$
+Then
+$A^{*}A=\begin{pmatrix}t^{2}&0\\ 0&t^{-2}\end{pmatrix},\qquad|A|=\begin{pmatrix}t&0\\ 0&t^{-1}\end{pmatrix}$,
+so $\kappa(A)=t^{2}=\kappa$.
+Also,
+$B_{\mu}^{*}B_{\mu}=(\mu^{2}+d^{2})I_{2},$
+hence
+$|B_{\mu}|=\sqrt{\mu^{2}+d^{2}}\,I_{2},\qquad\kappa(B_{\mu})=1.$
+Moreover,
+$A+B_{\mu}=\begin{pmatrix}\mu&m\\ m&\mu\end{pmatrix}\geq 0\qquad(\mu\geq m),$
+so
+$|A+B_{\mu}|=A+B_{\mu}=\begin{pmatrix}\mu&m\\ m&\mu\end{pmatrix}$.
+Suppose that
+$|A+B_{\mu}|\leq a\,|A|+|B_{\mu}|.$
+Writing
+$r_{\mu}:=\sqrt{\mu^{2}+d^{2}},\qquad\delta_{\mu}:=r_{\mu}-\mu>0,$
+this becomes
+$\begin{pmatrix}at+\delta_{\mu}&-m\\ -m&at^{-1}+\delta_{\mu}\end{pmatrix}\geq 0.$
+Hence its determinant must be nonnegative:
+$(at+\delta_{\mu})(at^{-1}+\delta_{\mu})-m^{2}\geq 0.$
+Since $tt^{-1}=1$ and $t+t^{-1}=2m$, this is
+$a^{2}+2m\delta_{\mu}\,a+\delta_{\mu}^{2}-m^{2}\geq 0.$
+Therefore
+$a\geq\rho_{\mu}:=-m\delta_{\mu}+\sqrt{m^{2}+(m^{2}-1)\delta_{\mu}^{2}}.$
+Now $\delta_{\mu}=r_{\mu}-\mu\to 0$ as $\mu\to\infty$, so
+$\rho_{\mu}\longrightarrow m=c(\kappa).$
+Given any $a<c(\kappa)$, we can choose $\mu$ sufficiently large so that
+$a<\rho_{\mu}.$
+For this choice of $\mu$, the inequality
+$|A+B_{\mu}|\leq a\,|A|+|B_{\mu}|$
+cannot hold. This proves the lemma. ∎
+Next, we give a proof of Theorem 1.17 based on Proposition 6.2 and Lemma 6.3.
+Proof of Theorem 1.17.
+For each $i$, set
+$\alpha_{i}:=\|A_{i}^{-1}\|_{\infty}^{-1},\qquad\beta_{i}:=\|A_{i}\|_{\infty}.$
+Then
+$\alpha_{i}I\leq\left|A_{i}\right|\leq\beta_{i}I,\qquad\frac{\beta_{i}}{\alpha_{i}}=\|A_{i}\|_{\infty}\,\|A_{i}^{-1}\|_{\infty}=\kappa_{i}.$
+Also, $\alpha_{i}I\leq\left|A_{i}^{*}\right|\leq\beta_{i}I,$ because $\|A_{i}^{*}\|_{\infty}=\|A_{i}\|_{\infty}$ and $\|(A_{i}^{*})^{-1}\|_{\infty}=\|A_{i}^{-1}\|_{\infty}$.
+By Proposition 6.2,
+$\left|A_{1}+\cdots+A_{m}\right|\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\left|A_{1}\right|+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\left|A_{m}\right|,$
+and similarly,
+$\left|A_{1}^{*}+\cdots+A_{m}^{*}\right|\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\left|A_{1}^{*}\right|+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\left|A_{m}^{*}\right|.$
+Thus,
+$\left|A_{1}+\cdots+A_{m}\right|_{\mathrm{sym}}\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\,\left|A_{1}\right|_{\mathrm{sym}}+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\,\left|A_{m}\right|_{\mathrm{sym}}.$
+Next, recall the lifting
+$\Gamma:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H}),\qquad\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}$.
+Then $\Gamma$ is linear and, for each $i$,
+$\left|\Gamma(A_{i})\right|^{2}=\left|A_{i}\right|_{\mathrm{qsym}}^{2}=\frac{\left|A_{i}\right|^{2}+\left|A_{i}^{*}\right|^{2}}{2}.$
+From
+$\alpha_{i}I\leq\left|A_{i}\right|,\ \left|A_{i}^{*}\right|\leq\beta_{i}I,$
+it follows that
+$\alpha_{i}I\leq\left|\Gamma(A_{i})\right|\leq\beta_{i}I.$
+Applying Proposition 6.2 to
+$\Gamma(A_{1}),\dots,\Gamma(A_{m})\in\mathcal{B}(\mathcal{H},\mathcal{H}\oplus\mathcal{H}),$
+yields
+$\left|\Gamma(A_{1}+\cdots+A_{m})\right|\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\,\left|\Gamma(A_{1})\right|+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\,\left|\Gamma(A_{m})\right|.$
+That is,
+$\left|A_{1}+\cdots+A_{m}\right|_{\mathrm{qsym}}\leq\frac{\kappa_{1}+1}{2\sqrt{\kappa_{1}}}\,\left|A_{1}\right|_{\mathrm{qsym}}+\cdots+\frac{\kappa_{m}+1}{2\sqrt{\kappa_{m}}}\,\left|A_{m}\right|_{\mathrm{qsym}}.$
+This completes the proof.
+We now prove the sharpness statement in Theorem 1.17.
+Fix $\kappa\geq 1$, and let $c(\kappa)=(\kappa+1)/(2\sqrt{\kappa})$.
+By Lemma 6.3, for every $a<c(\kappa)$ there exist invertible $2\times 2$ matrices $A,B$ such that
+$\kappa(A)=\kappa,\qquad\kappa(B)=1,$
+but
+$|A+B|\nleq a\,|A|+|B|.$
+Thus the coefficient $c(\kappa)$ is already sharp in the two-variable inequality for the usual modulus.
+To transfer this sharpness to the arithmetic and quadratic symmetric moduli, consider the self-adjoint block matrices
+$\widehat{A}:=\begin{pmatrix}0&A\\ A^{*}&0\end{pmatrix},\qquad\widehat{B}:=\begin{pmatrix}0&B\\ B^{*}&0\end{pmatrix}$.
+Then
+$\widehat{A}=\widehat{A}^{*},\qquad\widehat{B}=\widehat{B}^{*},$
+and
+$\kappa(\widehat{A})=\kappa(A)=\kappa,\qquad\kappa(\widehat{B})=\kappa(B)=1.$
+Since $\widehat{A}$ and $\widehat{B}$ are self-adjoint, all three moduli coincide:
+$\left|\widehat{A}\right|_{\mathrm{sym}}=\left|\widehat{A}\right|_{\mathrm{qsym}}=|\widehat{A}|=\begin{pmatrix}|A^{*}|&0\\ 0&|A|\end{pmatrix},\qquad\left|\widehat{B}\right|_{\mathrm{sym}}=\left|\widehat{B}\right|_{\mathrm{qsym}}=|\widehat{B}|=\begin{pmatrix}|B^{*}|&0\\ 0&|B|\end{pmatrix}$,
+and similarly
+$\left|\widehat{A}+\widehat{B}\right|_{\mathrm{sym}}=\left|\widehat{A}+\widehat{B}\right|_{\mathrm{qsym}}=|\widehat{A}+\widehat{B}|=\begin{pmatrix}|A^{*}+B^{*}|&0\\ 0&|A+B|\end{pmatrix}$.
+Suppose, for contradiction, that one could replace the coefficient $c(\kappa)$ in front of the first term by some smaller constant $a<c(\kappa)$ in the arithmetic symmetric modulus inequality.
+Applying that putative inequality to $\widehat{A},\widehat{B}$ would give
+$\left|\widehat{A}+\widehat{B}\right|_{\mathrm{sym}}\leq a\,\left|\widehat{A}\right|_{\mathrm{sym}}+\left|\widehat{B}\right|_{\mathrm{sym}}.$
+Comparing the lower-right block yields
+$|A+B|\leq a\,|A|+|B|,$
+contrary to Lemma 6.3. Hence the coefficient $c(\kappa)$ is sharp for $\left|\cdot\right|_{\mathrm{sym}}$.
+Exactly the same block comparison shows that $c(\kappa)$ is also sharp for $\left|\cdot\right|_{\mathrm{qsym}}$.
+Consequently, the coefficient function
+$c(\kappa)=\frac{\kappa+1}{2\sqrt{\kappa}}$
+is individually sharp for each of the three moduli
+$|\cdot|,\qquad|\cdot|_{\mathrm{sym}},\qquad|\cdot|_{\mathrm{qsym}}.$
+This completes the proof. ∎
+7. Clarkson–McCarthy type inequalities for the quadratic symmetric modulus
+Proof of Theorem 1.18.
+Recall that for every operator $T$,
+$|T|_{\mathrm{qsym}}^{2}=\frac{|T|^{2}+|T^{*}|^{2}}{2}=\frac{T^{*}T+TT^{*}}{2}.$
+For each $1\leq i\leq m$, set $T_{i}:=\sum_{j=1}^{n}u_{ij}A_{j}.$ Then
+$\sum_{i=1}^{m}|T_{i}|_{\mathrm{qsym}}^{2}=\frac{1}{2}\sum_{i=1}^{m}\bigl(T_{i}^{*}T_{i}+T_{i}T_{i}^{*}\bigr).$
+Assume first that $U^{*}U=I_{n}$. Then
+$\sum_{i=1}^{m}T_{i}^{*}T_{i}=\sum_{i=1}^{m}\left(\sum_{j=1}^{n}\overline{u_{ij}}A_{j}^{*}\right)\left(\sum_{k=1}^{n}u_{ik}A_{k}\right)=\sum_{j,k=1}^{n}\left(\sum_{i=1}^{m}\overline{u_{ij}}u_{ik}\right)A_{j}^{*}A_{k}.$
+Since
+$\sum_{i=1}^{m}\overline{u_{ij}}u_{ik}=(U^{*}U)_{jk}=\delta_{jk},$
+we obtain
+$\sum_{i=1}^{m}T_{i}^{*}T_{i}=\sum_{j=1}^{n}A_{j}^{*}A_{j}.$
+Similarly,
+$\sum_{i=1}^{m}T_{i}T_{i}^{*}=\sum_{i=1}^{m}\left(\sum_{j=1}^{n}u_{ij}A_{j}\right)\left(\sum_{k=1}^{n}\overline{u_{ik}}A_{k}^{*}\right)=\sum_{j,k=1}^{n}\left(\sum_{i=1}^{m}u_{ij}\overline{u_{ik}}\right)A_{j}A_{k}^{*}=\sum_{j=1}^{n}A_{j}A_{j}^{*}.$
+Therefore
+$\sum_{i=1}^{m}|T_{i}|_{\mathrm{qsym}}^{2}=\frac{1}{2}\sum_{j=1}^{n}\bigl(A_{j}^{*}A_{j}+A_{j}A_{j}^{*}\bigr)=\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}.$
+Conversely, assume that
+$\sum_{i=1}^{m}\left|\sum_{j=1}^{n}u_{ij}A_{j}\right|_{\mathrm{qsym}}^{2}=\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}$
+holds for all $A_{1},\ldots,A_{n}\in\mathbb{M}_{N}$. Let $P$ be any rank-one projection in $\mathbb{M}_{N}$, and choose arbitrary scalars $a_{1},\ldots,a_{n}\in\mathbb{C}$. Put
+$A_{j}:=a_{j}P\qquad(1\leq j\leq n).$
+Since $P=P^{*}=P^{2}$, we have
+$\left|\left(\sum_{j=1}^{n}u_{ij}a_{j}\right)P\right|_{\mathrm{qsym}}^{2}=\left|\sum_{j=1}^{n}u_{ij}a_{j}\right|^{2}P,\qquad|a_{j}P|_{\mathrm{qsym}}^{2}=|a_{j}|^{2}P.$
+Hence the assumed identity becomes
+$\sum_{i=1}^{m}\left|\sum_{j=1}^{n}u_{ij}a_{j}\right|^{2}P=\sum_{j=1}^{n}|a_{j}|^{2}P.$
+Since $P\neq 0$, it follows that
+$\sum_{i=1}^{m}\left|\sum_{j=1}^{n}u_{ij}a_{j}\right|^{2}=\sum_{j=1}^{n}|a_{j}|^{2}\qquad\text{for all }(a_{1},\ldots,a_{n})\in\mathbb{C}^{n}.$
+This is exactly the condition that $U$ is an isometry, namely, $U^{*}U=I_{n}.$ The proof is complete. ∎
+We now turn to a unitary-orbit version of Jensen’s inequality for monotone convex/concave functions.
+These orbit inequalities yield matrix-order dominations, and hence deduce Schatten $p$-norm consequences.
+The following unitary-orbit inequalities on monotone concave/convex functions will be used in the sequel.
+Lemma 7.1 (1) the first inequality is obtained by a straightforward induction on $n$ from the corresponding two-variable inequality [7, Corollary 3.2].
+Lemma 7.1 (2) the first inequality follows similarly from [7, Theorem 3.1] (equivalently, [3, Theorem 2.1]).
+Lemma 7.1 (1) the second inequality follows from the isometric-column Jensen inequality [7, Corollary 2.4] by choosing $Z_{i}=\sqrt{\alpha_{i}}\,I$, so that $\sum_{i=1}^{n}Z_{i}^{*}Z_{i}=I$ and $\sum_{i=1}^{n}Z_{i}^{*}A_{i}Z_{i}=\sum_{i=1}^{n}\alpha_{i}A_{i}$.
+Similarly, Lemma 7.1 (2) the second inequality follows from the same result, using the fact that the inequality in [7, Corollary 2.4] reverses for concave functions.
+Lemma 7.1.
+Let $A_{1},\dots,A_{n}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{n}$ be positive real numbers such that $\sum_{j=1}^{n}\alpha_{j}=1$. Then
+  1. (1)
+    for every monotone convex function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\leq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
+    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\geq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
+    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\leq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*};$
+  2. (2)
+    for every monotone concave function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\geq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
+    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\leq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
+    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\geq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*}.$
+Proof of Theorem 1.19.
+By Theorem 1.18, the assumption implies that
+$\sum_{i=1}^{m}|B_{i}|_{\mathrm{qsym}}^{2}=\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}.$
+Case 1: $p\geq 2$. Take $f(t)=t^{p/2}$ in Lemma 7.1 (1). Then there exist unitaries $U_{1},\ldots,U_{m}$ and $W$ such that
+$\sum_{i=1}^{m}U_{i}|B_{i}|_{\mathrm{qsym}}^{p}U_{i}^{*}$ $\leq\left(\sum_{i=1}^{m}|B_{i}|_{\mathrm{qsym}}^{2}\right)^{p/2}$
+$=\left(\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}\right)^{p/2}$
+$=n^{p/2}\left(\frac{1}{n}\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}\right)^{p/2}$
+$\leq n^{\frac{p}{2}-1}W\left(\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{p}\right)W^{*}.$
+Conjugating by $W^{*}$ and replacing each $U_{i}$ by $W^{*}U_{i}$, we obtain
+$\sum_{i=1}^{m}U_{i}|B_{i}|_{\mathrm{qsym}}^{p}U_{i}^{*}\leq n^{\frac{p}{2}-1}\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{p}.$
+The second inequality is proved in the same way, interchanging the roles of $(A_{1},\ldots,A_{n})$ and $(B_{1},\ldots,B_{m})$.
+Case 2: $0<p\leq 2$. The argument is the same, except that the inequalities in Lemma 7.1 reverse. ∎
+Next, we illustrate a Jensen-type principle for unitarily invariant norms, which allows us to convert the square-sum identity (1.9) into $p$-power estimates via the function $t\mapsto t^{p/2}$.
+We will use the following Jensen-type inequalities for unitarily invariant norms.
+Lemma 7.2 ([13, Lemma 2.1]).
+Let $m\geq 1$. Let $A_{1},\dots,A_{m}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{m}$ be positive real numbers such that $\sum_{j=1}^{m}\alpha_{j}=1$. Then
+  1. (1)
+    for every convex function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
+[…]
+[…]
+  2. (2)
+    for every concave function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
+[…]
+[…]
+Proof of Theorem 1.21.
+By Theorem 1.18, the condition $(B_{1},\ldots,B_{m})^{T}=U(A_{1},\ldots,A_{n})^{T}$ for some isometry $U=(u_{ij})$ implies that
+$\sum_{i=1}^{m}\left|B_{i}\right|_{\mathrm{qsym}}^{2}=\sum_{j=1}^{n}|A_{j}|_{\mathrm{qsym}}^{2}.$
+Case 1: $p\geq 2$. Take $f(t)=t^{\frac{p}{2}}$ in Lemma 7.2 (1), we have
+[…]
+[…]
+[…]
+Similarly, we have
+[…]
+Case 2: $0<p\leq 2$. The argument is analogous to that for the case $p\geq 2$, except that all the inequalities are reversed. ∎
 8. Proof of Theorem 1.27
@@
 9. Proof of Theorem 1.30
+In this section, we prove Theorem 1.30.
 9.1. Trace necessary condition
@@
 Compute
-$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c\\ -s\end{pmatrix}(c,\,-s)=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
+$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c&-s\\ 0&0\end{pmatrix}\begin{pmatrix}c&0\\ -s&0\end{pmatrix}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
 Hence $|Z_{\theta}|=(Z_{\theta}^{*}Z_{\theta})^{1/2}=P$. Similarly,
@@
 For $\operatorname{Im}\,Z_{\theta}$, the eigenvalues are $\pm s/2$, hence $\operatorname{Tr}|\operatorname{Im}\,Z_{\theta}|=s$. Adding yields (9.4). ∎
-Combining (3.1), Claims 3.1 and 3.2, we see that if (1.2) were to hold for $Z_{\theta}$, then necessarily
+Combining (9.1), Claims 9.1 and 9.2, we see that if (1.10) were to hold for $Z_{\theta}$, then necessarily
 $\left(\frac{1+c}{2}\right)^{1/p}+\left(\frac{1-c}{2}\right)^{1/p}\leq 1+s.$ (9.5)
@@
 If (1.10) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (9.1).
-Claims 3.1 and 3.2 reduce (3.1) to (3.5), which fails for all sufficiently small $\theta$ by Claim 3.3.
+Claims 9.1 and 9.2 reduce (9.1) to (9.5), which fails for all sufficiently small $\theta$ by Claim 9.3.
 Hence (1.10) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
 Remark 10.1.
-The construction is intentionally minimal and answers the existence question in Question 1.9.
+The construction is intentionally minimal and answers the existence question in Question 1.33.
 Here $Z$ has rank $2$, and the operator $\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}$ is diagonal and easy to read off explicitly.
 The strict inequality is obtained by selecting indices $(j,k)=(2,0)$ for which the corresponding singular value of $\operatorname{Re}Z$ vanishes at level $1+j=3$, while $\operatorname{Im}Z$ contributes only its top singular value $2^{-1/2}$, yet the third singular value on the left-hand side equals $2^{-1/p}>2^{-1/2}$ whenever $p>2$.
-5. Rectangular Thompson inequality and proof of Theorem 1.12
-In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.12 (ii).
-The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$.
-After recording this elementary fact (Lemma 5.1), we apply the classical (square) Thompson inequality to the compressed matrices $W^{*}A$ and $W^{*}B$ and then compare $|W^{*}A|,|W^{*}B|$ with $|A|,|B|$, respectively, to obtain the desired unitary-orbit domination.
-Finally, we use the rectangular Thompson inequality to settle Question 1.11, thereby proving Theorem 1.12.
-Lemma 5.1.
-Let $X\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
-$X=W|X|,\qquad|X|=(X^{*}X)^{1/2},$
-and
-$W^{*}W=\mathrm{supp}(|X|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|X|=|X|,$
-where $\mathrm{supp}(\cdot)$ means the support projection.
-Proof.
-Take a singular value decomposition $X=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
-$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(X).$
-Then
-$X^{*}X=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
-hence
-$|X|=(X^{*}X)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
-Define $W:=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\in\mathbb{M}_{m,n}.$ Clearly, $W^{*}W\leq I_{n},WW^{*}\leq I_{m}$ and
-$W|X|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=X.$
-Moreover,
-$(W^{*}W)|X|=P|X|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|X|.$
-That is, $W^{*}W$ is exactly the support projection of $|X|$. ∎
-Lemma 5.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|X|$.
-This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
-Lemma 5.2 (Rectangular Thompson inequality).
-Let $A,B\in\mathbb{M}_{m,n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
-$|A+B|\ \leq\ U|A|U^{*}+V|B|V^{*}.$ (5.1)
-Proof.
-By using Lemma 5.1, there exists a partial isometry $W$ such that A+B=W|A+B|A+B=W\,|A+B|, $W^{*}W$ is the support projection of $|A+B|$ and $W^{*}W\,|A+B|=|A+B|$.Therefore
-$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
-In particular,
-$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (5.2)
-Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices. Applying Thompson’s inequality (1.1) (the square case) to $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
-$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (5.3)
-Since the left-hand side equals $|A+B|$, it remains to compare $|W^{*}A|$ with $|A|$.
-Since $WW^{*}\leq I_{m}$, we have
-$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
-By the operator monotonicity of $t\mapsto t^{1/2}$ on $0,\infty)$, this implies $|W^{*}A|\leq|A|$. Similarly, $|W^{*}B|\leq|B|$. Conjugating preserves the order, hence
-$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
-Substituting into (5.3) and combining with (5.2) yields (5.1). ∎
-With Lemma 5.2 in hand, we now prove Theorem 1.12.
-Part (i) is verified by a concrete counterexample for the symmetric modulus, while part (ii) follows from an application of Lemma 5.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
-Proof of Theorem 1.12.
-(i) Take
-$X=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad Y=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in M_{2}(\mathbb{R}).$
-A direct computation gives
-$|X|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|Y|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|X+Y|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
-Hence
-$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
-so
-$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e.
-}\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
-If the inequality in Question 1.11 held for some unitaries $U,V$, then taking operator norms would yield
-$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty},$
-a contradiction. This proves (i).
-(ii) Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
-$\Gamma(Z):=\begin{pmatrix}Z\\ Z^{*}\end{pmatrix}$.
-Then
-$\Gamma(Z)^{*}\Gamma(Z)=Z^{*}Z+ZZ^{*},\qquad\text{hence}\qquad|\Gamma(Z)|=\bigl(Z^{*}Z+ZZ^{*}\bigr)^{1/2}=\sqrt{2}\,|Z|_{\mathrm{qsym}}.$
-Moreover, $\Gamma(X+Y)=\Gamma(X)+\Gamma(Y)$. Applying rectangular Thompson inequality (Lemma 5.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
-$|\Gamma(X+Y)|\leq U|\Gamma(X)|U^{*}+V|\Gamma(Y)|V^{*}.$
-Dividing by $\sqrt{2}$ and using $|\Gamma(Z)|=\sqrt{2}\,|Z|_{\mathrm{qsym}}$ gives
-$|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*},$
-as desired. ∎
 11. Euler-type identities and isometry orbits
@@
 We start with a Pythagoras-type theorem for partitioned matrices due to Bourin and Lee [9], which allows us to express the square modulus of a block matrix as a sum of isometric conjugations of the square moduli of its blocks.
-Lemma 6.1 ([11, Theorem 2.1]).
+Lemma 11.1 ([9, Theorem 2.1]).
 Let $m,n\geq 1$ and let $T=[T_{ij}]_{i,j=1}^{m}$ be an $m\times m$ block matrix with $T_{ij}\in\mathbb{M}_{n}$. Then there exist isometries $W_{ij}\in\mathbb{M}_{mn,n}$ such that
@@
 This lemma will serve as a building block for a higher-dimensional isometry decomposition.
-Lemma 6.3 ([9, Lemma 3.4]).
+Lemma 11.3 ([7, Lemma 3.4]).
 Let $\begin{pmatrix}X&Y\\ Y^{*}&Z\end{pmatrix}\in\mathbb{M}_{2n}$ be positive semidefinite, written in $n\times n$ blocks. Then there exist unitary matrices $U,V_{0}\in\mathbb{M}_{2n}$ such that
@@
 Conjugating by $W_{4}^{*}$ and setting $U_{k}=W_{4}^{*}V_{k}$ completes the proof. ∎
-7. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
-In this section, we exploit Euler’s operator identity (1.6) and the associated isometry-/unitary-orbit refinements obtained in Section 6 to derive sharp Clarkson–McCarthy type inequalities.
-We treat two parallel frameworks: first, inequalities for general unitarily invariant norms; second, stronger unitary-orbit dominations, from which Schatten $p$-norm inequalities follow by taking traces.
-Finally, we introduce a vector-valued complex interpolation scheme and deduce a mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimate, which will be used to prove the mixed-exponent Euler-type inequalities later in the paper.
-7.1. Unitarily invariant norm case
-We begin with a Jensen-type principle for unitarily invariant norms, which allows us to convert the square-sum identity (1.6) into $p$-power estimates via the function $t\mapsto t^{p/2}$.
-We will use the following Jensen-type inequalities for unitarily invariant norms.
-Lemma 7.1 ([14, Lemma 2.1]).
-Let $m\geq 1$. Let $A_{1},\dots,A_{m}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{m}$ be positive real numbers such that $\sum_{j=1}^{m}\alpha_{j}=1$. Then
-  1. (1)
-    for every convex function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
-[…]
-[…]
-  2. (2)
-    for every concave function $f:[0,\infty)\to[0,\infty)$ with $f(0)=0$,
-[…]
-[…]
-Proof of Theorem 1.25.
-Set $r:=p/2$ and $f(t):=t^{r}$ on $0,\infty)$. Note that $f(0)=0$.
-Define the positive semidefinite matrices
-$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
-and
-$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
-By Euler’s identity (1.6),
-$X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}=:S.$
-Moreover, $|A+B|^{p}=X_{1}^{r}=f(X_{1})$, etc., and $|A|^{p}=Y_{2}^{r}=f(Y_{2})$, etc.
-Step 1: case 1: $p\geq 2$. That is, $r\geq 1$ Then $f$ is convex.
-Applying Lemma 7.1 (1) the second inequality to $X_{1},X_{2},X_{3}$ yields
-[…]
-Next, applying Lemma 7.1 (1) the first inequality with $n=4$ and $\alpha_{j}=1/4$ to $Y_{1},Y_{2},Y_{3},Y_{4}$ gives
-[…]
-[…]
-[…]
-Hence
-[…]
-Combining the above inequalities yields the desired estimate for $p\geq 2$.
-Step 2: $0<p\leq 2$. That is, $0<r\leq 1$. Then $f$ is concave.
-Applying Lemma 7.1 (2) the second inequality to $X_{1},X_{2},X_{3}$ yields
-[…]
-Next, applying Lemma 7.1 (2) the first inequality with $n=4$ and $\alpha_{j}=1/4$ to $Y_{1},Y_{2},Y_{3},Y_{4}$ gives
-[…]
-Using homogeneity again,
-[…]
-and therefore
-[…]
-[…]
-Step 3: optimality. Taking $A=B=-C\neq 0$, we have $A+B=2A$, $B+C=0$, $C+A=0$, and $A+B+C=A$. Thus
-[…]
-while
-[…]
-Hence equality holds for all $p>0$, and the constant $2^{p-2}$ is best possible. ∎
-7.2. Unitary orbit case
-We next turn to a unitary-orbit version of Jensen’s inequality for monotone convex/concave functions.
-These orbit inequalities yield matrix-order dominations, and hence deduce Schatten $p$-norm consequences.
-The following unitary-orbit inequalities on monotone concave/convex functions will be used in the sequel.
-Lemma 7.2 (1) the first inequality is obtained by a straightforward induction on $n$ from the corresponding two-variable inequality [9, Corollary 3.2].
-Lemma 7.2 (2) the first inequality follows similarly from [9, Theorem 3.1] (equivalently, [2, Theorem 2.1]).
-Lemma 7.2 (1) the second inequality follows from the isometric-column Jensen inequality [9, Corollary 2.4] by choosing $Z_{i}=\sqrt{\alpha_{i}}\,I$, so that $\sum_{i=1}^{n}Z_{i}^{*}Z_{i}=I$ and $\sum_{i=1}^{n}Z_{i}^{*}A_{i}Z_{i}=\sum_{i=1}^{n}\alpha_{i}A_{i}$.
-Similarly, Lemma 7.2 (2) the second inequality follows from the same result, using the fact that the inequality in [9, Corollary 2.4] reverses for concave functions.
-Lemma 7.2.
-Let $A_{1},\dots,A_{n}\in\mathbb{M}_{n}$ be positive semidefinite and let $\alpha_{1},\dots,\alpha_{n}$ be positive real numbers such that $\sum_{j=1}^{n}\alpha_{j}=1$. Then
-  1. (1)
-    for every monotone convex function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\leq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
-    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\geq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
-    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\leq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*};$
-  2. (2)
-    for every monotone concave function $f:[0,\infty)\to\mathbb{R}$ with $f(0)\geq 0$, there exist unitaries $U,U_{1},\dots,U_{n}\in\mathbb{M}_{n}$ such that
-    $f\!\Bigl(\sum_{j=1}^{n}A_{j}\Bigr)\$$\leq\ \sum_{j=1}^{n}U_{j}f(A_{j})U_{j}^{*},$
-    $f\!\Bigl(\sum_{j=1}^{n}\alpha_{j}A_{j}\Bigr)\$$\geq\ U\Bigl(\sum_{j=1}^{n}\alpha_{j}f(A_{j})\Bigr)U^{*}.$
-We remark that Theorems 1.19, 1.26, and 7.10 can in fact be derived directly from the results of the author in [21, Theorem 1.12 and Corollary 1.14].
-For the reader’s convenience, and to keep the present paper self-contained, we nevertheless include complete proofs.
-Proof of Theorem 1.26.
-Let $r:=p/2$ and define $f(t):=t^{r}$ on $0,\infty)$.
-Note that $f$ is monotone.
-If $p\geq 2$, then $r\geq 1$ and $f$ is convex with $f(0)=0\leq 0$.
-If $0<p\leq 2$, then $0<r\leq 1$ and $f$ is concave with $f(0)=0\geq 0$.
-Set
-$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
-and
-$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
-By Euler’s identity (1.6),
-$S:=X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}.$
-Moreover, $f(X_{1})=|A+B|^{p}$, $f(X_{2})=|B+C|^{p}$, $f(X_{3})=|C+A|^{p}$, and $f(Y_{1})=|A+B+C|^{p}$, $f(Y_{2})=|A|^{p}$, $f(Y_{3})=|B|^{p}$, $f(Y_{4})=|C|^{p}$.
-Step 1: case $p\geq 2$. Apply Lemma 7.2 (1) the first inequality to $X_{1},X_{2},X_{3}$. Then there exist unitaries $W_{1},W_{2},W_{3}\in\mathbb{M}_{n}$ such that
-$\sum_{j=1}^{3}W_{j}f(X_{j})W_{j}^{*}\ \leq\ f(S).$ (7.1)
-Next apply Lemma 7.2 (1) the second inequality to $Y_{1},Y_{2},Y_{3},Y_{4}$ with weights $\alpha_{j}=1/4$. Then there exists a unitary $U\in\mathbb{M}_{n}$ such that
-$f\!\Bigl(\frac{1}{4}S\Bigr)\ \leq\ U\Bigl(\frac{1}{4}\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
-Using $f(tX)=t^{r}f(X)$ for $t>0$, we obtain
-$f(S)=4^{r}f\!\Bigl(\frac{1}{4}S\Bigr)\ \leq\ 4^{r-1}\,U\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
-Conjugating by $U^{*}$ yields
-$U^{*}f(S)U\ \leq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k}).$ (7.2)
-Now conjugate (7.1) by $U^{*}$ and set $U_{j}:=U^{*}W_{j}$ to get
-$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \leq\ U^{*}f(S)U.$
-Combining with (7.2) gives
-$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \leq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k})=2^{p-2}\bigl(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\bigr),$
-i.e.
-$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-2}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
-Step 2: case $0<p\leq 2$. Now apply Lemma 7.2 (2) the first inequality to $X_{1},X_{2},X_{3}$ to get
-[…]
-for some unitaries $W_{j}$. Also apply Lemma 7.2 (2) the second inequality to $Y_{1},\dots,Y_{4}$ with $\alpha_{k}=1/4$ to get a unitary $U$ such that
-$f(S)=4^{r}f\!\Bigl(\frac{1}{4}S\Bigr)\ \geq\ 4^{r-1}\,U\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)U^{*}.$
-Conjugating and combining exactly as above (setting $U_{j}:=U^{*}W_{j}$) yields the reversed inequality
-$\sum_{j=1}^{3}U_{j}f(X_{j})U_{j}^{*}\ \geq\ 4^{r-1}\sum_{k=1}^{4}f(Y_{k})=2^{p-2}\bigl(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\bigr).$
-Step 3: optimality. Take $A=B=-C\neq 0$. Then $A+B=2A$, $B+C=0$, $C+A=0$, and $A+B+C=A$. Hence the inequality becomes an equality for all $p>0$, and the constant $2^{p-2}=4^{r-1}$ is best possible. ∎
-Taking traces in the unitary-orbit inequality (or, equivalently, taking the trace norm in the unitarily invariant norm statement) immediately yields the Schatten $p$-norm formulation.
-Proof of Theorem 1.19.
-Taking trace norms in Theorem 1.25, or simply taking traces in Theorem 1.26, and then passing to general Schatten class operators by approximating with finite-rank operators, yields Theorem 1.19. ∎
-7.3. Complex interpolation theory
-We now introduce the tools from complex interpolation needed in this paper; for the basic theory of interpolation, we refer the interested reader to [3, Chapters 1–5].
-Definition 7.3.
-Let $A_{0}$ and $A_{1}$ be two topological vector spaces.
-We say that $A_{0}$ and $A_{1}$ are compatible if there exists a Hausdorff topological vector space $\mathcal{Z}$ such that $A_{0}$ and $A_{1}$ are subspaces of $\mathcal{Z}$.
-In this case, we can form their sum $A_{0}+A_{1}$ and their intersection $A_{0}\cap A_{1}$.
-Definition 7.4.
-Let $X$ and $Y$ be normed spaces, and let $T:X\to Y$ be a bounded linear operator. The operator norm of $T$ is defined by
-$\|T\|_{X\to Y}:=\sup_{\begin{subarray}{c}x\in X\\ x\neq 0\end{subarray}}\frac{\|Tx\|_{Y}}{\|x\|_{X}}.$
-Definition 7.5.
-Let $(X_{0},X_{1})$ be a compatible Banach couple and $0<\theta<1$. The space $(X_{0},X_{1})_{[\theta]}$ is defined as
-$(X_{0},X_{1})_{[\theta]}:=\{f(\theta):f\in\mathcal{F}(X_{0},X_{1})\},$
-where $\mathcal{F}(X_{0},X_{1})$ denotes the space of all bounded continuous functions $f:\{z\in\mathbb{C}:0\leq\Re z\leq 1\}\to X_{0}+X_{1}$ which are analytic on $\{z\in\mathbb{C}:0<\Re z<1\}$, satisfy $f(it)\in X_{0}$ for all $t\in\mathbb{R}$ and $f(1+it)\in X_{1}$ for all $t\in\mathbb{R}$, and are bounded on the boundary lines in the respective norms of $X_{0}$ and $X_{1}$.
-The norm on $(X_{0},X_{1})_{[\theta]}$ is given by
-$\|x\|_{(X_{0},X_{1})_{[\theta]}}:=\inf\left\{\|f\|_{\mathcal{F}(X_{0},X_{1})}:f(\theta)=x\right\}.$
-We recall the abstract complex interpolation theorem for Banach couples; see [3, Theorem 4.1.2].
-Theorem 7.6 ([3]).
-Let $(X_{0},X_{1})$ and $(Y_{0},Y_{1})$ be two compatible couples of Banach spaces, and let $0<\theta<1$. Suppose that
-$T:X_{0}+X_{1}\longrightarrow Y_{0}+Y_{1}$
-is a linear operator such that $T:X_{j}\to Y_{j}$ is bounded for $j=0,1$. Then $T$ is bounded from $(X_{0},X_{1})_{[\theta]}$ to $(Y_{0},Y_{1})_{[\theta]}$, and
-$\|T\|_{(X_{0},X_{1})_{[\theta]}\to(Y_{0},Y_{1})_{[\theta]}}\leq\|T\|_{X_{0}\to Y_{0}}^{\,1-\theta}\,\|T\|_{X_{1}\to Y_{1}}^{\,\theta}.$
-We will also use the classical Riesz–Thorin interpolation theorem.
-For $L^{p}$ spaces, see [3, Theorem 5.1.1].
-For certain noncommutative $L^{p}$ spaces, such as the Schatten $p$-classes $S_{p}$, a Riesz–Thorin interpolation theorem is also known to hold; see [18, Eq.
-(2.1)], since matrix algebras are semifinite von Neumann algebras.
-See also [13, Theorem 13.1], [19, Theorems 2.9–2.10], and [19, Remark 1, p.
-23].
-Theorem 7.7 ([18]).
-Assume that $1\leq p_{0},p_{1}\leq\infty$ and $0<\theta<1$. Then
-$(S_{p_{0}},S_{p_{1}})_{[\theta]}=S_{p}\quad\text{(with equal norms)},$
-if
-$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}.$
-We shall also use the complex interpolation formula for $\ell_{p}$-sums; see [3, Theorem 5.6.3, p.
-123].
-[^1] Following [3, p.
-121, Section 5.6], let $A$ be a Banach space, let $s\in\mathbb{R}$ and $q>0$.
-Denote by $\dot{\ell}_{q}^{\,s}(A)$ the space of all bi-infinite sequences $a=(a_{\nu})_{\nu\in\mathbb{Z}}$ with $a_{\nu}\in A$ such that $\|a\|_{\dot{\ell}_{q}^{\,s}(A)}:=\Bigg(\sum_{\nu\in\mathbb{Z}}\bigl(2^{\nu s}\|a_{\nu}\|_{A}\bigr)^{q}\Bigg)^{1/q}<\infty.$ In particular, taking $s=0$ and fixing $m\in\mathbb{N}$, we may identify an $m$-tuple $(a_{1},\dots,a_{m})$ with an element of this sequence space by extending it to a bi-infinite sequence $a=(a_{\nu})_{\nu\in\mathbb{Z}}$ via $a_{\nu}=0$ for all $\nu\notin\{1,\dots,m\}$; this agrees with our definition in Theorem 7.8.
-Theorem 7.8 ([3]).
-Let $(X_{0},X_{1})$ be a compatible couple of Banach spaces.
-Assume that $1\leq p_{0},p_{1}\leq\infty$ and $0<\theta<1$.
-For each $m\in\mathbb{N}$ and $k=0,1$, define $\ell_{p_{k}}^{m}(X_{k})$ as the space of $m$-tuples $a=(a_{1},\dots,a_{m})$ with $a_{j}\in X_{k}$, equipped with the norm
-$\|a\|_{\ell_{p_{k}}^{m}(X_{k})}=\begin{cases}\big(\sum_{j=1}^{m}\|a_{j}\|_{X_{k}}^{p_{k}}\big)^{1/p_{k}},&p_{k}<\infty,\\[5.69054pt] \max_{1\leq j\leq m}\|a_{j}\|_{X_{k}},&p_{k}=\infty.\end{cases}$
-Then
-[…]
-if
-$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}.$
-Combining Theorem 7.7 and Theorem 7.8 gives the following.
-Theorem 7.9.
-Assume that $1\leq p_{0},p_{1},q_{0},q_{1}\leq\infty$ and $0<\theta<1$. Then for each $m\in\mathbb{N}$,
-[…]
-if
-$\frac{1}{p}=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}}\quad\text{ and }\quad\frac{1}{q}=\frac{1-\theta}{q_{0}}+\frac{\theta}{q_{1}}.$
-With the interpolation tools in hand, we proceed to prove Theorem 1.23.
-We first obtain (1.8) for $1<p\leq 2$ via complex interpolation between the endpoints $(p,q)=(1,\infty)$ and $(2,2)$, and then deduce (1.9) from (1.8) applied to $U^{*}$.
-Proof of Theorem 1.23.
-We first prove (1.8) for $1<p\leq 2$ by complex interpolation between the endpoints $(p,q)=(1,\infty)$ and $(p,q)=(2,2)$, and then deduce (1.9) from (1.8) applied to $U^{*}$.
-Step 1: endpoint 1: $\ell_{1}^{t}(S_{1})\to\ell_{\infty}^{s}(S_{1})$. For each $i$,
-$\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{1}\leq\sum_{j=1}^{t}|u_{ij}|\,\|Z_{j}\|_{1}\leq\mu\sum_{j=1}^{t}\|Z_{j}\|_{1}.$
-Taking the maximum over $i$ yields
-$\|T_{U}(Z)\|_{\ell_{\infty}^{s}(S_{1})}\leq\mu\,\|Z\|_{\ell_{1}^{t}(S_{1})},$
-hence
-$\|T_{U}\|_{\ell_{1}^{t}(S_{1})\to\ell_{\infty}^{s}(S_{1})}\leq\mu.$ (7.3)
-Step 2: endpoint 2: $\ell_{2}^{t}(S_{2})\to\ell_{2}^{s}(S_{2})$.
-Equip $S_{2}$ with the Hilbert–Schmidt inner product $\langle X,Y\rangle=\mathrm{Tr}(Y^{*}X)$, so that $\|X\|_{2}^{2}=\langle X,X\rangle$.
-A direct computation gives
-$\|T_{U}(Z)\|_{\ell_{2}^{s}(S_{2})}^{2}=\sum_{j,k=1}^{t}(U^{*}U)_{jk}\,\langle Z_{k},Z_{j}\rangle=\operatorname{Tr}(U^{*}U\,G),\qquad G=[\langle Z_{k},Z_{j}\rangle]_{j,k=1}^{t}\geq 0.$
-If $\|U\|_{\infty}\leq 1$, then $U^{*}U\leq I_{t}$, and therefore
-$\|T_{U}(Z)\|_{\ell_{2}^{s}(S_{2})}^{2}\leq\sum_{j=1}^{t}\|Z_{j}\|_{2}^{2}=\|Z\|_{\ell_{2}^{t}(S_{2})}^{2},$
-so that
-$\|T_{U}\|_{\ell_{2}^{t}(S_{2})\to\ell_{2}^{s}(S_{2})}\leq 1.$ (7.4)
-Step 3: interpolation. Fix $0<\theta<1$. Applying Theorem 7.6 to the couples
-$X_{0}=\ell_{1}^{t}(S_{1}),\quad X_{1}=\ell_{2}^{t}(S_{2}),\qquad Y_{0}=\ell_{\infty}^{s}(S_{1}),\quad Y_{1}=\ell_{2}^{s}(S_{2}),$
-and using (7.3)–(7.4), we obtain
-$\|T_{U}\|_{(X_{0},X_{1})_{[\theta]}\to(Y_{0},Y_{1})_{[\theta]}}\leq\mu^{\,1-\theta}.$ (7.5)
-By Theorem 7.9,
-$(X_{0},X_{1})_{[\theta]}=\ell_{p}^{t}(S_{p}),\qquad(Y_{0},Y_{1})_{[\theta]}=\ell_{q}^{s}(S_{p}),$
-where
-$\frac{1}{p}=1-\frac{\theta}{2},\qquad\frac{1}{q}=\frac{\theta}{2},$
-so that $q$ is the conjugate exponent of $p$ and $1<p\leq 2$. Thus (7.5) yields
-$\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\leq\mu^{\,1-\theta}\,\|Z\|_{\ell_{p}^{t}(S_{p})}=\mu^{\,\frac{2}{p}-1}\,\|Z\|_{\ell_{p}^{t}(S_{p})},$
-which is exactly (1.8).
-Step 4: the case $p\geq 2$ (reverse inequality in (1.8)).
-Assume in addition that $U^{*}U=I_{t}$ (as required in the statement for the reverse direction).
-Let $q$ be the conjugate exponent of $p$, so that $1<q\leq 2$.
-Let $Z=(Z_{1},\dots,Z_{t})\in\ell_{p}^{t}(S_{p})$. By duality,
-$\|Z\|_{\ell_{p}^{t}(S_{p})}=\sup\Big\{\,|\langle Z,W\rangle|:\ W\in\ell_{q}^{t}(S_{q}),\ \|W\|_{\ell_{q}^{t}(S_{q})}\leq 1\,\Big\},\qquad\langle Z,W\rangle:=\sum_{j=1}^{t}\operatorname{Tr}(W_{j}^{*}Z_{j}).$
-For such a $W$, since $U^{*}U=I_{t}$ we have $T_{U^{*}}\circ T_{U}=\mathrm{Id}$, hence $Z=T_{U^{*}}(T_{U}(Z))$ and therefore
-$\langle Z,W\rangle=\langle T_{U^{*}}(T_{U}(Z)),\,W\rangle=\langle T_{U}(Z),\,T_{U}(W)\rangle.$
-By Hölder’s inequality for Schatten classes and for $\ell_{p}$-sums,
-$|\langle T_{U}(Z),T_{U}(W)\rangle|\leq\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\ \|T_{U}(W)\|_{\ell_{p}^{s}(S_{q})}.$
-Since $1<q\leq 2$ and $\|U\|_{\infty}\leq 1$, we may apply the already proved estimate (1.8) (with $q$ in place of $p$ and $p$ in place of $q$) to obtain
-$\|T_{U}(W)\|_{\ell_{p}^{s}(S_{q})}\leq\mu^{\frac{2}{q}-1}\,\|W\|_{\ell_{q}^{t}(S_{q})}=\mu^{\,1-\frac{2}{p}}\,\|W\|_{\ell_{q}^{t}(S_{q})}.$
-Hence, for $\|W\|_{\ell_{q}^{t}(S_{q})}\leq 1$,
-$|\langle Z,W\rangle|\leq\mu^{\,1-\frac{2}{p}}\,\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}.$
-Taking the supremum over all such $W$ yields
-$\|Z\|_{\ell_{p}^{t}(S_{p})}\leq\mu^{\,1-\frac{2}{p}}\,\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})},$
-equivalently,
-$\|T_{U}(Z)\|_{\ell_{q}^{s}(S_{p})}\geq\mu^{\,\frac{2}{p}-1}\,\|Z\|_{\ell_{p}^{t}(S_{p})}.$
-This is exactly the reversed inequality in (1.8) for $p\geq 2$.
-Step 5: the reverse-direction estimate. Assume now that $U^{*}U=I_{t}$. Consider the adjoint matrix $U^{*}\in\mathbb{M}_{t,s}$ and the associated map
-$T_{U^{*}}:\ \ell_{p}^{s}(S_{p})\longrightarrow\ell_{p}^{t}(S_{p}),\qquad T_{U^{*}}(W_{1},\dots,W_{s}):=\Big(\sum_{i=1}^{s}\overline{u_{i1}}\,W_{i},\ \dots,\ \sum_{i=1}^{s}\overline{u_{it}}\,W_{i}\Big).$
-Then $T_{U^{*}}\circ T_{U}=\mathrm{Id}$ on $\ell_{p}^{t}(S_{p})$. Applying (1.8) to $U^{*}$ (note that $\max_{i,j}|(U^{*})_{ij}|=\mu$) gives, for $1<p\leq 2$,
-$\|T_{U^{*}}(W)\|_{\ell_{q}^{t}(S_{p})}\leq\mu^{\,\frac{2}{p}-1}\,\|W\|_{\ell_{p}^{s}(S_{p})}\qquad(W\in\ell_{p}^{s}(S_{p})).$
-Taking $W=T_{U}(Z)$ and using $T_{U^{*}}(T_{U}(Z))=Z$ yields (1.9). The case $p\geq 2$ again follows by duality. ∎
-Proof of Theorem 1.21.
-Fix $1<p\leq 2$ and let $q$ be the conjugate exponent of $p$, i.e. $1/p+1/q=1$. Set
-$X_{1}:=A+B,\quad X_{2}:=B+C,\quad X_{3}:=C+A,\qquad Y_{1}:=A+B+C,\ Y_{2}:=A,\ Y_{3}:=B,\ Y_{4}:=C.$
-Step 1: A suborthogonal coefficient matrix with $U^{*}$ isometric. Consider the real $3\times 4$ matrix
-$U:=\frac{1}{2}\begin{pmatrix}1&1&1&-1\\ 1&-1&1&1\\ 1&1&-1&1\end{pmatrix}$.
-Then
-$\begin{pmatrix}X_{1}\\ X_{2}\\ X_{3}\end{pmatrix}=U\begin{pmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\\ Y_{4}\end{pmatrix},\qquad\text{i.e.}\qquad X_{i}=\sum_{j=1}^{4}u_{ij}Y_{j}\ (i=1,2,3).$
-Moreover, a direct computation gives $UU^{*}=I_{3}$; hence $U^{*}$ is an isometry. Also $\max_{i,j}|u_{ij}|=1/2=:\mu$.
-Step 2: The relations $X=UY$ and $Y=U^{*}X$. By construction,
-$\begin{pmatrix}X_{1}\\ X_{2}\\ X_{3}\end{pmatrix}=U\begin{pmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\\ Y_{4}\end{pmatrix}$.
-Moreover, a direct computation gives $Y=U^{*}X$ and explicitly
-$Y_{1}=\frac{X_{1}+X_{2}+X_{3}}{2},\quad Y_{2}=\frac{X_{1}-X_{2}+X_{3}}{2},\quad Y_{3}=\frac{X_{1}+X_{2}-X_{3}}{2},\quad Y_{4}=\frac{-X_{1}+X_{2}+X_{3}}{2},$
-which are easily verified from $X_{1}=A+B$, $X_{2}=B+C$, $X_{3}=C+A$.
-Step 3: Apply Theorem 1.23. Since $\|U\|_{\infty}\leq 1$, Theorem 1.23 (i) gives for $1<p\leq 2$,
-$\Big(\|X_{1}\|_{p}^{q}+\|X_{2}\|_{p}^{q}+\|X_{3}\|_{p}^{q}\Big)^{1/q}\leq\mu^{\frac{2}{p}-1}\Big(\|Y_{1}\|_{p}^{p}+\|Y_{2}\|_{p}^{p}+\|Y_{3}\|_{p}^{p}+\|Y_{4}\|_{p}^{p}\Big)^{1/p}.$
-Raising both sides to the power $q$ and using $\mu=1/2$ yields
-$\|X_{1}\|_{p}^{q}+\|X_{2}\|_{p}^{q}+\|X_{3}\|_{p}^{q}\leq 2^{\,1-\frac{q}{p}}\Big(\|Y_{1}\|_{p}^{p}+\|Y_{2}\|_{p}^{p}+\|Y_{3}\|_{p}^{p}+\|Y_{4}\|_{p}^{p}\Big)^{q/p},$
-which is the second inequality.
-Applying the same estimate to $U^{*}$ and using $Y=U^{*}X$ gives
-$\Big(\|Y_{1}\|_{p}^{q}+\|Y_{2}\|_{p}^{q}+\|Y_{3}\|_{p}^{q}+\|Y_{4}\|_{p}^{q}\Big)^{1/q}\leq\mu^{\frac{2}{p}-1}\Big(\|X_{1}\|_{p}^{p}+\|X_{2}\|_{p}^{p}+\|X_{3}\|_{p}^{p}\Big)^{1/p},$
-and raising to $q$ again yields the first inequality with the same constant $2^{\,1-\frac{q}{p}}$.
-Step 4: optimality. Taking $A=B=-C\neq 0$ gives $A+B=2A$ and $B+C=C+A=0$, while $A+B+C=A$. Thus equality holds for all $p>1$, and the constant $2^{1-\frac{q}{p}}$ is optimal.
-7.4. Derivation of Audenaert–Kittaneh’s inequality and its sharp complement
-In this subsection, we show that the Audenaert–Kittaneh inequality (1.10) and the complementary estimate (1.11) both follow from Theorem 1.23 by a suitable choice of the coefficient matrix $U$.
-Let $n\geq 2$ and set
-$s:=1+\binom{n}{2}$.
-Fix an enumeration of the pairs $(i,j)$ with $1\leq i<j\leq n$:
-$(i_{1},j_{1}),\ (i_{2},j_{2}),\ \ldots,\ (i_{\binom{n}{2}},j_{\binom{n}{2}})$.
-Define
-$U=\frac{1}{\sqrt{n}}\begin{pmatrix}\mathbf{1}^{\top}\\ (e_{i_{1}}^{\top}-e_{j_{1}}^{\top})\\ (e_{i_{2}}^{\top}-e_{j_{2}}^{\top})\\ \vdots\\ (e_{i_{\binom{n}{2}}}^{\top}-e_{j_{\binom{n}{2}}}^{\top})\end{pmatrix}\in\mathbb{M}_{s,n},$
-where $\mathbf{1}=(1,\dots,1)^{\top}\in\mathbb{C}^{n}$ and $e_{k}$ denotes the $k$th standard basis vector in $\mathbb{C}^{n}$. Then
-$\mu:=\max_{\alpha,j}|u_{\alpha j}|=\frac{1}{\sqrt{n}}.$
-We claim $U$ is an isometry, i.e. $U^{*}U=I_{n}$ (hence $\|U\|_{\infty}=1$ and $U$ is a contraction). Indeed,
-$U^{*}U=\frac{1}{n}\left(\mathbf{1}\mathbf{1}^{*}+\sum_{1\leq i<j\leq n}(e_{i}-e_{j})(e_{i}-e_{j})^{*}\right),$
-and one checks the identity
-$\sum_{1\leq i<j\leq n}(e_{i}-e_{j})(e_{i}-e_{j})^{*}=nI_{n}-\mathbf{1}\mathbf{1}^{*}.$
-Therefore $U^{*}U=\frac{1}{n}(\mathbf{1}\mathbf{1}^{*}+nI_{n}-\mathbf{1}\mathbf{1}^{*})=I_{n}$.
-Given $A_{1},\dots,A_{n}\in S_{p}$, consider the $n$-tuple $Z:=(A_{1},\dots,A_{n})\in\ell_{p}^{n}(S_{p})$. By construction,
-$T_{U}(Z)=\Big(\tfrac{1}{\sqrt{n}}\sum_{k=1}^{n}A_{k},\ \tfrac{1}{\sqrt{n}}(A_{i_{1}}-A_{j_{1}}),\ \dots,\ \tfrac{1}{\sqrt{n}}(A_{i_{\binom{n}{2}}}-A_{j_{\binom{n}{2}}})\Big)\in\ell_{q}^{s}(S_{p}).$
-Consequently, for $r\in\{p,q\}$ we have
-$\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{r}=\frac{1}{n^{r/2}}\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{r}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{r}\right).$ (7.6)
-In particular, (7.6) with $r=q$ gives the left side of (1.8), and with $r=p$ gives the right side of (1.9).
-Step 1: Audenaert–Kittaneh inequality from Theorem 1.23(i). Let $1<p\leq 2$ and let $q$ be the conjugate exponent of $p$. Since $U$ is a contraction, we may apply (1.8) to $Z=(A_{1},\dots,A_{n})$:
-$\Bigg(\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\leq\mu^{\frac{2}{p}-1}\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{1/p}.$
-Using (7.6) with $r=q$ and $\mu=1/\sqrt{n}$, we obtain
-$\frac{1}{\sqrt{n}}\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\right)^{1/q}\leq\Big(\frac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{1/p}.$
-Multiplying by $\sqrt{n}$ and noting that
-$\sqrt{n}\Big(\tfrac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}=n^{1-\frac{1}{p}},\qquad\text{and}\qquad q\Bigl(1-\frac{1}{p}\Bigr)=1,$
-we get
-$\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{p}\Bigg)^{q/p},$
-which is exactly (1.10). For $p\geq 2$, the inequality is reversed by Theorem 1.23(i) since $U^{*}U=I_{n}$.
-Step 2: the complementary estimate from Theorem 1.23(ii). Since $U^{*}U=I_{n}$, we may apply (1.9) to the same $U$ and $Z$. For $1<p\leq 2$,
-$\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{q}\Bigg)^{1/q}\leq\mu^{\frac{2}{p}-1}\Bigg(\sum_{\alpha=1}^{s}\Big\|\sum_{j=1}^{n}u_{\alpha j}A_{j}\Big\|_{p}^{p}\Bigg)^{1/p}.$
-Using (7.6) with $r=p$ and $\mu=1/\sqrt{n}$ yields
-$\Bigg(\sum_{j=1}^{n}\|A_{j}\|_{p}^{q}\Bigg)^{1/q}\leq\Big(\frac{1}{\sqrt{n}}\Big)^{\frac{2}{p}-1}\cdot\Big(\frac{1}{\sqrt{n}}\Big)\left(\Big\|\sum_{k=1}^{n}A_{k}\Big\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{1/p}.$
-Since $\big(\tfrac{1}{\sqrt{n}}\big)^{\frac{2}{p}-1}\cdot\tfrac{1}{\sqrt{n}}=n^{-1/p}$, raising both sides to the power $q$ gives (1.11). For $p\geq 2$, the inequality is reversed by Theorem 1.23(ii).
-Step 3: optimality of the constant in (1.11). Let $A_{1}=\cdots=A_{n}=A\neq 0$. Then $\sum_{k=1}^{n}A_{k}=nA$ and $A_{i}-A_{j}=0$, hence
-[…]
-Thus equality holds for all $1<p<\infty$, and the constant $n^{-q/p}$ is best possible.
-7.5. A weaker bound on Conjecture 1.20
-We record a non-sharp estimate with the explicit constant $3^{\frac{p}{2}-1}$, which follows directly from the Euler-type identities; however, this coefficient is not optimal.
-Theorem 7.10.
-Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
-$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\ \leq\ 3^{\frac{p}{2}-1}\Bigl(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Bigr).$
-For $0<p\leq 2$, the inequality is reversed.
-Proof.
-Set $r:=p/2$ and $f(t):=t^{r}$ on $0,\infty)$. Let
-$X_{1}:=|A+B|^{2},\quad X_{2}:=|B+C|^{2},\quad X_{3}:=|C+A|^{2},$
-and
-$Y_{1}:=|A+B+C|^{2},\quad Y_{2}:=|A|^{2},\quad Y_{3}:=|B|^{2},\quad Y_{4}:=|C|^{2}.$
-By Euler’s identity (1.6),
-$S:=X_{1}+X_{2}+X_{3}=Y_{1}+Y_{2}+Y_{3}+Y_{4}.$
-Assume first that $p\geq 2$, i.e.
-$r\geq 1$, so that $f$ is convex and $f(0)=0$.
-Using the trace form of the Jensen/Rotfel’d type inequality for convex $f$ (equivalently, Lemma 7.1 (1) with the trace norm), we have
-$\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)\ \leq\ \operatorname{Tr}\bigl(f(S)\bigr).$ (7.7)
-Next apply Jensen’s inequality with weights $\alpha_{j}=1/3$ to $X_{1},X_{2},X_{3}$:
-$\operatorname{Tr}\Bigl(f\!\bigl(\tfrac{1}{3}S\bigr)\Bigr)\leq\operatorname{Tr}\Bigl(\tfrac{1}{3}\sum_{j=1}^{3}f(X_{j})\Bigr).$
-Since $f(tX)=t^{r}f(X)$ for $t>0$, this becomes
-$3^{-r}\operatorname{Tr}\bigl(f(S)\bigr)\ \leq\ \tfrac{1}{3}\,\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr),$
-or equivalently
-$\operatorname{Tr}\bigl(f(S)\bigr)\ \leq\ 3^{r-1}\,\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr).$ (7.8)
-Combining (7.7)–(7.8) gives
-$\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}|Y_{k}|^{r}\Bigr)=\operatorname{Tr}\!\Bigl(\sum_{k=1}^{4}f(Y_{k})\Bigr)\leq 3^{r-1}\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}f(X_{j})\Bigr)=3^{r-1}\operatorname{Tr}\!\Bigl(\sum_{j=1}^{3}|X_{j}|^{r}\Bigr).$
-Finally, passing to general Schatten class operators by approximating with finite-rank operators, we obtain exactly the desired inequality with constant $3^{r-1}=3^{\frac{p}{2}-1}$.
-For $0<p\leq 2$ (i.e. $0<r\leq 1$), the function $f$ is concave and all inequalities above reverse, yielding the reversed statement. ∎
 Remark 11.5.
-The constant $3^{\frac{p}{2}-1}$ in Theorem 7.10 is generally not optimal; numerical evidence suggests that the sharp constant should be the one proposed in Conjecture 1.20.
-8. Unitary-orbit refinements of Euler’s identity and spectral consequences
-8.1. A unitary-orbit bound for the Euler modulus
-To convert the square-sum identity (1.6) into a linear (modulus) estimate in unitary orbits, we use the following unitary subadditivity result for concave functions due to Aujla and Bourin [2, Theorem 2.1], which is also included in Lemma 7.2 (2) the first inequality.
-Lemma 8.1 (Aujla–Bourin).
-Let $H,K\in\mathbb{M}_{n}$ be positive semidefinite and let $f:[0,\infty)\to[0,\infty)$ be monotone concave with $f(0)\geq 0$. Then there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
-$f(H+K)\leq Uf(H)U^{*}+Vf(K)V^{*}.$
-We now apply Lemma 8.1 with the concave function $f(t)=t^{1/2}$ to the Euler square-sum relation (1.6). This yields a three-term unitary-orbit domination of the Euler modulus by the pairwise moduli.
-Theorem 8.2.
-Let $A,B,C\in\mathbb{M}_{n}$. Then there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
-$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$ (8.1)
-Proof.
-By (1.6),
-$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}=\sqrt{|A+B|^{2}+|B+C|^{2}+|C+A|^{2}}.$
-Apply Lemma 8.1 with $f(t)=\sqrt{t}$ to $H=|A+B|^{2}$ and $K=|B+C|^{2}+|C+A|^{2}$ to get unitaries $U_{1},V_{1}$ with
-$\sqrt{H+K}\leq U_{1}|A+B|U_{1}^{*}+V_{1}\sqrt{K}V_{1}^{*}.$
-Apply Lemma 8.1 again to $|B+C|^{2}$ and $|C+A|^{2}$ to obtain unitaries $U_{2},V_{2}$ such that
-$\sqrt{K}\leq U_{2}|B+C|U_{2}^{*}+V_{2}|C+A|V_{2}^{*}.$
-Substitute and absorb conjugation by $V_{1}$ into $U_{2},V_{2}$ to obtain (8.1). ∎
-As a first immediate consequence, taking any unitarily invariant norm in (8.1) and using the triangle inequality yields a norm inequality.
-Corollary 8.3.
-Let $A,B,C\in\mathbb{M}_{n}$. Then
-[…]
-We next record a singular-value refinement, obtained by combining Theorem 8.2 with the Weyl inequality for sums of positive semidefinite matrices.
-Corollary 8.4.
-Let $A,B,C\in\mathbb{M}_{n}$. Then for all integers $j,k,\ell\in\{0,1,\dots,n-1\}$,
-$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j}^{\downarrow}(A+B)+\mu_{1+k}^{\downarrow}(B+C)+\mu_{1+\ell}^{\downarrow}(C+A).$
-Proof.
-By Theorem 8.2, there exist unitaries $U,V,W\in\mathbb{M}_{n}$ such that
-$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq U|A+B|U^{*}+V|B+C|V^{*}+W|C+A|W^{*}.$
-Set $X:=U|A+B|U^{*}$, $Y:=V|B+C|V^{*}$ and $Z:=W|C+A|W^{*}$. Then $X,Y,Z\geq 0$ and $\mu_{j}^{\downarrow}(X)=\mu_{j}^{\downarrow}(|A+B|)=\mu_{j}^{\downarrow}(A+B)$, and similarly for $Y,Z$.
-We use the following singular-value inequality (see, e.g., [4]): for $P,Q\geq 0$ and $r,s\geq 0$,
-$\mu_{1+r+s}^{\downarrow}(P+Q)\leq\mu_{1+r}^{\downarrow}(P)+\mu_{1+s}^{\downarrow}(Q).$ (8.2)
-Applying (8.2) to $P=X$ and $Q=Y+Z$ gives
-$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k+\ell}^{\downarrow}(Y+Z).$
-Next apply (8.2) to $P=Y$ and $Q=Z$ to obtain
-$\mu_{1+k+\ell}^{\downarrow}(Y+Z)\leq\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
-Combining the last two inequalities yields
-$\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z)\leq\mu_{1+j}^{\downarrow}(X)+\mu_{1+k}^{\downarrow}(Y)+\mu_{1+\ell}^{\downarrow}(Z).$
-Finally, since
-$\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\leq X+Y+Z$
-and both sides are positive semidefinite, the monotonicity of singular values gives
-$\mu_{1+j+k+\ell}^{\downarrow}\!\Big(\sqrt{|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}}\Big)\leq\mu_{1+j+k+\ell}^{\downarrow}(X+Y+Z),$
-which implies the desired estimate. ∎
-8.2. Symmetric norms, antinorms, and Ky Fan consequences
-Finally, since Euler’s identity (1.6) is an exact equality of positive operators, it also yields further consequences for symmetric norms and antinorms (in the sense of [8]) on the positive semidefinite cone of $\mathbb{M}_{n}$, and in particular for Ky Fan sums of singular values.
-Corollary 8.5.
-Let $A,B,C\in\mathbb{M}_{n}$.
-  1. (a)
-[…]
-[…]
-  2. (b)
-    For every symmetric antinorm $\|\cdot\|_{!}$,
-    $\Big\||A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big\|_{!}\geq\big\||A+B|^{2}\big\|_{!}+\big\||B+C|^{2}\big\|_{!}+\big\||C+A|^{2}\big\|_{!}.$
-Applying Corollary 8.5 to Ky Fan $m$-norms and Ky Fan $m$-antinorms gives the following eigenvalue-sum inequalities.
-Corollary 8.6.
-Let $A,B,C\in\mathbb{M}_{n}$. Then for $m=1,\dots,n$,
-$\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\leq\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\downarrow}\!\big(|C+A|^{2}\big),$
-$\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\Big(|A+B+C|^{2}+|A|^{2}+|B|^{2}+|C|^{2}\Big)$ $\geq\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|A+B|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|B+C|^{2}\big)+\sum_{j=1}^{m}\mu_{j}^{\uparrow}\!\big(|C+A|^{2}\big).$
-Acknowledgments
-The author is very grateful to Professor J.-C.
-Bourin for reading this manuscript and for his valuable comments.
-This work is supported by the China Scholarship Council, the Young Elite Scientists Sponsorship Program for PhD Students (China Association for Science and Technology), and the Fundamental Research Funds for the Central Universities at Xi’an Jiaotong University (Grant No.
-xzy022024045).
+In the proof of Theorem 1.39 above, the choice of the unitarily conjugating matrix is not unique. For instance, we could pick a Hadamard matrix
+$H_{4}=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{pmatrix}$
+instead of the $4\times 4$ Fourier transform $F_{4}$: in such a case, dealing with real matrices yields real isometries.